### v1

# Exploring Test-Time Adaptation for Object Detection in Continually Changing Environments

Abstract.
For real-world applications, neural network models are commonly deployed in dynamic environments, where the distribution of the target domain undergoes temporal changes. Continual Test-Time Adaptation (CTTA) has recently emerged as a promising technique to gradually adapt a source-trained model to test data drawn from a continually changing target domain. Despite recent advancements in addressing CTTA, two critical issues remain: 1) The use of a fixed threshold for pseudo-labeling in existing methodologies leads to the generation of low-quality pseudo-labels, as model confidence varies across categories and domains; 2) While current solutions utilize stochastic parameter restoration to mitigate catastrophic forgetting, their capacity to preserve critical information is undermined by its intrinsic randomness. To tackle these challenges, we present CTAOD, aiming to enhance the performance of detection models in CTTA scenarios. Inspired by prior CTTA works for effective adaptation, CTAOD is founded on the mean-teacher framework, characterized by three core components. Firstly, the object-level contrastive learning module tailored for object detection extracts object-level features using the teacher’s region of interest features and optimizes them through contrastive learning. Secondly, the dynamic threshold strategy updates the category-specific threshold based on predicted confidence scores to improve the quality of pseudo-labels. Lastly, we design a data-driven stochastic restoration mechanism to selectively reset inactive parameters using the gradients as weights for a random mask matrix, thereby ensuring the retention of essential knowledge. We demonstrate the effectiveness of our approach on four CTTA tasks for object detection, where CTAOD outperforms existing methods, especially achieving a 3.0 mAP improvement on the Cityscapes-to-Cityscapes-C CTTA task. The code of this work will be released soon.
copyright: none
1. Introduction
Figure 1. Motivation of our proposed method. The source-trained model adapts to dynamic environments without source data. (1) DT: The bar chart above represents the mean model scores for different categories across four domains of the ACDC dataset (Sakaridis et al., 2021). While existing methods produce the pseudo-labels via a fixed threshold, CTAOD utilizes DT to update the category-specific threshold based on the mean predicted scores. (2) DSR: The bottom describes the neuron reset process, where neurons with deeper shades of purple indicate increased activity levels and blue neurons represent corresponding neurons from the source model. While existing methods restore parameters randomly, CTAOD resets inactive parameters with higher possibilities.
Deep learning models have demonstrated immense potential across various data modalities such as image (He et al., 2016; Dosovitskiy et al., 2020; Ren et al., 2015), audio (Aytar et al., 2016; Amodei et al., 2016), language (Devlin et al., 2018; Vaswani et al., 2017), and video (Tran et al., 2015; Carreira and Zisserman, 2017). However, these models experience pronounced degradation in performance when confronted with training data (i.e., source domain) and testing data (i.e., target domain) originating from disparate distributions. This phenomenon, commonly referred to as distribution shifts, poses a significant challenge (Geirhos et al., 2018; Hendrycks and Dietterich, 2019; Luo et al., 2020; Deng et al., 2014). In such a scenario, unsupervised domain adaptation (UDA) becomes crucial, which typically involves aligning the source and target data distributions, thereby mitigating the impact of distribution shifts (Ganin et al., 2016; Tzeng et al., 2017; Saito et al., 2018). Still, UDA falls short by necessitating access to source data, which is often inaccessible due to privacy constraints or data transmission barriers (VS et al., 2023b; Huang et al., 2021).
This limitation catalyzes the exploration of Test-Time Adaptation (TTA) where the source-trained model directly adapts toward unlabeled test samples encountered during evaluation in an online manner, without the reliance on the source data (Mummadi et al., 2021; Sinha et al., 2023; You et al., 2021; Wang et al., 2020). Nonetheless, the aforementioned methods, which assume a static target domain, face a more challenging and realistic problem, as real-world machine learning systems work in non-stationary and continually evolving environments (Wang et al., 2022; Döbler et al., 2023). Existing TTA methods are vulnerable to catastrophic forgetting of previously learned source knowledge and error accumulation when adaptation faces more than one distribution shift (Wang et al., 2022; Döbler et al., 2023; Niloy et al., 2024). For example, a vehicle may encounter various continuous environmental changes such as fog, night, rain, and snow during its journey.
Recently, Wang et al. (2022) introduce CoTTA by applying stochastic parameters restoration to mitigate catastrophic forgetting in Continual Test-Time Adaptation (CTTA) scenarios (see Table 1), where the model is continually adapted to sequences of target domains. Although the strategy of randomly selecting parameters for restoration partially helps alleviate the forgetting of source knowledge (Wang et al., 2022; Zhu et al., 2023), its randomness may also contribute to the loss of crucial knowledge specific to the current domain (refer to Figure 1). Additionally, existing CTTA methods normally rely on self-training methods by utilizing pseudo-labels through a fixed threshold to supervise the training of the model (Wang et al., 2022; Döbler et al., 2023; Sójka et al., 2023). Given that model confidence could vary across categories and domains (refer to Figure 1), employing a uniformly fixed threshold could result in excluding high-quality pseudo-labels while incorporating incorrect ones. Erroneous or missing pseudo-labels lead to error accumulation as they provide negative feedback to the model.
Therefore, despite recent efforts to handle CTTA, two significant challenges persist: Current self-training-based methods suffer from noisy and low-quality pseudo-labels, leading to error accumulation (Challenge 1); Continual adaptation to dynamic environments in existing methods struggles to effectively retain valuable knowledge about the current domain while mitigating noise specific to the previous domain (Challenge 2).
To tackle the identified challenges in the CTTA setting for object detection, we introduce CTAOD (Continual Test-time Adaption for Object Detection). Aligning with previous CTTA works for robust adaptation (Wang et al., 2022; Gong et al., 2022; Döbler et al., 2023), CTAOD is constructed upon the mean-teacher framework (Tarvainen and Valpola, 2017) where the teacher model is an exponential moving average of the student model. Particularly, CTAOD comprises Object-level Contrastive Learning (OCL), Dynamic Threshold (DT), and Data-driven Stochastic Restoration (DSR).
More specifically, we propose OCL, which is tailored for object detection, to acquire fine-grained and localized feature representation, motivated by the effectiveness of contrastive learning (CL) in self-supervised learning (Chen et al., 2020; He et al., 2020; Grill et al., 2020; Chen and He, 2021). OCL extracts features based on proposals generated by the Region Proposal Network (RPN), which provides multiple cropped views around the object instance at different locations and scales. Subsequently, CL loss is applied on RPN cropped views to guide the model to encourage similar object instances to remain close while pushing dissimilar ones apart. The OCL is well integrated into the mean-teacher paradigm as a drop-in enhancement for feature adaptation. Furthermore, to improve the quality of pseudo-labels across various categories, i.e., addressing Challenge 1, we devise a DT strategy to dynamically adjust the threshold for each category individually based on the predicted confidence scores. The dynamic nature of the DT method makes it better suited to address the effects of continuously changing distributions compared to fixed thresholds. Finally, the DSR mechanism is utilized to reset inactive parameters with a higher possibility than active ones by employing the gradients of parameters as weights for a random mask matrix. On the one hand, it not only helps alleviate the impact of noise from the previous domain but also preserves important information, i.e., addressing Challenge 2. On the other hand, its randomness also allows potentially falsely activated parameters to be reset, thereby making the model more stable.
We demonstrate the effectiveness of our proposed approach on four continual test-time adaptation benchmark tasks for object detection. These datasets are associated with synthetic and real-world distribution shifts in the short-term and long-term adaption (i.e. Cityscapes (Cordts et al., 2016) $\rightarrow$ Cityscapes-C (Hendrycks and Dietterich, 2019), SHIFT (Sun et al., 2022), and Cityscapes (Cordts et al., 2016) $\rightarrow$ ACDC (Sakaridis et al., 2021) adaptation). Experimental results indicate that our method significantly improves performance over existing state-of-the-art methods, with gains of up to 3.0 mAP on these tasks.
In summary, the main contributions of this paper are as follows:
  • This study introduces a novel method named CTAOD, which pioneers exploring CTTA for detection models. Specifically, we propose to leverage object-level features for contrastive learning to refine feature representation for object detection.
  • To address two challenges in CTTA, we dynamically update the category-specific threshold based on predicted scores to improve the quality of pseudo-labels and reset the inactive parameter with higher possibilities to mitigate forgetting.
  • Empirical experiments demonstrate that our proposed method surpasses existing methods and effectively facilitates short-term and long-term adaptation in dynamic environments.
2. RELATED WORK
2.1. Source-Free Domain Adaptation
UDA tackles the inter-domain divergence by aligning the distributions of source and target data (Ganin et al., 2016; Li et al., 2019; Luo et al., 2020; Jiang et al., 2020). Despite its effectiveness, the limitation of UDA lies in its requirement for access to the source domain data, which often raises concerns regarding data privacy, data portability, and data transmission efficiency. As a result, Source-Free Domain Adaptation (SFDA) received extensive research attention, especially in object detection, where the source-trained detector is re-trained on the target training data (VS et al., 2023b, a; Li et al., 2022b, 2021a, 2021b) before evaluation. For instance, MemCLR (VS et al., 2023a) employs a cross-attention transformer-based memory bank for contrastive learning, while IRG (VS et al., 2023a) utilizes the object relations with instance relation graph network to explore the SFDA setting for object detection without the source data. However, the standard SFDA setting requires prior knowledge of the target domain, which is impractical in most real-world applications.
2.2. Test-Time Adaptation
TTA adapts the source-trained model to the target test data during inference time without access to the source data. Since both TTA and SFDA involve adapting the source-trained model to the unlabeled target data without utilizing source data, some works also refer to TTA as SFDA (Wang et al., 2022; Brahma and Rai, 2023). In this paper, we distinguish between TTA and SFDA based on evaluation protocol, although they can be transformed into each other in experiments. Furthermore, TTA methods improve the model performance under distribution shift commonly through techniques like pseudo-labeling (Sinha et al., 2023; Iwasawa and Matsuo, 2021; Sun et al., 2020; Zeng et al., 2023), batchnorm statistics updating (Hu et al., 2021; You et al., 2021), or entropy regularization (Wang et al., 2020; Iwasawa and Matsuo, 2021; Niu et al., 2022) during testing. For example, Tent (Wang et al., 2020) updates the batchnorm parameters with entropy minimization and demands a large batch size for optimization during test-time adaptation, which is unsuitable for real-time detection model deployment where images are processed sequentially. The above approaches assume a specific and static target domain where all the test inputs come from a single domain. However, in numerous practical scenarios, the distribution of test input may exhibit a continual shift over time.
2.3. Continual Test-time Adaptation
Early works consider adaptation to evolving and continually changing domains by aligning the source and target data (Hoffman et al., 2014; Wulfmeier et al., 2018). These methods rely on source data during inference, which limits their applicability. A significant advancement in this area is the proposal of CoTTA (Wang et al., 2022), marking the first work tailored to the demands of CTTA by adapting a pre-trained model to sequences of domains without using any source data. Subsequently, research efforts have been dedicated to exploring CTTA, primarily focused on classification (Döbler et al., 2023; Niloy et al., 2024; Brahma and Rai, 2023; Yu et al., 2023) and segmentation (Niloy et al., 2024; Song et al., 2023; Zhu et al., 2023) tasks. For instance, similar to CoTTA (Wang et al., 2022), Zhu et al. (2023) apply a stochastic reset mechanism for CTTA in the segmentation task of medical images to prevent forgetting. In contrast, PETAL (Brahma and Rai, 2023) utilizes the Fisher Information Matrix (FIM) as a metric of parameter importance to reset only the most irrelevant parameters. Nevertheless, these two methods are sub-optimal, as the randomness might lead to losing essential information, and pure data-driven restoration may retain false active parameters resulting from noise.
Moreover, the mean-teacher framework (Tarvainen and Valpola, 2017) serves as a base architecture for most CTTA works (Wang et al., 2022; Gong et al., 2022; Döbler et al., 2023; Niloy et al., 2024), where the teacher commonly generates pseudo-labels via a fixed threshold to supervise the student. Nonetheless, these methods suffer from low-quality pseudo-labels with a uniform threshold since the model confidence varies across categories and domains. Furthermore, Wang et al. (2024) design a dynamic thresholding technique to update the threshold based on the maximum predicted score in a batch for classification tasks, which requires a relatively large batch size to set an appropriate threshold. However, it is not suitable for object detection, where a smaller batch size is often preferred for computational efficiency. Alternatively, Gan et al. (2023) presents a cloud-device collaborative continual adaptation paradigm to accommodate dynamic environment in object detection. Nevertheless, its dependency on source data also constrains its practical utility. Therefore, a gap still exists in exploring the CTTA setting in object detection to improve the quality of pseudo-labels (Challenge 1) and to effectively reset noisy neurons (Challenge 2), without reliance on the source data.
2.4. Continual Learning
Continual learning, also known as incremental learning or life-long learning, typically involves enabling the model to retain previously acquired knowledge while learning from a sequential series of tasks, i.e. preventing catastrophic forgetting (De Lange et al., 2021; Parisi et al., 2019). It is commonly categorized into replay methods (Rebuffi et al., 2017; Tiwari et al., 2022; Rolnick et al., 2019), parameter isolation method (Aljundi et al., 2017; Xu and Zhu, 2018), and regularization-based methods (Kirkpatrick et al., 2017; Li and Hoiem, 2017; Zenke et al., 2017). As an illustration, Elastic weight consolidation (EWC) (Kirkpatrick et al., 2017) is a regularization-based technique that penalizes the changing of parameters with a significant impact on prediction, based on the Fisher Information Matrix (FIM). In this paper, motivated by (Kirkpatrick et al., 2017; Brahma and Rai, 2023), we utilize gradients to approximate FIM as a metric of parameter importance for resetting noisy parameters. Furthermore, we introduce randomness to enhance model robustness during continual adaptation. Additionally, while the continual learning approaches aim to tackle catastrophic forgetting in sequences of new tasks, our work focuses on learning from different domains for a single task.
3. METHOD
3.1. Preliminary
3.1.1. Problem Statement
Table 1. Comparisons between different problem settings. ‘Online’ indicates whether the model has access to the target data or only predicts the incoming test samples immediately.
Setting Source Data Target Training Data Target Distribution Train Loss Test Loss Online
Continual Learning $\times$ $(x^{t},y^{t})$ Dynamic $\mathcal{L}(x^{t},y^{t})$$\times$ $\times$
Unsupervised Domain Adaptation $(x^{s},y^{s})$ $(x^{t})$ Static $\mathcal{L}(x^{s},y^{s})+\mathcal{L}(x^{s},x^{t})$$\times$ $\times$
Source-free Domain Adaptation $\times$ $(x^{t})$ Static $\mathcal{L}(x^{t})$$\times$ $\times$
Test-time Adaptation $\times$ $\times$ Static $\times$ $\mathcal{L}(x^{t})$checkmark
Continual Test-Time Adaptation $\times$ $\times$ Dynamic $\times$ $\mathcal{L}(x^{t})$checkmark
Given a sequence of domain $D=\{D^{0},D^{1},...,D^{n}\}$, we define the first domain $D^{0}$ as the source domain and the subsequent domains as the target domain. The objective of CTTA is to enhance the performance of the model $f_{\theta_{0}}(x)$, where the parameters $\theta_{0}$ are pre-trained on source data $x^{D^{0}},y^{D^{0}}$ from $D^{0}$, in a continually changing target domain during inference time without using source data. For simplicity, we denote the data without the superscript about the specific domain in the following discussion. At time step $t$, unlabeled target data $x_{t}$ is provided sequentially, following the domain group order. The model is required to make a prediction $\hat{y}_{t}=f_{\theta_{t-1}}(x_{t})$ using the parameters $\theta_{t-1}$ which have been updated based on previous target data $x_{1},...,x_{t-1}$. Subsequently, $\hat{y}_{t}$ serves as the evaluation output at time step t, and the model will adapt itself toward $x_{t}$ as $\theta_{t-1}\rightarrow\theta_{t}$, which will only influence future inputs $x_{t+n}$.
Moreover, We compare CTTA with other adaptation settings, as detailed in Table 1. These settings are developed to meet the diverse prerequisites and requirements for real-world applications.
3.1.2. Mean-Teacher Framework
In alignment with previous CTTA works (Wang et al., 2022; Gong et al., 2022; Döbler et al., 2023; Niloy et al., 2024), we build our method based on the mean-teacher framework (Tarvainen and Valpola, 2017). This framework is characterized by the interplay between a teacher model and a student model. Both networks are initialized with the source-trained model at the beginning of adaptation. The teacher model produces the pseudo-labels for the unlabeled target data, which serve as labels of the unlabeled data to supervise the student model. While the parameters of the student model are optimized via gradient descent, the parameters of the teacher model are updated following an Exponential Moving Average (EMA) strategy based on the student model. Formally, this process can be expressed as follows:
(1) $\mathcal{L}_{pl}(x_{t})=\mathcal{L}_{rpn}(x_{t},\hat{y}_{t})+\mathcal{L}_{rcnn}(x_{t},\hat{y}_{t}),$
(2) $\theta_{t}^{std}\leftarrow\theta_{t-1}^{std}+\gamma\frac{\partial(\mathcal{L}_{pl}(x_{t}))}{\partial\theta_{t-1}^{std}},$
(3) $\theta_{t}^{tch}\leftarrow\alpha\theta_{t-1}^{tch}+(1-\alpha)\theta_{t}^{std},$
where $x_{t}$ and $\hat{y}_{t}$ denote the unlabeled target data and corresponding pseudo-labels generated by the teacher network. $\theta^{std}_{t}$ and $\theta^{tch}_{t}$ symbolize the parameters of student and teacher networks at time step $t$, respectively. Moreover, the supervision loss $\mathcal{L}_{pl}$, which consists of the rpn loss $\mathcal{L}_{rpn}$ and rcnn loss $\mathcal{L}_{rcnn}$ in faster-rcnn (Ren et al., 2015), is utilized for pseudo-labeling. Additionally, $\gamma$ represents the student’s learning rate, and the EMA rate is denoted by $\alpha$. Hence, the teacher can be regarded as an ensemble of historical students, providing stable supervision. Despite the effectiveness of the mean-teacher framework under the static distribution shift, this framework encounters error accumulation and catastrophic forgetting in non-stationary environments. Therefore, we design CTAOD to improve the robustness of feature adaptation in the CTTA setting.
Figure 2. The overview of the proposed CTAOD, which follows the mean-teacher framework. The teacher generates proposals and predictions for the weakly augmented image, whereas the student receives the strongly enhanced images. Our proposed CTAOD includes: (1) The Object-level Contrastive Learning (OCL) module compares region of interest features extracted from feature maps of both networks based on teacher proposals for contrastive learning; (2) The dynamic threshold (DT) strategy adjusts category-specific thresholds based on prediction scores; These pseudo-labels are utilized to supervise the student through the threshold. (3) The Data-driven stochastic restoration (DSR) reset the inactive parameter based on the gradient.
3.2. Proposed Method
3.2.1. Overview
As indicated in Figure 2, this subsection presents a concise overview of our proposed CTAOD, consisting of Object-level Contrastive Learning (OCL), Dynamic Threshold (DT), and Data-driven Stochastic Restoration (DSR). Inspired by UDA methods for object detection (Li et al., 2022a; Liu et al., 2021), we adopt the Weak-Strong augmentation to enable the teacher model to generate reliable pseudo-labels without being affected by heavy augmentation. Specifically, the teacher model receives input with weak augmentations while we input student networks with strong enhanced images. The overall loss of the student model for gradient descent is defined as:
(4) $\mathcal{L}_{all}=\mathcal{L}_{pl}(x_{t})+\lambda\mathcal{L}_{cl}(x_{t})+\mu\mathcal{L}_{kl}(x_{t}),$
where $\mathcal{L}_{cl}(x_{t})$ denotes the contrastive learning (CL) loss, to be elaborated in section 3.2.2. The $\mathcal{L}_{kl}$ represents the Kullback-Leibler (KL) Divergence (Kullback and Leibler, 1951) loss, used to quantify the distinction between two probability distributions. The parameters $\lambda$ and $\mu$ are corresponding weights of the losses. The KL loss is defined as:
(5) $\mathcal{L}_{kl}\left(P\parallel Q\right)=\sum_{x\in\mathcal{X}}P(x)\mathrm{~{}}\log\left(\frac{P(x)}{Q(x)}\right),$
where $P(x)$ and $Q(x)$ represent different distribution. We employ the KL divergence loss to encourage the student model to approximate the teacher model closely.
3.2.2. Object-Level Contrastive Learning
In this subsection, we first briefly introduce SimCLR, a widely used CL approach for self-supervised visual representation learning (Chen et al., 2020). SimCLR aims to learn high-quality feature representation across differently augmented views of the same image. For the $i^{th}$ image $x_{i}$ in a batch, the CL loss in SimCLR is formulated as follows:
(6) $\mathcal{L}_{cl}(x_{i})=-\log\left(\frac{\exp(sim(z_{i},z_{j})/\tau)}{\sum_{k=1,\ni_{k\neq i}}^{2N}\exp(sim(z_{i},z_{k})/\tau)}\right),$
where $\tau>0$ is a temperature hyper-parameter. $z_{i}$ and $z_{j}$ denote the features of two different augmentations of the same images, serving as the positive pair. At the same time, $z_{k}$ represents one of the features of a different augmented sample in a batch size of $N$. Here, the $sim(\cdot)$ indicates the similarity function, i.e. cosine similarity. Notably, the SimCLR is initially designed for classification tasks, assuming each image pertains to a single category, and requires large batch sizes to ensure sufficient positive and negative pairs for feature representation learning. Consequently, the original SimCLR is not well-suited for object detection, where images typically contain multiple instances, thus requiring significant computational resources to accommodate large batch sizes.
Motivated by SimCLR (Chen et al., 2020), we present an OCL module to extract teacher and student features for CL based on proposals generated from the region proposal network (RPN). Since proposals provide multiple cropped views around the object instance, this strategy does not require a large batch size, thus computationally efficient for online detection model updates. Specifically, given a weakly augmented image $f_{weak}(x_{t})$ at time step t, the teacher produces ROI proposals $p_{t}=\{p_{1},...,p_{l}\}$ via region proposal network. We then apply RoIAlign (He et al., 2017), a pooling operation for ROI, to extract corresponding teacher and student object-level features $F_{t}^{T}=\{f_{i}^{T}\in\mathbb{R}^{1\times C}\}_{i=1}^{l}$ and $F_{t}^{S}=\{f_{i}^{S}\in\mathbb{R}^{1\times C}\}_{i=1}^{l}$ based on the feature map from the backbone, respectively. Subsequently, the features are projected into the final object-level features with the same encoder using two fully connected layers. For simplicity, we still denote these two final features as $F_{t}^{T}$ and $F_{t}^{S}$. The features associated with the same proposal are considered positive pairs, otherwise negative pairs. we employ the CL loss to these features $F_{t}^{T}$ and $F_{t}^{S}$ derived from weak and strong enhanced images respectively by minimizing:
(7) $\mathcal{L}_{cl}(x_{t})=\frac{1}{l}\sum_{i=1}^{l}-\log\frac{\exp(f_{i}^{T}\cdot f_{i}^{S}/\tau)}{\sum_{j=1}^{l}\exp(f_{i}^{T}\cdot f_{j}^{S}/\tau)},$
where $l$ indicates the number of the features. This strategy encourages the model to learn fine-grained and localized representations on the target domain, without relying on accurate pseudo-labels. Moreover, the OCL is well integrated into the mean-teacher self-training paradigm as a drop-in enhancement for feature adaptation.
3.2.3. Dynamic Threshold
Due to the presence of domain shifts, pseudo-labels inevitably suffer from label noise. Additionally, the common practice is to use a fixed threshold to remove suspected noisy labels. However, model confidence could differ between different categories and domains, thereby hurting the performance using a fixed threshold. Taking inspiration from the threshold strategy (Wang et al., 2024; Zhao et al., 2023), we design the DT strategy to update the category-specific threshold based on the predicted scores dynamically to minimize noise and select a suitable threshold for each category across different domains in test-time. As shown in Figure 2, the teacher network are first to make the prediction $\hat{y}_{t}=f_{\theta_{t-1}}(f_{wek}(x_{t}))$ for the weakly enhanced images $f_{wek}(x_{t})$ The thresholds for each category are initialized with the same value $\delta_{0}$ at the start of the adaptation, which is updated every iteration by:
(8) $\delta_{t}^{c}\leftarrow\beta\cdot\delta_{t-1}^{c}+(1-\beta)\cdot\epsilon\cdot(\overline{l_{t}^{c}})^{\frac{1}{2}},$
where $\delta_{t}^{c}$ denotes the threshold of category $c$ at time step t, $\overline{l_{t}^{c}}$ is the mean predicted scores of the category $c$ at time step $t$, $\beta$ represents the update rate, and $\epsilon$ provides a linear projection. Furthermore, the threshold $\delta_{t}^{c}$ will not change if class $c$ does not exist in prediction $\hat{y}_{t}$, and we set a fixed upper and lower bound $\delta_{max}$ and $\delta_{mini}$. For unlabeled data from dynamic environments, this mechanism effectively prevents the threshold from being too high or too low while generating an appropriate threshold for each category.
3.2.4. Data-Driven Stochastic Restoration
Existing methods address catastrophic forgetting through stochastic reset mechanisms (Wang et al., 2022; Zhu et al., 2023) or pure data-driven restoration (Brahma and Rai, 2023). Nevertheless, the former may reset valuable parameters, potentially erasing essential knowledge relevant to the current domain. On the other hand, the latter restores the most unimportant parameters, which may retain noise parameters, thereby leading to error accumulation. Taking inspiration from (Wang et al., 2022; Brahma and Rai, 2023), we propose the DSR strategy to reset irrelevant parameters with higher possibilities, enabling robust adaptation. Following (Brahma and Rai, 2023), we adopt the diagonal approximation of Fisher Information Matrix (FIM) (Kirkpatrick et al., 2017) based on gradient, as a metric of parameter importance. The FIM approximates parameter importance by considering the sensitivity of the loss function to parameter changes (Kirkpatrick et al., 2017). Parameters with higher values in the FIM are those for which small changes result in significant increases in loss, indicating their importance for domain at hand.
Specifically, consider a convolution layer of the student model with weight after gradient descent $W_{t-1}$ and its corresponding gradient matrix $G_{t}$ with the same shape as $W_{t-1}$ at time step $t$. First, we use a random matrix $R_{t}\sim\operatorname{Uniform}(0,1)$ with the same shape as the weights matrix, where each element follows a uniform distribution between 0 and 1. Moreover, we utilize the element-wise multiplication of the square of the gradient and the random matrix to approximate the FIM (Kirkpatrick et al., 2017):
(9) $F_{t}=G_{t}\odot G_{t}\odot R_{t},$
where $\odot$ represents the element-wise multiplication. Furthermore, the proposed DSR updates the parameters based on the FIM values, which is formulated as:
(10) $M_{t}=F_{t}<\eta,$
(11) $W_{t}=M_{t}\odot W_{0}+(1-M_{t})\odot W_{t-1},$
where $M_{t}$ indicates the mask matrix of the same shape as $W_{t}$, $<$ represents the element-wise less than operation, and $\eta$ is the threshold value which is acquired by the $q$-quantile of $F_{t}$, i.e. $\eta=quantile(F_{t},q)$. Consequently, the elements in $M_{t}$ are set to 1, when the corresponding FIM value is less than $\eta$, indicating the corresponding parameters should be reset to the source weight $W_{0}$. This strategy enables the model to retain essential knowledge while introducing randomness to enhance robustness, thereby mitigating the forgetting issue in CTTA.
4. EXPERIMENTS
Table 2. Experimental results (mAP@0.5) of Cityscapes-to-Cityscapes-C short-term CTTA task. We evaluate the short-term adaptation performance by continually adapting the source-trained model to the twelve corruptions with the largest corruption severity level 5.
Time $t\xrightarrow{\hskip 325.215pt}$ All
Condition Defocus Glass Motion Zoom Snow Frost Fog Brightness Contrast Elastic Pixelate Jpeg Mean Gain
Source (Ren et al., 2015) 6.8 8.1 8.0 1.5 0.2 6.8 34.6 30.7 3.0 50.2 17.6 13.5 15.1 /
Tent (Wang et al., 2020) 6.8 7.8 7.7 1.3 0.2 6.1 33.1 28.0 2.2 51.1 14.8 11.0 14.2 -0.9
CoTTA (Wang et al., 2022) 7.8 9.0 8.9 1.8 0.3 7.1 38.4 31.1 8.6 49.6 16.2 13.1 16.0 +0.9
SVDP (Yang et al., 2024) 7.7 10.1 9.7 2.3 0.7 13.0 42.4 45.2 15.4 47.2 21.2 14.8 19.1 +4.0
IRG (VS et al., 2023a) 8.0 11.0 9.3 3.4 1.2 13.0 37.9 41.3 15.9 38.9 16.9 13.4 17.5 +2.4
MemCLR (VS et al., 2023b) 8.5 10.4 10.6 2.7 1.1 12.2 41.4 41.6 16.4 43.1 15.4 12.7 18.0 +2.9
Ours 8.6 12.1 11.7 3.6 1.5 16.7 44.7 48.1 16.7 47.4 22.5 13.9 20.6 +5.5
Table 3. Experimental results (mAP@0.5) of Cityscapes-to-Cityscapes-C long-term CTTA task. We evaluate the long-term adaptation performance by continually adapting the source model to the five corruption ten times with the largest corruption severity level 5. To save space, we display selected rounds of results, and full results are provided in the supplementary material.
Time $t\xrightarrow{\hskip 390.25534pt}$
Round 1 5 10 All
Condition Fog Motion Snow Brightness Defocus Fog Motion Snow Brightness Defocus Fog Motion Snow Brightness Defocus Mean Gain
Source (Ren et al., 2015) 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 16.4 /
Tent (Wang et al., 2020) 35.8 8.1 0.2 29.2 6.2 30.7 10.2 0.8 27.9 12.4 20.4 5.0 0.1 11.5 2.4 12.0 -4.4
CoTTA (Wang et al., 2022) 38.2 10.6 0.4 33.8 9.3 40.8 10.4 0.6 36.5 9.2 40.8 10.4 0.5 36.3 9.6 19.1 +2.7
SVDP (Yang et al., 2024) 36.8 8.8 0.6 43.5 11.2 45.0 15.4 3.8 47.8 19.9 41.5 16.2 4.6 44.6 20.2 25.3 +8.9
IRG (VS et al., 2023a) 37.9 9.0 0.7 44.3 12.2 45.6 17.7 6.0 46.7 21.9 37.0 16.7 7.1 38.5 20.5 25.6 +9.2
MemCLR (VS et al., 2023b) 37.7 8.9 0.8 45.5 13.5 45.0 18.1 5.1 46.8 22.4 36.8 18.5 7.5 38.4 21.8 26.0 +9.6
Ours 39.0 10.4 0.8 48.0 13.8 49.4 18.6 7.7 51.7 25.0 45.7 20.0 11.8 46.4 25.8 29.0 +12.6
In this study, we rigorously evaluate our methodology across four benchmark tasks tailored for CTTA in object detection. Four tasks encompass continual adaptation to synthetic and real-world distribution shifts, evaluated over short and long-term periods. Inspired by the foundational work CoTTA (Wang et al., 2022), the short-term CTTA task entails the sequential adaptation to various target domains once. In contrast, the long-term CTTA task involves continually adapting the model toward groups of target domains cyclically.
4.1. Datasets
Cityscapes. The Cityscapes dataset (Cordts et al., 2016) is collected for urban scene understanding, encompassing 2,975 training images and 500 validation images with eight object types, i.e. person, rider, car, truck, bus, train, motorcycle, and bicycle. We utilize the model pre-trained on the training set of Cityscapes as the source model, and the source data is discarded during adaptation.
Cityscapes-C. The Cityscapes-C (Hendrycks and Dietterich, 2019) is initially designed to assess robustness against various corruptions, introducing 15 types of corruption with 5 levels of severity. We create the dataset by applying these corruptions at the maximum severity level 5 to the validation set of clean cityscapes, treating each corruption type as an individual target domain comprising 500 images.
For Cityscapes-C, both short-term and long-term CTTA tasks are explored. Our short-term CTTA task selectively focuses on the latter 12 corruptions, i.e. Defocus Blur, Frosted Glass Blur, Motion Blur, Zoom Blur, Snow, Frost, Fog, Brightness, Contrast, Elastic, Pixelate, and JPEG. For the long-term CTTA task, we prioritize the five corruptions with direct relevance to autonomous driving scenarios as the target domain groups following (Gan et al., 2023), namely Fog, Motion, Snow, Brightness, and Defocus. To mimic the scenario in real life where similar environments might be revisited, and to evaluate the forgetting effect of our methods in the long term, we repeat the adaptation to the target domain group 10 times.
SHIFT. The SHIFT (Sun et al., 2022) is a synthetic dataset for autonomous driving, featuring real-world environmental changes. SHIFT can be categorized as clear, cloudy, overcast, rainy, and foggy, where each condition contains images taken at various times ranging from daytime to night. For SHIFT, short-term CTTA tasks are considered. We designate the clear condition as the source domain, with the remaining four conditions as the target domain groups including nearly 20k images in total.
ACDC. The ACDC dataset (Sakaridis et al., 2021), namely the adverse conditions dataset, shares the same class types as Cityscapes and is collected in four different adverse visual conditions, including Fog, Night, Rain, and Snow. Following (Wang et al., 2022; Gan et al., 2023), we use these four conditions as the target domain group for the long-term CTTA task, with 400 unlabeled images from each condition. Similarly, we adapt the source-trained model to continually to the target domain groups 10 cycles for the long-term CTTA task.
4.2. Implementation Details
We adopt the Faster R-CNN (Ren et al., 2015) with ResNet50 (He et al., 2016) pre-trained on ImageNet (Krizhevsky et al., 2012) as the backbone. Aligning with (VS et al., 2023b; Wang et al., 2022), we maintain a batch size of 1, thereby emulating a real-world application scenario where the object detection model is required to adapt toward a continuous influx of images. The source and student models are trained using an SGD optimizer with a learning rate of 0.001 and a momentum of 0.9. Algorithms are implemented leveraging the Detectron2 platform (Wu et al., 2019). The metric of mAP at an IoU threshold of 0.5 (mAP@0.5) is employed for evaluation. Each experiment is conducted on 1 NVIDIA A800 GPU. More implementation details are provided in the supplementary material.
4.3. Baselines and Compared Approaches
To establish the efficacy of our proposed CTAOD, we compare it with five baselines across various setting types for a fair comparison, including Source (Ren et al., 2015), IRG (VS et al., 2023a), Memclr (VS et al., 2023b), Tent (Wang et al., 2020), and CoTTA (Wang et al., 2022). Specifically, ”Source” represents the source model, i.e., Faster R-CNN (Ren et al., 2015), solely pre-trained on the source domain. Moreover, Memclr (VS et al., 2023b) and IRG (Ren et al., 2015) are SFDA object detection methods. MemCLR (VS et al., 2023b) integrates cross-attention with contrastive learning (CL), while IRG (VS et al., 2023a) incorporates instance relation graph and CL. Furthermore, Tent (Wang et al., 2020) updates the affine parameters through entropy minimization in TTA. CoTTA (Wang et al., 2022) utilizes weight-averaged predictions and random recovery neurons to tackle CTTA, while SVDP (Yang et al., 2024) explores sparse visual prompts for CTTA dense prediction.
Table 4. Experimental results (mAP@0.5) of SHIFT short-term CTTA task. We evaluate the adaptation performance by continually adapting the source model to the four conditions.
Time $t\xrightarrow{\hskip 195.12767pt}$ All
Condition Cloudy Overcast Rainy Foggy Mean Gain
Source 51.8 41.5 43.8 33.9 42.7 /
Tent 50.9 39.5 36.3 23.1 37.5 -5.2
CoTTA 51.1 40.1 40.7 29.9 40.5 -2.2
SVDP 52.0 41.0 43.8 35.9 43.2 +0.5
IRG 51.9 40.6 42.7 34.3 42.4 -0.3
MemCLR 51.8 40.4 42.6 35 42.4 -0.3
Ours 52.3 41.3 44.1 36.8 43.6 +0.9
4.4. Results and Analysis
Table 5. Experimental results (mAP@0.5) of Cityscapes-to-ACDC long-term CTTA task. We evaluate the long-term adaptation performance by continually adapting the source model to the four conditions ten times. To save space, we display selected rounds of results, and full results are provided in the supplementary material.
Time $t\xrightarrow{\hskip 368.57964pt}$
Round 1 4 7 10 All
Condition Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow Mean Gain
Source (Ren et al., 2015) 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 36.0 /
Tent (Wang et al., 2020) 52.4 18.6 33.4 38.9 51.7 17.4 31.4 36.0 45.8 14.6 28.1 28.5 35.0 9.5 21.5 18.3 30.5 -5.5
CoTTA (Wang et al., 2022) 53.7 19.7 38.0 42.4 53.1 19.7 37.7 42.9 51.3 19.0 36.5 41.8 50.9 19.1 36.0 42.4 37.8 +1.8
SVDP (Yang et al., 2024) 52.8 20.0 35.6 42.0 54.6 23.5 38.7 43.8 52.8 24.0 38.6 43.9 51.8 23.6 38.2 43.0 39.5 +3.5
IRG (VS et al., 2023a) 52.7 20.6 36.0 42.9 53.6 23.2 38.1 45.6 51.0 23.1 37.1 42.8 49.7 22.5 36.5 40.7 38.9 +2.9
MemCLR (VS et al., 2023b) 52.9 21.2 35.2 42.8 52.8 23.2 38.2 44.4 51.4 23.2 37.1 42.9 49.9 22.9 36.6 41.0 38.7 +2.7
Ours 52.9 20.4 34.5 42.9 54.5 23.8 39.6 45.8 53.5 25.1 39.0 45.3 52.5 24.3 38.8 44.7 40.3 +4.3
4.4.1. Synthetic Continual Distribution Shift
We first evaluate the effectiveness of the proposed method on the synthetic continual shift dataset (i.e. Cityscapes-to-Cityscapes-C adaptation tasks) in the short term. As depicted in Table 2, Tent (Wang et al., 2020) undergoes a slight decline in performance, dropping from 15.1 to 14.2 mAP relative to the source model. This downturn may be attributed to its dependency on a large batch size to update the parameters of the batchnorm layer, making it suboptimal for online adaptation. In contrast, CoTTA (Wang et al., 2022), IRG (VS et al., 2023a), and Memclr (VS et al., 2023b) exhibit enhancements in performance, achieving 16.0, 17.5, and 18.0 mAP respectively. Although the SFDA methods like IRG and MemCLR assume a static target domain, employing a large momentum update rate $\alpha$ for the teacher enables a relatively stable adaptation in the short term. Consequently, CL leads to a better performance than CoTTA which employs weight-averaged predictions. However, SVDP that utilizes visual prompts achieves the sub-optimal performance of 19.1 mAP. Remarkably, our proposed method improves the performance to 20.6 mAP, consistently outperforming the above approaches. CTAOD ensures a more reliable adaptation to target domains characterized by intense and continual changes in the short term with high-quality pseudo-labels and feature representation.
As presented in Table 9, the long-term task outcomes reveal the source model’s relatively poor performance, with an average mAP of 16.4. Tent’s performance is markedly deteriorative across the timeline. Despite the improvement of Memclr and IRG, their performance also begins to decline in the later rounds, failing to maintain stability over long-term adaptation. We believe this is due to the aforementioned methods not accounting for continual distribution shifts, resulting in error accumulation and catastrophic forgetting. Furthermore, CoTTA utilizes a stochastic restoration mechanism to mitigate forgetting, but its randomness may result in losing crucial information, thus limiting its performance at 19.1 mAP. In contrast, SVDP that employ sparse visual prompts raise the performance to 25.3 mAP. Particularly, CTAOD yields a remarkable 12.6 mAP enhancement over the Source and surpasses all comparative baselines, which employs the DSR mechanism to conserve valuable knowledge while eliminating noise from prior domains. These findings empirically validate the effectiveness of our proposed method in ensuring stable adaptation amidst synthetic continual distribution shifts, across both short-term and long-term adaptation scenarios.
(a) Source
(b) MemCLR
(c) CTAOD (ours)
Figure 3. Qualitative results. We compare the detection results of the Source model, MemCLR, and CTAOD in the $10^{th}$ round of adaption to brightness (row 1) and defocus (row 2) corruption on the long-term Cityscapes-to-Cityscapes-C task.
4.4.2. Real-World Continual Distribution Shift
We also evaluate our method on the real-world continual distribution shift dataset (i.e. SHIFT and Cityscapes-to-ACDC adaptation tasks). For the SHIFT short-term CTTA task, where each domain includes more complex images ranging from daytime to night, the results is shown in Table 4. While other baseline method suffer from performance decline, SVDP that employs fixed threshold for pseudo-labeling and prompts learning achieve a 0.5 mAP improvements. Moreover, CTAOD that introduces a DT strategy shows a superior performance of 43.6 mAP.
The experimental results of ACDC long-term CTTA task are summarized in Table 5. Tent, CoTTA, IRG, and MemCLR potentially suffer from error accumulation and catastrophic forgetting, manifesting in a rapid decline in performance during the later stages. Similarly, SVDP achieves a sub-optimal performance of 39.5 mAP. Notably, CTAOD achieves a performance of 40.3 mAP, yielding an absolute improvement of 0.8 mAP over baselines. While CTAOD may not display a significant advantage over other methods in the initial round, its performance continually improves and remains stable throughout the long-term adaptation. This outcome underscores the capability of our method to foster robust adaptation toward the real-world continual distribution shift in the long term.
4.4.3. Qualitative Results
As shown in Figure 3, we provide some examples of detection results. We selectively compare the source model, MemCLR, and CTAOD in the final round adaptation to the brightness and defocus corruption on the long-term Cityscapes-to-Cityscapes-C tasks. Compared with Source and MemCLR, CTAOD guides the model to learn better feature representations while effectively mitigating forgetting in the long-term adaptation. Therefore, CTAOD assists the detector in distinguishing more foreground object categories and better locating them. More visualization analysis is presented in the supplementary material.
Table 6. Alation experiment. “Mean-Teacher” represents the base mean teacher framework with weak-strong augmentation and KL divergence distillation. All experiments are done on long-term Cityscapes-to-Cityscapes-C tasks.
Mean-Teacher OCL DT DSR Mean Gain
1 $\checkmark$ 25.9 /
2 $\checkmark$ $\checkmark$ 27.4 +1.5
3 $\checkmark$ $\checkmark$ 27.1 +1.2
4 $\checkmark$ $\checkmark$ 26.7 +0.8
5 $\checkmark$ $\checkmark$ $\checkmark$ 28.5 +12.1
6 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 29.0 +12.6
4.4.4. Ablation Study
We conduct an ablation study to empirically assess the impact of the proposed OCL, DT, and DSR. As shown in Table 6, the integration of the mean-teacher architecture, which incorporates weak-strong augmentation and KL divergence distillation, elevates the performance to 25.9 mAP. Building upon the mean-teacher foundation, the proposed OCL, DT, and DSR achieve an additional 1.5, 1.2, and 0.8 mAP improvement individually. Moreover, the combination of the OCL and DT modules raises the performance to 28.5 mAP. Furthermore, the inclusion of the DSR mechanism brings a 0.5 mAP boost in performance, culminating in a comprehensive model performance of 29.0 mAP
5. CONCLUSION
In this work, we proposed CTAOD, aimed at addressing the challenges of low-quality pseudo-labels and random errors resulting from stochastic recovery in continual test-tiem adaptation. Firstly, object-level contrastive learning leverages teacher ROI features for contrastive learning to refine the feature representation, tailored for object detection. Secondly, the dynamic threshold strategy selects higher-quality pseudo-labels by dynamically updating the category-specific threshold based on the predicted confidence scores. Lastly, the data-driven stochastic restoration mechanism selectively reset inactive parameters to mitigate forgetting while retaining valuable knowledge. The empirical results of four continual test-time adaptation tasks for object detection demonstrate the efficacy of CTAOD in both short-term and long-term adaption scenarios.
Appendix A Appendix Overview
In this supplementary material, we provide additional details and analysis. In Section B, we provide implementation details on CTAOD and other baseline methodologies. In Section C, we furnish complete results for the quantitative research and additional visualization examples for the qualitative research. In Section D, we discuss the limitations and prospective directions of our research.
Appendix B Additional Implementation Details
Following (VS et al., 2023b), we trained a Faster R-CNN (Ren et al., 2015) with ResNet50 (He et al., 2016) pre-trained on ImageNet (Krizhevsky et al., 2012) on the training set of Cityscapes (Cordts et al., 2016) as the source-trained model. We utilized an SGD optimizer with a learning rate of 0.001, a momentum of 0.9, and a batch size of 1 and 70k iterations. We will specify the implementation details of the CTAOD and other baseline methods in the following subsections.
B.1. Implememtation of CTAOD
Detailed hyper-parameters of CTAOD for each benchmark are listed in Table 8. Furthermore, we present a pseudo-code outline of our CTAOD framework in Algorithm 1. The three key components of CTAOD are highlighted, i.e., Object-level Contrastive Learning, Dynamic Threshold, and Data-driven Stochastic Restoration.
B.2. Implememtation of Baseline Methods
The Tent (Wang et al., 2020), CoTTA (Wang et al., 2022), and SVDP (Yang et al., 2024) are originally designed for the classification or segmentation tasks. We implement them based on their public code, with some modifications to fit into the objection detection. Specifically, the Tent approach is implemented by updating the parameters of the batch normalization in Faster R-CNN through entropy minimization. Meanwhile, we implement the Cotta method by constructing a mean-teacher framework based on Faster R-CNN, which incorporates a stochastic restoration mechanism. Similarly, we implement SVDP with visual prompts on the area of the bounding box in pseudo-labels based on mean-teacher framework. Further, the MemCLR (VS et al., 2023b) and IRG (VS et al., 2023a) are source-free domain adaptation methods for object detection. Therefore, we implement the MemCLR and IRG based on their public code, with little modifications to follow the CTTA setting (i.e., updating the model parameters during evaluation). For a fair comparison, we set the same momentum $\alpha$ of the Exponential Moving Average (EMA) for CoTTA, SVDP, MemCLR, and IRG as CTAOD, all of which apply the mean-teacher framework (Tarvainen and Valpola, 2017). Other hyper-parameters are retained as specified in their public code.
Appendix C Additional Experiment Results
C.1. Quantitative Results
We present the complete experiment results on the two long-term CTTA tasks, i.e., Cityscapes-to-Cityscapes-C and Cityscapes-to-ACDC adaptation tasks, in Table 9 and 10. Experiments show that the proposed CTAOD can largely maintain strong performance in the long term.
C.2. Qualitative Results
As shown in figure 4 - 8, we present more visualization results to compare the CTAOD and other baseline methods based on the final round adaption on the long-term Cityscapes-to-Cityscapes-C task. The results further validate the effectiveness of CTAOD in distinguishing more foreground object categories and better locating them, compared with other baseline methods.
C.3. Additional Ablation Study
Table 7. Additional Ablation Study. The “FT” represents replacing dynamic threshold in CTAOD with a fixed threshold. The “SR” denotes the stochastic restoration (Wang et al., 2022), and the “DR” indicates the data-driven restoration (Brahma and Rai, 2023). All experiments are done on long-term Cityscapes-to-Cityscapes-C tasks.
CTAOD FT of 0.9 FT of 0.8 FT of 0.7 SR DR Mean
1 $\checkmark$ $\checkmark$ 27.7
2 $\checkmark$ $\checkmark$ 28.2
3 $\checkmark$ $\checkmark$ 28.5
4 $\checkmark$ $\checkmark$ 27.1
5 $\checkmark$ $\checkmark$ 28.3
6 $\checkmark$ 29.0
As presented in Table 7, we perform an additional ablation study to evaluate the effectiveness of the proposed Dynamic Threshold (DT) and Data-driven Stochastic Restoration (DSR). The findings indicate that the DT approach results in improvements of 1.3, 0.8, and 0.5 mAP over fixed thresholds set at 0.9, 0.8, and 0.7, respectively. While the stochastic restoration and data-driven restoration limit the performance to 27.1 and 28.3 mAP, the DSR mechanism raises the performance to 29.0 mAP.
Appendix D Limitation
Firstly, our experiments are primarily based on a single detector backbone, i.e. Faster R-CNN (Ren et al., 2015). Future work will extend to other advanced detector backbone networks to assess the universality and generalization capabilities of CTAOD. Moreover, the proposed DSR mechanism to reset the model every iteration increases the computational cost. Future work could explore more efficient mechanisms for restoration, such as restoring the model only when significant changes in the distribution of the target domain are detected. Finally, the evaluation tasks in our work are designed to simulate real-world adaptation scenarios by incorporating datasets affected by corruption and adverse conditions. However, real-world data distributions are inherently more complex. Consequently, a promising avenue for future research would be to apply our methodology in practical, real-world systems to further validate its effectiveness.
Table 8. Detailed hyper-parameters for each benchmark. “City2City-C-Short” denotes the short-term Cityscapes-to-Cityscapes-C task, “City2City-C-Long” denotes long-term Cityscapes-to-Cityscapes-C task, and “ City2ACDC-Long” denotes long-term Cityscapes-to-ACDC task.
Hyper-parameter Description City2City-C-Short City2City-C-Long SHIFT-Short City2ACDC-Long
$\gamma$ Learning rate 0.001 0.001 0.0001 0.001
$\alpha$ Momentum in EMA 0.9996 0.9998 0.9999 0.9998
$\tau$ Temperature in CL 0.07 0.07 0.07 0.07
$\delta_{0}$ Initial valuse in DT 0.8 0.8 0.8 0.8
$\beta$ Hyper-parameter in DT 0.95 0.95 0.95 0.95
$\epsilon$ Hyper-parameter in DT 1.3 1.3 1.3 1.3
$\delta_{max}$ The upper-bound in DT 0.9 0.9 0.9 0.9
$\delta_{mini}$ The lower-bound in DT 0.7 0.7 0.7 0.7
$q$ Quantile value 0.05 0.001 0.001 0.01
Algorithm 1 Pseudo-code for the CTAOD Method
1:Initialization: Object detectors: The student and teacher model parameters $\theta_{0}^{std}$ and $\theta_{0}^{tch}$ initialized from source-trained model $\theta_{0}$. Hyper-parameters: learning rates $\gamma$, Momentum $\alpha$ in exponential moving average (EMA), temperature $\tau$, thresholds $\delta_{0}^{c}$, $\beta$ and $\epsilon$ in dynamic threshold, and quantile value $q$.
2:Input: Unlabeled test data $x_{t}$ for the target domain
3:for each iteration $t$ do
4:       Generate weakly and strongly augmented images $f_{weak}(x_{t})$ and $f_{strong}(x_{t})$
5:       Generate predictions: $y_{t}=f_{\theta_{t-1}^{tch}}(f_{weak}(x_{t}))$ and proposal $p_{t}$ using the teacher model
6:       // 1. Object-level Contrastive Learning
7:       Extract teacher features $F_{t}^{T}$ and student features $F_{t}^{S}$ based on $p_{t}$ using RoIAlign.
8:       Compute contrastive learning loss $\mathcal{L}_{cl}$:
9:             $\mathcal{L}_{cl}(x_{t})=\frac{1}{l}\sum_{i=1}^{l}-\log\frac{\exp(f_{i}^{T}\cdot f_{i}^{S}/\tau)}{\sum_{j=1}^{l}\exp(f_{i}^{T}\cdot f_{j}^{S}/\tau)}$
10:       // 2. Dynamic Threshold
11:       for each category $c$ do
12:             Compute the mean predicted scores $\overline{l_{t}^{c}}$ of category $c$
13:             Update dynamic thresholds based on predictions:
14:                   $\delta_{t}^{c}\leftarrow\beta\cdot\delta_{t-1}^{c}+(1-\beta)\cdot\epsilon\cdot(\overline{l_{t}^{c}})^{\frac{1}{2}}$
15:       end for
16:       Generate the pseudo-label $\hat{y}_{t}$ through the category-specific threshold $\delta_{0}^{c}$
17:       Compute the supervised loss $\mathcal{L}_{pl}(x_{t})=\mathcal{L}_{rpn}(x_{t},\hat{y}_{t})+\mathcal{L}_{rcnn}(x_{t},\hat{y}_{t})$
18:       Perform gradient descent to update the student model $\theta_{t}^{std}$ and get the gradient matrix $G_{t}$
19:       Update teacher model $\theta_{t}^{tch}$ via EMA strategy
20:       // 3. Data-driven Stochastic Restoration
21:       Generate random matrix $R_{t}\sim\operatorname{Uniform}(0,1)$
22:       Generate the fisher information matrix: $F_{t}=G_{t}\odot G_{t}\odot R_{t}$
23:       Find the $q$-quantile of $F_{t}$: $\eta=quantile(F_{t},q)$
24:       Generate the mask matrix: $M_{t}=F_{t}<\eta$
25:       Reset the updated student model: $\theta_{t}^{std}=M_{t}\odot\theta_{0}^{std}+(1-M_{t})\odot\theta_{t}^{std}$
26:end for
27:Output: Teacher predictions $y_{t}$
Table 9. Complete experimental results (mAP@0.5) of Cityscapes-to-Cityscapes-C long-term CTTA task. We evaluate the long-term adaptation performance by continually adapting the source model to the five corruption ten times with the largest corruption severity level 5. “F.”, “M.”, “S.”, “B.”, and “D.” represent Fog, Motion, Snow, Brightness, and Defocus, respectively.
Time $t\xrightarrow{\hskip 403.26341pt}$
Round 1 2 3 4 5 /
Condition F. M. S. B. D. F. M. S. B. D. F. M. S. B. D. F. M. S. B. D. F. M. S. B. D. / /
Source 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 / /
Tent 35.8 8.1 0.2 29.2 6.2 34.6 7.7 0.1 27.5 5.5 33.5 7.8 0.1 25.5 5.1 32 7.3 0.1 23.2 4.6 30.5 6.8 0.1 21.0 3.9 / /
CoTTA 38.2 10.6 0.4 33.8 9.3 39.9 10.7 0.5 34.0 9.9 40.6 10.4 0.6 34.4 9.8 40.9 10.4 0.6 35.4 9.1 40.8 10.4 0.6 36.5 9.2 / /
SVDP 36.8 8.8 0.6 43.5 11.2 43.9 11.8 1.7 47.9 14.9 46.2 13.9 2.5 50.0 17.8 46.2 14.4 3.2 49.0 19.6 45.0 15.4 3.8 47.8 19.9 / /
IRG 37.9 9.0 0.7 44.3 12.2 46.1 12.9 2.0 49.2 16.6 47.6 15.2 3.0 49.5 19.5 45.2 17.3 4.6 48.6 21.4 45.6 17.7 6.0 46.7 21.9 / /
MemCLR 37.7 8.9 0.8 45.5 13.5 45.8 13.6 2.3 49.8 17.7 47.3 16.4 4.1 50.2 20.3 46.4 17.4 4.8 47.8 21.9 45.0 18.1 5.1 46.8 22.4 / /
Ours 39.0 10.4 0.8 48.0 13.8 48.2 13.7 3.1 52.9 18.1 49.3 15.9 5.3 53.3 21.3 48.5 17.1 7.1 52.2 23.5 49.4 18.6 7.7 51.7 25.0 / /
Round 6 7 8 9 10 Mean Gain
Source 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31 6.7 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 16.4 /
Tent 28.6 6.4 0.1 19.0 3.5 26.6 6.3 0.1 16.9 3.0 24.8 5.9 0.1 15.2 2.7 22.9 5.6 0.1 13.7 2.7 20.4 5.0 0.1 11.5 2.4 12.0 -4.4
CoTTA 39.6 10.3 0.6 34.9 9.4 40.2 10.8 0.5 35.3 10.0 40.0 10.8 0.5 35 9.7 40.1 10.5 0.5 34.1 9.5 40.8 10.4 0.5 36.3 9.6 19.1 +2.7
SVDP 44.4 15.6 4.2 47.6 20.5 44.1 16.1 4.3 46.8 20.4 43.0 15.9 4.6 45.7 20.7 42.6 16.0 4.8 45.0 20.2 41.5 16.2 4.6 44.6 20.2 25.3 +8.9
IRG 43.9 18 5.9 44.8 22.1 42.3 17.5 6.6 42.9 22.2 41.7 16.9 7.2 42.1 21.8 38.8 17 7.2 40.1 20.6 37.0 16.7 7.1 38.5 20.5 25.6 +9.2
MemCLR 43.4 19.3 5.9 44.6 23.3 41.1 19.5 6.6 43.1 23.9 39.6 19.1 7.0 40.9 24.3 38.6 18.9 7.2 39.1 22.4 36.8 18.5 7.5 38.4 21.8 26.0 +9.6
Ours 48.5 19.7 8.9 50.7 26.3 47.8 19.6 9.8 49.0 26.1 45.7 19.3 10.5 49.4 26.9 45.2 19.7 11.2 47.4 26.9 45.7 20.0 11.8 46.4 25.8 29.0 +12.6
Table 10. Complete experimental results (mAP@0.5) of Cityscapes-to-ACDC long-term CTTA task. We evaluate the long-term adaptation performance by continually adapting the source model to the four conditions ten times with the largest corruption severity level 5.
Time $t\xrightarrow{\hskip 390.25534pt}$
Round 1 2 3 4 5 /
Condition Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow / /
Source 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 / /
Tent 52.4 18.6 33.4 38.9 52.4 18.1 32.5 37.7 52.2 17.8 32.7 36.7 51.7 17.4 31.4 36 50.4 16.5 31.7 32.9 / /
CoTTA 53.7 19.7 38 42.4 53.6 20.3 37.7 42.2 53.7 20.1 38.1 43.3 53.1 19.7 37.7 42.9 52.6 19.5 37 42.2 / /
SVDP 52.8 20.0 35.6 42.0 53.4 22.0 37.4 43.8 54.0 23.0 38.8 44.0 54.6 23.5 38.7 43.8 53.4 24.0 38.0 44.4 / /
IRG 52.7 20.6 36 42.9 53.2 22.4 37.4 44.6 53.8 23.2 38.1 45.5 53.6 23.2 38.1 45.6 53.5 23.2 38.1 45.3 / /
MemCLR 52.9 21.2 35.2 42.8 52.5 23.1 37.3 44.3 53.9 23.2 38.4 44.8 52.8 23.2 38.2 44.4 51.5 23.5 37.2 43.8 / /
Ours 52.9 20.4 34.5 42.9 53 22.9 38.6 44.6 54.2 24 39.5 45.3 54.5 23.8 39.6 45.8 54.4 24.4 39.2 45.4 / /
Round 6 7 8 9 10 Mean Gain
Source 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 36 /
Tent 49 15.8 29.5 30.6 45.8 14.6 28.1 28.5 42.4 12.6 25.8 24.9 38.9 11 23.6 21.6 35 9.5 21.5 18.3 30.5 -5.5
CoTTA 52.3 19.1 36.3 42 51.3 19 36.5 41.8 51.2 18.9 36.1 41.4 51.5 19.2 36.4 41.3 50.9 19.1 36 42.4 37.8 +1.8
SVDP 53.3 24.0 38.9 43.8 52.8 24.0 38.6 43.9 52.0 23.8 38.7 43.4 51.8 23.8 38.6 43.6 51.8 23.6 38.2 43.0 39.5 +3.6
IRG 52.6 23.5 37.5 43.7 51 23.1 37.1 42.8 50.2 23.1 37 42.2 49.8 23.2 36.5 41.1 49.7 22.5 36.5 40.7 38.9 +2.9
MemCLR 50.2 23.3 36.6 43 51.4 23.2 37.1 42.9 50.5 23.1 36.6 42.6 49.9 22.8 36.8 42 49.9 22.9 36.6 41 38.7 +2.7
Ours 53.9 24.9 39.2 45.1 53.5 25.1 39 45.3 53.3 24.6 39.3 45.3 53.2 24.5 39.3 44.9 52.5 24.3 38.8 44.7 40.3 +4.3
(a) Source
(b) Tent
(c) CoTTA
(d) IRG
(e) MemCLR
(f) CTAOD (ours)
Figure 4. Complete qualitative results. We compare the detection results of the CTAOD and other baseline methods in the $10^{th}$ round of adaption to Brightness corruption on the long-term Cityscapes-to-Cityscapes-C task.
(a) Source
(b) Tent
(c) CoTTA
(d) IRG
(e) MemCLR
(f) CTAOD (ours)
Figure 5. Complete qualitative results. We compare the detection results of the CTAOD and other baseline methods in the $10^{th}$ round of adaption to Defocus corruption on the long-term Cityscapes-to-Cityscapes-C task.
(a) Source
(b) Tent
(c) CoTTA
(d) IRG
(e) MemCLR
(f) CTAOD (ours)
Figure 6. Complete qualitative results. We compare the detection results of the CTAOD and other baseline methods in the $10^{th}$ round of adaption to Fog corruption on the long-term Cityscapes-to-Cityscapes-C task.
(a) Source
(b) Tent
(c) CoTTA
(d) IRG
(e) MemCLR
(f) CTAOD (ours)
Figure 7. Complete qualitative results. We compare the detection results of the CTAOD and other baseline methods in the $10^{th}$ round of adaption to Motion corruption on the long-term Cityscapes-to-Cityscapes-C task.
(a) Source
(b) Tent
(c) CoTTA
(d) IRG
(e) MemCLR
(f) CTAOD (ours)
Figure 8. Complete qualitative results. We compare the detection results of the CTAOD and other baseline methods in the $10^{th}$ round of adaption to Snow corruption on the long-term Cityscapes-to-Cityscapes-C task.

### v3

@@
 Abstract.
-For real-world applications, neural network models are commonly deployed in dynamic environments, where the distribution of the target domain undergoes temporal changes.
-Continual Test-Time Adaptation (CTTA) has recently emerged as a promising technique to gradually adapt a source-trained model to test data drawn from a continually changing target domain.
-Despite recent advancements in addressing CTTA, two critical issues remain: 1) The use of a fixed threshold for pseudo-labeling in existing methodologies leads to the generation of low-quality pseudo-labels, as model confidence varies across categories and domains; 2) While current solutions utilize stochastic parameter restoration to mitigate catastrophic forgetting, their capacity to preserve critical information is undermined by its intrinsic randomness.
-To tackle these challenges, we present CTAOD, aiming to enhance the performance of detection models in CTTA scenarios.
-Inspired by prior CTTA works for effective adaptation, CTAOD is founded on the mean-teacher framework, characterized by three core components.
-Firstly, the object-level contrastive learning module tailored for object detection extracts object-level features using the teacher’s region of interest features and optimizes them through contrastive learning.
-Secondly, the dynamic threshold strategy updates the category-specific threshold based on predicted confidence scores to improve the quality of pseudo-labels.
-Lastly, we design a data-driven stochastic restoration mechanism to selectively reset inactive parameters using the gradients as weights for a random mask matrix, thereby ensuring the retention of essential knowledge.
-We demonstrate the effectiveness of our approach on four CTTA tasks for object detection, where CTAOD outperforms existing methods, especially achieving a 3.0 mAP improvement on the Cityscapes-to-Cityscapes-C CTTA task.
-The code of this work will be released soon.
+Real-world application models are commonly deployed in dynamic environments, where the target domain distribution undergoes temporal changes.
+Continual Test-Time Adaptation (CTTA) has recently emerged as a promising technique to gradually adapt a source-trained model to continually changing target domains.
+Despite recent advancements in addressing CTTA, two critical issues remain: 1) Fixed thresholds for pseudo-labeling in existing methodologies generate low-quality pseudo-labels, as model confidence varies across categories and domains; 2) Stochastic parameter restoration methods for mitigating catastrophic forgetting fail to effectively preserve critical information due to their intrinsic randomness.
+To tackle these challenges for detection models in CTTA scenarios, we present CTAOD, featuring three core components.
+Firstly, the object-level contrastive learning module extracts object-level features for contrastive learning to refine the feature representation in the target domain.
+Secondly, the adaptive monitoring module dynamically skips unnecessary adaptation and updates the category-specific threshold based on predicted confidence scores to enable efficiency and improve the quality of pseudo-labels.
+Lastly, the data-driven stochastic restoration mechanism selectively reset inactive parameters with higher possibilities, ensuring the retention of essential knowledge.
+We demonstrate the effectiveness of CTAOD on four CTTA object detection tasks, where CTAOD outperforms existing methods, especially achieving a 3.2 mAP improvement and a 20% increase in efficiency on the Cityscapes-to-Cityscapes-C CTTA task.
+The code will be released.
 copyright: none
@@
 Figure 1.
-Motivation of our proposed method.
+Motivation.
 The source-trained model adapts to dynamic environments without source data.
-(1) DT: The bar chart above represents the mean model scores for different categories across four domains of the ACDC dataset (Sakaridis et al., 2021).
-While existing methods produce the pseudo-labels via a fixed threshold, CTAOD utilizes DT to update the category-specific threshold based on the mean predicted scores.
-(2) DSR: The bottom describes the neuron reset process, where neurons with deeper shades of purple indicate increased activity levels and blue neurons represent corresponding neurons from the source model.
+(1) AM: The bar chart shows mean model scores varying across categories and domains.
+Unlike existing methods that continually adapt to all test data using fixed thresholds, CTAOD dynamically skips unnecessary adaptations and updates the category-specific threshold.
+(2) DSR: The bottom illustrates the neuron reset process, where darker purple indicates higher activity, and blue represents source knowledge.
 While existing methods restore parameters randomly, CTAOD resets inactive parameters with higher possibilities.
-Deep learning models have demonstrated immense potential across various data modalities such as image (He et al., 2016; Dosovitskiy et al., 2020; Ren et al., 2015), audio (Aytar et al., 2016; Amodei et al., 2016), language (Devlin et al., 2018; Vaswani et al., 2017), and video (Tran et al., 2015; Carreira and Zisserman, 2017).
+Deep learning models have demonstrated immense potential across various data modalities such as image (He et al., 2016; Dosovitskiy et al., 2020), audio (Aytar et al., 2016), and language (Devlin et al., 2018).
 However, these models experience pronounced degradation in performance when confronted with training data (i.e., source domain) and testing data (i.e., target domain) originating from disparate distributions.
-This phenomenon, commonly referred to as distribution shifts, poses a significant challenge (Geirhos et al., 2018; Hendrycks and Dietterich, 2019; Luo et al., 2020; Deng et al., 2014).
-In such a scenario, unsupervised domain adaptation (UDA) becomes crucial, which typically involves aligning the source and target data distributions, thereby mitigating the impact of distribution shifts (Ganin et al., 2016; Tzeng et al., 2017; Saito et al., 2018).
-Still, UDA falls short by necessitating access to source data, which is often inaccessible due to privacy constraints or data transmission barriers (VS et al., 2023b; Huang et al., 2021).
-This limitation catalyzes the exploration of Test-Time Adaptation (TTA) where the source-trained model directly adapts toward unlabeled test samples encountered during evaluation in an online manner, without the reliance on the source data (Mummadi et al., 2021; Sinha et al., 2023; You et al., 2021; Wang et al., 2020).
-Nonetheless, the aforementioned methods, which assume a static target domain, face a more challenging and realistic problem, as real-world machine learning systems work in non-stationary and continually evolving environments (Wang et al., 2022; Döbler et al., 2023).
-Existing TTA methods are vulnerable to catastrophic forgetting of previously learned source knowledge and error accumulation when adaptation faces more than one distribution shift (Wang et al., 2022; Döbler et al., 2023; Niloy et al., 2024).
+This phenomenon, commonly referred to as distribution shifts, poses a significant challenge (Geirhos et al., 2018; Hendrycks and Dietterich, 2019).
+In such a scenario, unsupervised domain adaptation (UDA) becomes crucial, which typically involves aligning the source and target data distributions, thereby mitigating the impact of distribution shifts (Ganin et al., 2016; Tzeng et al., 2017).
+Still, UDA falls short by necessitating access to source data, which is often inaccessible due to privacy constraints or data transmission barriers (Huang et al., 2021).
+This limitation catalyzes the exploration of Test-Time Adaptation (TTA) where the source-trained model directly adapts toward unlabeled test samples encountered during evaluation in an online manner, without the reliance on the source data (Wang et al., 2020).
+Nonetheless, the TTA methods, which assume a static target domain, face a more challenging and realistic problem, as real-world systems work in non-stationary environments.
 For example, a vehicle may encounter various continuous environmental changes such as fog, night, rain, and snow during its journey.
+Existing TTA methods are vulnerable to catastrophic forgetting of source knowledge and error accumulation when adaptation faces more than one distribution shift (Wang et al., 2022).
 Recently, Wang et al.
-(2022) introduce CoTTA by applying stochastic parameters restoration to mitigate catastrophic forgetting in Continual Test-Time Adaptation (CTTA) scenarios (see Table 1), where the model is continually adapted to sequences of target domains.
-Although the strategy of randomly selecting parameters for restoration partially helps alleviate the forgetting of source knowledge (Wang et al., 2022; Zhu et al., 2023), its randomness may also contribute to the loss of crucial knowledge specific to the current domain (refer to Figure 1).
-Additionally, existing CTTA methods normally rely on self-training methods by utilizing pseudo-labels through a fixed threshold to supervise the training of the model (Wang et al., 2022; Döbler et al., 2023; Sójka et al., 2023).
-Given that model confidence could vary across categories and domains (refer to Figure 1), employing a uniformly fixed threshold could result in excluding high-quality pseudo-labels while incorporating incorrect ones.
-Erroneous or missing pseudo-labels lead to error accumulation as they provide negative feedback to the model.
-Therefore, despite recent efforts to handle CTTA, two significant challenges persist: Current self-training-based methods suffer from noisy and low-quality pseudo-labels, leading to error accumulation (Challenge 1); Continual adaptation to dynamic environments in existing methods struggles to effectively retain valuable knowledge about the current domain while mitigating noise specific to the previous domain (Challenge 2).
-To tackle the identified challenges in the CTTA setting for object detection, we introduce CTAOD (Continual Test-time Adaption for Object Detection).
-Aligning with previous CTTA works for robust adaptation (Wang et al., 2022; Gong et al., 2022; Döbler et al., 2023), CTAOD is constructed upon the mean-teacher framework (Tarvainen and Valpola, 2017) where the teacher model is an exponential moving average of the student model.
-Particularly, CTAOD comprises Object-level Contrastive Learning (OCL), Dynamic Threshold (DT), and Data-driven Stochastic Restoration (DSR).
-More specifically, we propose OCL, which is tailored for object detection, to acquire fine-grained and localized feature representation, motivated by the effectiveness of contrastive learning (CL) in self-supervised learning (Chen et al., 2020; He et al., 2020; Grill et al., 2020; Chen and He, 2021).
-OCL extracts features based on proposals generated by the Region Proposal Network (RPN), which provides multiple cropped views around the object instance at different locations and scales.
-Subsequently, CL loss is applied on RPN cropped views to guide the model to encourage similar object instances to remain close while pushing dissimilar ones apart.
-The OCL is well integrated into the mean-teacher paradigm as a drop-in enhancement for feature adaptation.
-Furthermore, to improve the quality of pseudo-labels across various categories, i.e., addressing Challenge 1, we devise a DT strategy to dynamically adjust the threshold for each category individually based on the predicted confidence scores.
-The dynamic nature of the DT method makes it better suited to address the effects of continuously changing distributions compared to fixed thresholds.
-Finally, the DSR mechanism is utilized to reset inactive parameters with a higher possibility than active ones by employing the gradients of parameters as weights for a random mask matrix.
-On the one hand, it not only helps alleviate the impact of noise from the previous domain but also preserves important information, i.e., addressing Challenge 2.
+(2022) introduce CoTTA by applying stochastic parameters restoration to mitigate catastrophic forgetting in Continual Test-Time Adaptation (CTTA) scenarios, where the model is continually adapted to sequences of target domains.
+Although randomly resetting parameters partially helps alleviate the forgetting of source knowledge (Wang et al., 2022; Zhu et al., 2023), its randomness may also contribute to losing crucial knowledge specific to the current domain (refer to Figure 1).
+Additionally, existing CTTA methods normally utilize pseudo-labeling through a fixed threshold for self-supervision (Wang et al., 2022; Döbler et al., 2023).
+Given that model confidence may vary across categories and domains (refer to Figure 1), employing a uniformly fixed threshold could exclude high-quality pseudo-labels while incorporating incorrect ones.
+Furthermore, these methods adapt to every new incoming data, resulting in computational inefficiency and potential performance degradation.
+Low-quality pseudo-labels and unstable adaptation to noisy data lead to error accumulation, as they provide negative feedback to the model.
+Therefore, despite recent efforts to handle CTTA, two significant challenges persist: Current self-training-based methods suffer from low-quality pseudo-labels, leading to error accumulation (Challenge 1); Continual adaptation to dynamic environments struggles to effectively retain valuable knowledge about the current domain while mitigating the forgetting of source knowledge (Challenge 2).
+To tackle these challenges, we introduce CTAOD (Continual Test-time Adaption for Object Detection).
+Aligning with previous CTTA works for robust adaptation (Wang et al., 2022), CTAOD is constructed upon the mean-teacher framework (Tarvainen and Valpola, 2017) where the teacher model is an exponential moving average of the student model.
+Particularly, CTAOD comprises Object-level Contrastive Learning (OCL), Adaptive Monitoring (AM), and Data-driven Stochastic Restoration (DSR).
+More specifically, OCL extracts features based on proposals generated by the Region Proposal Network (RPN), which provides multiple cropped views around the object at different locations and scales.
+Subsequently, Contrastive Learning (CL) loss is applied on RPN cropped views to guide the model to encourage similar object instances to remain close while pushing dissimilar ones apart.
+The OCL is well integrated into the mean-teacher paradigm as a drop-in enhancement to acquire more fine-grained and localized feature representation.
+Furthermore, to make adaptation more efficient and stable and to improve the quality of pseudo-labels, i.e., addressing Challenge 1, we design an AM module to decide whether to pause or resume the adaptation and dynamically adjust the category-specific thresholds, based on the predicted confidence scores.
+The dynamic nature of the AM method makes it better suited to address the effects of continuously changing distributions.
+Finally, the DSR mechanism is proposed to reset inactive parameters with a higher possibility than active ones by employing the gradients of parameters as weights for a random mask matrix.
+DSR not only helps prevent forgetting but preserves important information, i.e., addressing Challenge 2.
 On the other hand, its randomness also allows potentially falsely activated parameters to be reset, thereby making the model more stable.
 We demonstrate the effectiveness of our proposed approach on four continual test-time adaptation benchmark tasks for object detection.
-These datasets are associated with synthetic and real-world distribution shifts in the short-term and long-term adaption (i.e.
-Cityscapes (Cordts et al., 2016) $\rightarrow$ Cityscapes-C (Hendrycks and Dietterich, 2019), SHIFT (Sun et al., 2022), and Cityscapes (Cordts et al., 2016) $\rightarrow$ ACDC (Sakaridis et al., 2021) adaptation).
-Experimental results indicate that our method significantly improves performance over existing state-of-the-art methods, with gains of up to 3.0 mAP on these tasks.
-In summary, the main contributions of this paper are as follows:
-  • This study introduces a novel method named CTAOD, which pioneers exploring CTTA for detection models.
-Specifically, we propose to leverage object-level features for contrastive learning to refine feature representation for object detection.
-  • To address two challenges in CTTA, we dynamically update the category-specific threshold based on predicted scores to improve the quality of pseudo-labels and reset the inactive parameter with higher possibilities to mitigate forgetting.
-  • Empirical experiments demonstrate that our proposed method surpasses existing methods and effectively facilitates short-term and long-term adaptation in dynamic environments.
+These datasets are associated with synthetic and real-world distribution shifts in the short-term and long-term adaption, i.e.
+Cityscapes (Cordts et al., 2016), Cityscapes-C (Hendrycks and Dietterich, 2019), SHIFT (Sun et al., 2022), and ACDC (Sakaridis et al., 2021).
+Experimental results indicate that our method significantly improves performance over existing methods, with gains of up to 3.2 mAP and 20% in computational efficiency.
+Our main contributions are summarized as follows:
+  • This study introduces CTAOD, which pioneers exploring CTTA for detection models.
+Specifically, we propose to leverage object-level features for contrastive learning to refine feature representation in CTTA object detection.
+  • To address the two challenges in CTTA, our proposed CTAOD enables dynamic skipping and category-specific threshold updates based on predicted scores and reset the inactive parameter with higher possibilities.
+  • Empirical experiments demonstrate that CTAOD surpasses existing methods and facilitates efficient short-term and long-term adaptation in dynamic environments.
 2. RELATED WORK
 2.1. Source-Free Domain Adaptation
-UDA tackles the inter-domain divergence by aligning the distributions of source and target data (Ganin et al., 2016; Li et al., 2019; Luo et al., 2020; Jiang et al., 2020).
-Despite its effectiveness, the limitation of UDA lies in its requirement for access to the source domain data, which often raises concerns regarding data privacy, data portability, and data transmission efficiency.
-As a result, Source-Free Domain Adaptation (SFDA) received extensive research attention, especially in object detection, where the source-trained detector is re-trained on the target training data (VS et al., 2023b, a; Li et al., 2022b, 2021a, 2021b) before evaluation.
-For instance, MemCLR (VS et al., 2023a) employs a cross-attention transformer-based memory bank for contrastive learning, while IRG (VS et al., 2023a) utilizes the object relations with instance relation graph network to explore the SFDA setting for object detection without the source data.
+UDA tackles the inter-domain divergence by aligning the distributions of source and target data (Ganin et al., 2016; Luo et al., 2020).
+Despite its effectiveness, the limitation of UDA lies in its requirement for access to the source domain data, which often raises concerns regarding data privacy and transmission efficiency.
+As a result, Source-Free Domain Adaptation (SFDA) received extensive research attention in object detection, where the source-trained detector is re-trained on the target training data (VS et al., 2023b, a; Li et al., 2021) before evaluation.
+For instance, MemCLR (VS et al., 2023a) employs a cross-attention-based memory bank for CL, while IRG (VS et al., 2023a) utilizes the object relations with instance relation graph network to explore the SFDA setting for object detection without the source data.
 However, the standard SFDA setting requires prior knowledge of the target domain, which is impractical in most real-world applications.
@@
 In this paper, we distinguish between TTA and SFDA based on evaluation protocol, although they can be transformed into each other in experiments.
-Furthermore, TTA methods improve the model performance under distribution shift commonly through techniques like pseudo-labeling (Sinha et al., 2023; Iwasawa and Matsuo, 2021; Sun et al., 2020; Zeng et al., 2023), batchnorm statistics updating (Hu et al., 2021; You et al., 2021), or entropy regularization (Wang et al., 2020; Iwasawa and Matsuo, 2021; Niu et al., 2022) during testing.
+Furthermore, TTA methods improve the model performance under distribution shift commonly through pseudo-labeling (Sinha et al., 2023; Iwasawa and Matsuo, 2021), batchnorm statistics updating (Hu et al., 2021; You et al., 2021), or entropy regularization (Wang et al., 2020; Iwasawa and Matsuo, 2021) during testing.
 For example, Tent (Wang et al., 2020) updates the batchnorm parameters with entropy minimization and demands a large batch size for optimization during test-time adaptation, which is unsuitable for real-time detection model deployment where images are processed sequentially.
-The above approaches assume a specific and static target domain where all the test inputs come from a single domain.
-However, in numerous practical scenarios, the distribution of test input may exhibit a continual shift over time.
+TTA methods assume a static target domain where the target data come from a single domain.
+However, the target distribution may exhibit a continual shift over time in practical scenarios.
 2.3. Continual Test-time Adaptation
-Early works consider adaptation to evolving and continually changing domains by aligning the source and target data (Hoffman et al., 2014; Wulfmeier et al., 2018).
-These methods rely on source data during inference, which limits their applicability.
-A significant advancement in this area is the proposal of CoTTA (Wang et al., 2022), marking the first work tailored to the demands of CTTA by adapting a pre-trained model to sequences of domains without using any source data.
-Subsequently, research efforts have been dedicated to exploring CTTA, primarily focused on classification (Döbler et al., 2023; Niloy et al., 2024; Brahma and Rai, 2023; Yu et al., 2023) and segmentation (Niloy et al., 2024; Song et al., 2023; Zhu et al., 2023) tasks.
-For instance, similar to CoTTA (Wang et al., 2022), Zhu et al.
-(2023) apply a stochastic reset mechanism for CTTA in the segmentation task of medical images to prevent forgetting.
+Recently, Wang et al.
+(2022) introduce CoTTA, marking the first work tailored to the demands of CTTA by adapting a pre-trained model to sequences of domains without source data.
+Subsequently, research efforts have been dedicated to exploring CTTA, primarily focused on classification (Döbler et al., 2023; Brahma and Rai, 2023; gan2023decorate) and segmentation (Song et al., 2023; Ni et al., 2023; Liu et al., 2023) tasks.
+For instance, similar to CoTTA, Zhu et al.
+(2023) apply a stochastic reset mechanism to prevent forgetting in the CTTA medical segmentation task.
 In contrast, PETAL (Brahma and Rai, 2023) utilizes the Fisher Information Matrix (FIM) as a metric of parameter importance to reset only the most irrelevant parameters.
 Nevertheless, these two methods are sub-optimal, as the randomness might lead to losing essential information, and pure data-driven restoration may retain false active parameters resulting from noise.
-Moreover, the mean-teacher framework (Tarvainen and Valpola, 2017) serves as a base architecture for most CTTA works (Wang et al., 2022; Gong et al., 2022; Döbler et al., 2023; Niloy et al., 2024), where the teacher commonly generates pseudo-labels via a fixed threshold to supervise the student.
+Moreover, the mean-teacher framework (Tarvainen and Valpola, 2017) serves as a base architecture for most CTTA works (Wang et al., 2022; Gong et al., 2022), where the teacher generates pseudo-labels via a fixed threshold to supervise the student.
 Nonetheless, these methods suffer from low-quality pseudo-labels with a uniform threshold since the model confidence varies across categories and domains.
 Furthermore, Wang et al.
-(2024) design a dynamic thresholding technique to update the threshold based on the maximum predicted score in a batch for classification tasks, which requires a relatively large batch size to set an appropriate threshold.
-However, it is not suitable for object detection, where a smaller batch size is often preferred for computational efficiency.
-Alternatively, Gan et al.
-(2023) presents a cloud-device collaborative continual adaptation paradigm to accommodate dynamic environment in object detection.
-Nevertheless, its dependency on source data also constrains its practical utility.
-Therefore, a gap still exists in exploring the CTTA setting in object detection to improve the quality of pseudo-labels (Challenge 1) and to effectively reset noisy neurons (Challenge 2), without reliance on the source data.
-2.4. Continual Learning
+(2024) design a dynamic thresholding technique to update the threshold in a batch for classification tasks, requiring a large batch size.
+However, this approach is unsuitable for object detection where smaller batch sizes are preferred for computational efficiency.
+To improve efficiency, yoo2024and presents a CTTA approach for object detection to update specific parts of the model in specific adaptation, which utilizes source statistics.
+Although it effectively improved efficiency, its performance is limited.
+Additionally, Gan et al.
+(2023) presents a cloud-device collaborative continual adaptation paradigm in object detection by aligning source and target distribution.
+Nevertheless, the reliance on source data in both methods restricts their practical applicability.
+Therefore, a gap still exists in exploring the CTTA in object detection to improve the quality of pseudo-label (Challenge 1) and effectively reset noisy neurons (Challenge 2), without relying on any source data.
+3. METHOD
+3.1. Preliminary
+3.1.1. Problem Statement
+Given a sequence of domain $D=\{D^{0},D^{1},...,D^{n}\}$, we define $D^{0}$ as the source domain and the subsequent domains as the target domain.
+The objective of CTTA is to enhance the performance of the model $f_{\theta_{0}}(x)$, where the parameters $\theta_{0}$ are pre-trained on source data $x^{D^{0}},y^{D^{0}}$ from $D^{0}$, in a continually changing target domain during inference time without using source data.
+For simplicity, we denote the data without the superscript about the specific domain in the following discussion.
+At time step $t$, unlabeled target data $x_{t}$ is provided sequentially, following the domain group order.
+The model is required to make a prediction $\hat{y}_{t}=f_{\theta_{t-1}}(x_{t})$ using the parameters $\theta_{t-1}$ which have been updated based on previous target data $x_{1},...,x_{t-1}$.
+Subsequently, $\hat{y}_{t}$ serves as the evaluation output at time step t, and the model will adapt itself toward $x_{t}$ as $\theta_{t-1}\rightarrow\theta_{t}$, which will only influence future inputs $x_{t+n}$.
+Figure 2.
+Overview.
+CTAOD follows the mean-teacher framework, featuring: (1) The OCL module compares region of interest features extracted from feature maps of both networks based on teacher proposals for contrastive learning; (2) The AM module dynamically skips unnecessary adaptation and adjusts category-specific thresholds based on prediction scores; (3) The DSR reset the inactive parameter with higher possibilities based on the FIM.
+3.1.2. Mean-Teacher Framework
+In alignment with previous CTTA works (Wang et al., 2022; Gong et al., 2022), we build our method based on the mean-teacher framework (Tarvainen and Valpola, 2017).
+Specifically, both networks are first initialized with the source-trained model.
+The teacher model produces the pseudo-labels for the unlabeled target data to supervise the student model.
+While the parameters of the student model are optimized via gradient descent, the teacher model is updated following an Exponential Moving Average (EMA) strategy based on the student model.
+Formally, this process can be expressed as follows:
+(1) $\mathcal{L}_{pl}(x_{t})=\mathcal{L}_{rpn}(x_{t},\hat{y}_{t})+\mathcal{L}_{rcnn}(x_{t},\hat{y}_{t}),$
+(2) $\theta_{t}^{std}\leftarrow\theta_{t-1}^{std}+\gamma\frac{\partial(\mathcal{L}_{pl}(x_{t}))}{\partial\theta_{t-1}^{std}},$
+(3) $\theta_{t}^{tch}\leftarrow\alpha\theta_{t-1}^{tch}+(1-\alpha)\theta_{t}^{std},$
+where $x_{t}$ and $\hat{y}_{t}$ denote the unlabeled target data and corresponding pseudo-labels.
+$\theta^{std}_{t}$ and $\theta^{tch}_{t}$ symbolize the parameters of student and teacher networks at time step $t$.
+Moreover, the supervision loss $\mathcal{L}_{pl}$ in Faster-RCNN (Ren et al., 2015), which consists of the RPN loss $\mathcal{L}_{rpn}$ and RCNN loss $\mathcal{L}_{rcnn}$, is utilized for pseudo-labeling.
+Additionally, $\gamma$ represents the student’s learning rate, and the EMA rate is denoted by $\alpha$.
+Hence, the teacher can be regarded as an ensemble of historical students, providing stable supervision.
+However, this framework encounters error accumulation and catastrophic forgetting in dynamic environments.
+Therefore, we design CTAOD to fit in the CTTA setting.
+3.2. Proposed Method
+3.2.1. Overview
+Figure 2 presents a concise overview of CTAOD, consisting of Object-level Contrastive Learning (OCL), Adaptive Monitoring (AM), and Data-driven Stochastic Restoration (DSR).
+Inspired by UDA object detection methods (Li et al., 2022; Liu et al., 2021), we adopt the Weak-Strong augmentation to enable the teacher model to generate reliable pseudo-labels without being affected by heavy augmentation.
+Specifically, the teacher model receives input with weak augmentations while we input student networks with strong enhanced images.
+The overall loss of the student model for gradient descent is defined as:
+(4) $\mathcal{L}_{all}=\mathcal{L}_{pl}(x_{t})+\lambda\mathcal{L}_{cl}(x_{t})+\mu\mathcal{L}_{kl}(x_{t}),$
+where $\mathcal{L}_{cl}(x_{t})$ denotes the contrastive learning (CL) loss, to be elaborated later.
+The $\mathcal{L}_{kl}$ represents the Kullback-Leibler (KL) Divergence (Kullback and Leibler, 1951) loss, used to quantify the distinction between two probability distributions.
+The parameters $\lambda$ and $\mu$ are corresponding weights of the losses.
+The KL loss is defined as:
+(5) $\mathcal{L}_{kl}\left(P\parallel Q\right)=\sum_{x\in\mathcal{X}}P(x)\mathrm{~{}}\log\left(\frac{P(x)}{Q(x)}\right),$
+where $P(x)$ and $Q(x)$ represent different distribution. We employ the KL divergence loss to encourage the student model to approximate the teacher model closely.
+3.2.2. Object-Level Contrastive Learning
+SimCLR is a widely used CL approach for self-supervised visual representation learning (Chen et al., 2020).
+SimCLR learns high-quality feature representations from differently augmented views of the same image.
+Designed for classification tasks, it requires large batch sizes to ensure sufficient positive and negative pairs.
+Consequently, SimCLR is not well-suited for object detection, where images contain multiple instances and large batch sizes demand significant computational resources.
+Motivated by SimCLR (Chen et al., 2020), we present an OCL module to extract teacher and student features for CL based on proposals generated from the RPN.
+This strategy provides multiple cropped views around the object instance, eliminating the need for large batch sizes and ensuring computational efficiency for online detection model updates.
+Specifically, given a weakly augmented image $f_{weak}(x_{t})$ at time step t, the teacher produces ROI proposals $p_{t}=\{p_{1},...,p_{l}\}$ via region proposal network.
+We then apply RoIAlign (He et al., 2017), a pooling operation for ROI, to extract corresponding teacher and student object-level features $F_{t}^{T}=\{f_{i}^{T}\in\mathbb{R}^{1\times C}\}_{i=1}^{l}$ and $F_{t}^{S}=\{f_{i}^{S}\in\mathbb{R}^{1\times C}\}_{i=1}^{l}$ based on the feature map from the backbone, respectively.
+The features associated with the same proposal are considered positive pairs, otherwise negative pairs.
+we employ the CL loss to these features $F_{t}^{T}$ and $F_{t}^{S}$ derived from weak and strong enhanced images respectively by minimizing:
+(6) $\mathcal{L}_{cl}(x_{t})=\frac{1}{l}\sum_{i=1}^{l}-\log\frac{\exp(f_{i}^{T}\cdot f_{i}^{S}/\tau)}{\sum_{j=1}^{l}\exp(f_{i}^{T}\cdot f_{j}^{S}/\tau)},$
+where $\tau>0$ is the temperature, $l$ indicates the number of the features, and $f_{i}^{T}$ and $f_{i}^{S}$ denote the features of two different augmentations of the same object, serving as the positive pair.
+This strategy encourages the model to learn fine-grained and localized representations on the target domain, without relying on accurate pseudo-labels.
+3.2.3. Adaptive Monitoring
+The AM module monitors the model’s status using the predicted scores, featuring dynamic skipping and dynamic threshold.
+The former halts unnecessary adaptations to save computational resources, while the latter updates the category-specific threshold.
+As shown in Figure 2, the teacher network first make predictions $\hat{y}_{t}=f_{\theta_{t-1}}(f_{wek}(x_{t}))$ for the weakly enhanced images.
+Subsequently, We compute the mean scores of all category $\overline{l}_{t}$ in teacher’s predictions and maintain its exponentially moving average $\overline{l}_{t}^{ema}$ through:
+(7) $\overline{l}_{t}^{ema}\leftarrow\beta_{s}\cdot\overline{l}_{t-1}^{ema}+(1-\beta_{s})\cdot\overline{l}_{t},$
+where $\beta_{s}$ represents the update rate.
+We introduce a new index $\frac{\overline{l}_{t}}{\overline{l}_{t}^{ema}}$.
+Adaptation resumes when the $\frac{\overline{l}_{t}}{\overline{l}_{t}^{ema}}$ lies within the range $(\frac{1}{\delta_{s}},\delta_{s})$, indicating stability, and pauses when it deviates beyond this range, suggesting changes in data characteristics.
+This dynamic skipping strategy effectively prevents unstable and noisy adaptation, thus conserving computational resources and mitigating error accumulation.
+For dynamic threshold, the thresholds for each category are initialized with the same value $\delta_{0}$ at first, which is updated every iteration by:
+(8) $\delta_{t}^{c}\leftarrow\beta_{t}\cdot\delta_{t-1}^{c}+(1-\beta_{t})\cdot\epsilon\cdot(\overline{l}_{t}^{c})^{\frac{1}{2}},$
+where $\delta_{t}^{c}$ denotes the threshold of category $c$ at time step t, $\overline{l}_{t}^{c}$ is the mean predicted scores of the category $c$ at time step $t$, $\beta_{t}$ represents the update rate, and $\epsilon$ provides a linear projection.
+Furthermore, the threshold $\delta_{t}^{c}$ will not change if class $c$ does not exist in prediction $\hat{y}_{t}$, and we set a fixed upper and lower bound $\delta_{max}$ and $\delta_{mini}$.
+For unlabeled data from dynamic environments, this mechanism effectively generates an appropriate threshold for each category.
+3.2.4. Data-Driven Stochastic Restoration
+Existing methods address catastrophic forgetting through stochastic reset mechanisms (Wang et al., 2022) or pure data-driven restoration (Brahma and Rai, 2023).
+Nevertheless, the former may reset valuable parameters, potentially erasing essential knowledge relevant to the current domain.
+On the other hand, the latter restores the most unimportant parameters, which may retain noise parameters, thereby leading to error accumulation.
+Therefore, we propose the DSR strategy to reset irrelevant parameters with higher possibilities, enabling robust adaptation.
+Following (Brahma and Rai, 2023), we adopt the diagonal approximation of Fisher Information Matrix (FIM) (Kirkpatrick et al., 2017) based on gradient, as a metric of parameter importance.
+The FIM approximates parameter importance by considering the sensitivity of the loss function to parameter changes (Kirkpatrick et al., 2017).
+Parameters with higher values in the FIM are those for which small changes result in significant increases in loss, indicating their importance for the domain at hand.
+Specifically, consider a convolution layer with weight $W_{t-1}$ and its corresponding gradient matrix after gradient descent $G_{t}$ with the same shape as $W_{t-1}$ at time step $t$.
+First, we use a random matrix $R_{t}\sim\operatorname{Uniform}(0,1)$ with the same shape as the weights matrix, where each element follows a uniform distribution between 0 and 1.
+Moreover, we utilize the element-wise multiplication of the square of the gradient and the random matrix to approximate the FIM: $F_{t}=G_{t}\odot G_{t}\odot R_{t}$, where $\odot$ represents the element-wise multiplication.
+Furthermore, the proposed DSR updates the parameters based on the FIM values, which is formulated as:
+(9) $M_{t}=F_{t}<\eta,$
+(10) $W_{t}=M_{t}\odot W_{0}+(1-M_{t})\odot W_{t-1},$
+where $M_{t}$ indicates the mask matrix of the same shape as $W_{t}$, $<$ represents the element-wise less than operation, and $\eta$ is the threshold value which is acquired by the $q$-quantile of $F_{t}$, i.e.
+$\eta=quantile(F_{t},q)$.
+Consequently, the elements in $M_{t}$ are set to 1, when the corresponding FIM value is less than $\eta$, indicating the corresponding parameters should be reset to the source weight $W_{0}$.
+This strategy enables the model to retain essential knowledge while introducing randomness to enhance robustness to mitigate forgetting.
+4. EXPERIMENTS
+Time $t\xrightarrow{\hskip 368.57964pt}$ All
+Condition Defocus Glass Motion Zoom Snow Frost Fog Brightness Contrast Elastic Pixelate Jpeg Mean Gain Iter.
+Source 6.8 8.1 8.0 1.5 0.2 6.8 34.6 30.7 3.0 50.2 17.6 13.5 15.1 / /
+CoTTA 7.8 9.0 8.9 1.8 0.3 7.1 38.4 31.1 8.6 49.6 16.2 13.1 16.0 +0.9 6k
+SVDP 7.7 10.1 9.7 2.3 0.7 13.0 42.4 45.2 15.4 47.2 21.2 14.8 19.1 +4.0 6k
+IRG 8.0 11.0 9.3 3.4 1.2 13.0 37.9 41.3 15.9 38.9 16.9 13.4 17.5 +2.4 6k
+MemCLR 8.5 10.4 10.6 2.7 1.1 12.2 41.4 41.6 16.4 43.1 15.4 12.7 18.0 +2.9 6k
+Ours-unstop 8.6 12.1 11.7 3.6 1.5 16.7 44.7 48.1 16.7 47.4 22.5 13.9 20.6 +5.5 6k
+Ours 8.7 11.7 12.4 3.6 1.5 13.5 43.2 47.3 18.4 47.0 24.8 17.2 20.8 +5.7 4.7k
+Table 1.
+Experimental results (mAP@0.5) and adaptation iterations of Cityscapes-to-Cityscapes-C short-term CTTA task.
+We evaluate the performance by continually adapting the source model to the twelve corruptions.
+“-unstop” denotes no skipping.
+Time $t\xrightarrow{\hskip 368.57964pt}$
+Round 1 5 10 All
+Condition Fog Motion Snow Brightness Defocus Fog Motion Snow Brightness Defocus Fog Motion Snow Brightness Defocus Mean Gain Iter.
+Source 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 16.4 / /
+CoTTA 38.2 10.6 0.4 33.8 9.3 40.8 10.4 0.6 36.5 9.2 40.8 10.4 0.5 36.3 9.6 19.1 +2.7 25k
+SVDP 36.8 8.8 0.6 43.5 11.2 45.0 15.4 3.8 47.8 19.9 41.5 16.2 4.6 44.6 20.2 25.3 +8.9 25k
+IRG 37.9 9.0 0.7 44.3 12.2 45.6 17.7 6.0 46.7 21.9 37.0 16.7 7.1 38.5 20.5 25.6 +9.2 25k
+MemCLR 37.7 8.9 0.8 45.5 13.5 45.0 18.1 5.1 46.8 22.4 36.8 18.5 7.5 38.4 21.8 26.0 +9.6 25k
+Ours-unstop 39.0 10.4 0.8 48.0 13.8 49.4 18.6 7.7 51.7 25.0 45.7 20.0 11.8 46.4 25.8 29.0 +12.6 25k
+Ours 39.2 9.9 0.8 46.7 13.8 49.4 18.5 8.0 52.9 24.7 46.2 19.5 11.3 47.3 29.1 29.2 +12.8 20.1k
+Table 2.
+Experimental results (mAP@0.5) and adaptation iterations of Cityscapes-to-Cityscapes-C long-term CTTA task.
+We evaluate the performance by continually adapting the source model to the five corruption ten times.
+To save space, we display selected rounds of results, and full results are provided in the appendix.
+“-unstop” denotes no skipping.
+In this study, we rigorously evaluate our methodology across four benchmark tasks in CTTA object detection.
+These tasks encompass continual adaptation to synthetic and real-world distribution shifts, evaluated over short and long-term periods.
+Inspired by the foundational work CoTTA (Wang et al., 2022), the short-term CTTA task entails the sequential adaptation to various target domains once.
+In contrast, the long-term CTTA task involves continually adapting the model toward groups of target domains cyclically.
+4.1. Datasets
+Cityscapes.
+The Cityscapes dataset (Cordts et al., 2016) is collected for urban scene understanding, encompassing 2,975 training images and 500 validation images with eight object types.
+We utilize the model pre-trained on the training set of Cityscapes as the source model, and the source data is discarded during adaptation.
+Cityscapes-C.
+The Cityscapes-C (Hendrycks and Dietterich, 2019) is initially designed to assess robustness against various corruptions, introducing 15 types of corruption with 5 levels of severity.
+We create the dataset by applying these corruptions at the maximum severity level 5 to the validation set of clean cityscapes, treating each corruption type as an individual target domain comprising 500 images.
+Our short-term CTTA task selectively focuses on the latter 12 corruptions.
+For the long-term CTTA task, we prioritize the five corruptions related to autonomous driving scenarios as the target domain groups following (Gan et al., 2023), namely Fog, Motion, Snow, Brightness, and Defocus.
+We repeat the adaptation to the target domain group 10 times to evaluate long-term performance.
+SHIFT.
+The SHIFT (Sun et al., 2022) is a synthetic dataset for autonomous driving, featuring real-world environmental changes.
+SHIFT can be categorized as clear, cloudy, overcast, rainy, and foggy, where each condition contains images taken at various times ranging from daytime to night.
+For SHIFT, short-term CTTA tasks are considered.
+We designate the clear condition as the source domain, with the remaining four conditions as the target domain groups including nearly 20k images in total.
+ACDC.
+The ACDC dataset (Sakaridis et al., 2021) shares the same class types as Cityscapes and is collected in four different adverse visual conditions, including Fog, Night, Rain, and Snow.
+Following (Wang et al., 2022), we use these four conditions as the target domain group for the long-term CTTA task, with 400 unlabeled images per condition.
+Similarly, the source-trained model is continually adapted to the target domain group for 10 cycles.
+4.2. Implementation Details
+We adopt the Faster R-CNN with ResNet50 (He et al., 2016) pre-trained on ImageNet (Krizhevsky et al., 2012) as the backbone.
+Aligning with (VS et al., 2023b; Wang et al., 2022), we maintain a batch size of 1, thereby emulating a real-world application scenario where the detection model adapts toward a continuous influx of images.
+The source models are trained using an SGD optimizer with a learning rate of 0.001 and a momentum of 0.9.
+Algorithms are implemented leveraging the Detectron2 platform (Wu et al., 2019).
+The metric of mAP at an IoU threshold of 0.5 is employed for evaluation.
+Each experiment is conducted on 1 NVIDIA A800 GPU.
+More hyper-parameters and implementation details are provided in the appendix.
+4.3. Baselines and Compared Approaches
+We compare CTAOD with six baselines across various settings, including Source (Ren et al., 2015), IRG (VS et al., 2023a), Memclr (VS et al., 2023b), CoTTA (Wang et al., 2022), SVDP (Yang et al., 2024), and CTAOD-upstop.
+Specifically, ”Source” represents the source model without adaptation.
+Memclr and IRG are SFDA object detection methods, also referred to as TTA methods (Wang et al., 2022).
+MemCLR integrates cross-attention with CL, while IRG incorporates instance relation graph and CL.
+CoTTA utilizes weight-averaged predictions and random neurons recovery to tackle CTTA, while SVDP explores sparse visual prompts for CTTA dense prediction.
+“CTAOD-unstop” represents CTAOD without dynamic skipping.
+4.4. Results and Analysis
+Time $t\xrightarrow{\hskip 368.57964pt}$
+Round 1 4 7 10 All
+Condition Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow Mean Gain Iter.
+Source 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 36.0 / /
+CoTTA 53.7 19.7 38.0 42.4 53.1 19.7 37.7 42.9 51.3 19.0 36.5 41.8 50.9 19.1 36.0 42.4 37.8 +1.8 16k
+SVDP 52.8 20.0 35.6 42.0 54.6 23.5 38.7 43.8 52.8 24.0 38.6 43.9 51.8 23.6 38.2 43.0 39.5 +3.5 16k
+IRG 52.7 20.6 36.0 42.9 53.6 23.2 38.1 45.6 51.0 23.1 37.1 42.8 49.7 22.5 36.5 40.7 38.9 +2.9 16k
+MemCLR 52.9 21.2 35.2 42.8 52.8 23.2 38.2 44.4 51.4 23.2 37.1 42.9 49.9 22.9 36.6 41.0 38.7 +2.7 16k
+Ours-unstop 52.9 20.4 34.5 42.9 54.5 23.8 39.6 45.8 53.5 25.1 39.0 45.3 52.5 24.3 38.8 44.7 40.3 +4.3 16k
+Ours 52.8 20.1 35.2 42.9 54.4 24.6 38.6 45.5 53.5 25.6 38.5 46.2 53.5 25.4 37.4 45.3 40.4 +4.4 13.2k
+Table 3.
+Experimental results (mAP@0.5) and adaptation iterations of Cityscapes-to-ACDC long-term CTTA task.
+We evaluate the performance by continually adapting the source model to the four conditions ten times.
+To save space, we display selected rounds of results, and full results are provided in the appendix.
+“-unstop” denotes no skipping.
+(a) Source
+(b) MemCLR
+(c) CTAOD (ours)
+Figure 3.
+Qualitative results.
+We compare the detection results of the Source model, MemCLR, and CTAOD in the $10^{th}$ round of adaption to defocus corruption on the long-term Cityscapes-to-Cityscapes-C task.
+4.4.1. Synthetic Continual Distribution Shift
+We first evaluate the effectiveness of CTAOD on the short-term Cityscapes-to-Cityscapes-C adaptation tasks.
+As depicted in Table 1, CoTTA, IRG, and Memclr exhibit enhancements in performance, achieving 16.0, 17.5, and 18.0 mAP respectively.
+Although the SFDA methods like IRG and MemCLR assume a static target domain, employing a large momentum update rate $\alpha$ for the teacher enables a relatively stable adaptation in the short term.
+Consequently, CL leads to a better performance than CoTTA which solely employs weight-averaged predictions.
+Moreover, SVDP, which utilizes visual prompts, achieves a sub-optimal performance of 19.1 mAP.
+Remarkably, our proposed method improves the performance to 20.8 mAP, consistently outperforming the above approaches.
+Additionally, the dynamic skipping strategy in AM reduces adaptation iterations by 20% while achieving a 0.2 mAP improvement compared to CTAOD-unstop.
+CTAOD ensures a more reliable and efficient adaptation to target domains characterized by intense and continual changes in the short term.
+As presented in Table 8, the long-term task outcomes reveal the source model’s poor performance, with an average mAP of 16.4.
+Despite the improvement of Memclr and IRG, their performance also begins to decline in the later rounds, failing to maintain stability over long-term adaptation.
+We believe this is due to the aforementioned methods not accounting for continual distribution shifts, resulting in error accumulation and catastrophic forgetting.
+Furthermore, CoTTA utilizes a stochastic restoration mechanism to mitigate forgetting, but its randomness may result in losing crucial information, thus limiting its performance at 19.1 mAP.
+In contrast, SVDP which employs sparse visual prompts raises the performance to 25.3 mAP.
+Particularly, CTAOD yields a remarkable 12.8 mAP enhancement over the Source and surpasses all comparative baselines using 80% adaptation iterations, which employs the DSR mechanism to conserve valuable knowledge while eliminating noise from prior domains.
+These findings empirically validate the effectiveness of our proposed method in ensuring stable and efficient adaptation amidst synthetic continual distribution shifts, across both short-term and long-term scenarios.
+Time $t\xrightarrow{\hskip 195.12767pt}$ All
+Condition Cloudy Overcast Rainy Foggy Mean Gain Iter.
+Source 51.8 41.5 43.8 33.9 42.7 / /
+CoTTA 51.1 40.1 40.7 29.9 40.5 -2.2 19.6k
+SVDP 52.0 41.0 43.8 35.9 43.2 +0.5 19.6k
+IRG 51.9 40.6 42.7 34.3 42.4 -0.3 19.6k
+MemCLR 51.8 40.4 42.6 35 42.4 -0.3 19.6k
+Ours-unstop 52.3 41.3 44.1 36.8 43.6 +0.9 19.6k
+Ours 52.3 41.6 44.4 37.3 43.9 +1.2 16.4k
+Table 4. Experimental results (mAP@0.5) of SHIFT short-term CTTA task. We evaluate the performance by continually adapting the source model to the four conditions.
+4.4.2. Real-World Continual Distribution Shift
+We also evaluate our method on the real-world continual distribution shift dataset.
+The results of the SHIFT short-term CTTA task are shown in Table 4.
+While other baseline methods suffer from performance decline, SVDP which employs a fixed threshold for pseudo-labeling and prompts learning achieves a 0.5 mAP improvement.
+Moreover, CTAOD which introduces a dynamic threshold strategy in AM shows a superior performance of 43.9 mAP with 16% improvement in efficiency.
+The experimental results of ACDC long-term CTTA task are summarized in Table 3.
+CoTTA, IRG, and MemCLR potentially suffer from error accumulation and catastrophic forgetting, manifesting in a rapid decline in performance during the later stages.
+Similarly, SVDP achieves a sub-optimal performance of 39.5 mAP.
+Notably, CTAOD achieves a performance of 40.4 mAP, yielding an absolute improvement of 0.9 mAP over baselines.
+While CTAOD may not display a significant advantage over other methods in the initial round, its performance continually improves and remains stable throughout the long-term adaptation.
+These results underscore the capability of our method to foster robust adaptation toward the real-world continual distribution shift in both the short and long term.
+4.4.3. Qualitative Results
+As shown in Figure 3, we selectively compare the source model, MemCLR, and CTAOD in the final round adaptation to the brightness on the long-term Cityscapes-to-Cityscapes-C tasks.
+CTAOD guides the model in learning better feature representations while effectively mitigating forgetting in the long-term adaptation.
+Therefore, CTAOD assists the detector in distinguishing more foreground object categories and better locating them.
+More visualization analysis is presented in the appendix.
+Mean-Teacher OCL AM DSR Mean Gain
+1 $\checkmark$ 25.9 /
+2 $\checkmark$ $\checkmark$ 26.7 +0.8
+3 $\checkmark$ $\checkmark$ 27.3 +1.4
+4 $\checkmark$ $\checkmark$ 27.4 +1.5
+5 $\checkmark$ $\checkmark$ $\checkmark$ 28.6 +2.7
+6 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 29.2 +3.3
+Table 5.
+Alation experiment.
+“Mean-Teacher” represents the base mean teacher framework with weak-strong augmentation and KL divergence distillation.
+All experiments are done on long-term Cityscapes-to-Cityscapes-C tasks.
+4.4.4. Ablation Study
+We conduct an ablation study to empirically assess the impact of the OCL, AM, and DSR.
+As shown in Table 5, the integration of the mean-teacher architecture elevates the performance to 25.9 mAP.
+Building upon the mean-teacher foundation, the proposed OCL, AM, and DSR achieve an additional 1.5, 1.4, and 0.8 mAP improvement individually.
+Moreover, the combination of the OCL and AM modules raises the performance to 28.6 mAP.
+Furthermore, the inclusion of the DSR mechanism brings a 0.6 mAP boost in performance, culminating in a comprehensive model performance of 29.2 mAP.
+More fine-grained comparisons with AM and DSR are provided in the appendix.
+5. CONCLUSION
+In this work, we proposed CTAOD, aimed at addressing the challenges of low-quality pseudo-labels and random errors resulting from stochastic recovery in continual test-tiem adaptation.
+Firstly, object-level contrastive learning leverages teacher ROI features for contrastive learning to refine the feature representation, tailored for object detection.
+Secondly, the dynamic threshold strategy selects higher-quality pseudo-labels by dynamically updating the category-specific threshold based on the predicted confidence scores.
+Lastly, the data-driven stochastic restoration mechanism selectively reset inactive parameters to mitigate forgetting while retaining valuable knowledge.
+The empirical results of four continual test-time adaptation tasks for object detection demonstrate the efficacy of CTAOD in both short-term and long-term adaption scenarios.
+Appendix A Appendix Overview
+In this appendix, we provide additional details and analysis.
+In the first section, we provide implementation details on CTAOD and other baseline methodologies.
+In the second section C, we furnish complete results for the quantitative research and additional visualization examples for the qualitative research.
+In the third section, we further discuss the continual learning.
+In the last section E, we discuss the limitations and prospective directions of our research.
+Appendix B Additional Implementation Details
+Following (VS et al., 2023b), we trained a Faster R-CNN (Ren et al., 2015) with ResNet50 (He et al., 2016) pre-trained on ImageNet (Krizhevsky et al., 2012) on the training set of Cityscapes (Cordts et al., 2016) as the source-trained model of Cityscapes.
+We utilized an SGD optimizer with a learning rate of 0.001, a momentum of 0.9, and a batch size of 1 and 70k iterations.
+Similarly, we trained another source model on the clear condition of the SHIFT (Sun et al., 2022) with 30 epochs.
+We will specify the implementation details of the CTAOD and other baseline methods in the following subsections.
+B.1. Implememtation of CTAOD
+Detailed hyper-parameters of CTAOD for each benchmark are listed in Table 7.
+Furthermore, we present a pseudo-code outline of our CTAOD framework in Algorithm 1.
+The three key components of CTAOD are highlighted, i.e., Object-level Contrastive Learning, Adaptive Monitoring, and Data-driven Stochastic Restoration.
+B.2. Implememtation of Baseline Methods
+The CoTTA (Wang et al., 2022) and SVDP (Yang et al., 2024) are originally designed for the classification or segmentation tasks.
+We implement them based on their public code, with some modifications to fit into the objection detection.
+Specifically, we implement the Cotta method by constructing a mean-teacher framework based on Faster R-CNN, which incorporates a stochastic restoration mechanism.
+Similarly, we implement SVDP with visual prompts on the area of the bounding box in pseudo-labels based on mean-teacher framework.
+Further, the MemCLR (VS et al., 2023b) and IRG (VS et al., 2023a) are source-free domain adaptation methods for object detection.
+Therefore, we implement the MemCLR and IRG based on their public code, with little modifications to follow the CTTA setting (i.e., updating the model parameters during evaluation).
+For a fair comparison, we set the same momentum $\alpha$ of the Exponential Moving Average (EMA) for CoTTA, SVDP, MemCLR, and IRG as CTAOD, all of which apply the mean-teacher framework (Tarvainen and Valpola, 2017).
+Other hyper-parameters are retained as specified in their public code.
+Appendix C Additional Experiment Results
+C.1. Quantitative Results
+We present the complete experiment results on the two long-term CTTA tasks, i.e., Cityscapes-to-Cityscapes-C and Cityscapes-to-ACDC adaptation tasks, in Table 8 and 9.
+Experiments show that the proposed CTAOD can largely maintain strong performance in the long term.
+C.2. Qualitative Results
+As shown in figure 4 - 8, we present more visualization results to compare the CTAOD and other baseline methods based on the final round adaption on the long-term Cityscapes-to-Cityscapes-C task.
+The results further validate the effectiveness of CTAOD in distinguishing more foreground object categories and better locating them, compared with other baseline methods.
+C.3. Additional Ablation Study
+CTAOD-unstop FT of 0.9 FT of 0.8 FT of 0.7 SR DR Mean
+1 $\checkmark$ $\checkmark$ 27.7
+2 $\checkmark$ $\checkmark$ 28.2
+3 $\checkmark$ $\checkmark$ 28.5
+4 $\checkmark$ $\checkmark$ 27.1
+5 $\checkmark$ $\checkmark$ 28.3
+6 $\checkmark$ 29.0
+Table 6.
+Additional Ablation Study.
+The “CTAOD-unstop” denotes CTAOD without dynamic skipping.
+The “FT” represents replacing dynamic threshold in CTAOD with a fixed threshold.
+The “SR” denotes the stochastic restoration (Wang et al., 2022), and the “DR” indicates the data-driven restoration (Brahma and Rai, 2023).
+All experiments are done on long-term Cityscapes-to-Cityscapes-C tasks.
+As presented in Table 6, we perform an additional ablation study to evaluate the effectiveness of the proposed Dynamic Threshold (DT) strategy in Adaptive Monitoring and Data-driven Stochastic Restoration (DSR).
+The findings indicate that the DT approach results in improvements of 1.3, 0.8, and 0.5 mAP over fixed thresholds set at 0.9, 0.8, and 0.7, respectively.
+While the stochastic restoration and data-driven restoration limit the performance to 27.1 and 28.3 mAP, the DSR mechanism raises the performance to 29.0 mAP.
+Hyper-parameter Description City2City-C-Short City2City-C-Long SHIFT-Short City2ACDC-Long
+$\gamma$ Learning rate 0.001 0.001 0.0001 0.001
+$\alpha$ Momentum in EMA 0.9996 0.9998 0.9999 0.9998
+$\tau$ Temperature in CL 0.07 0.07 0.07 0.07
+$\delta_{0}$ Initial valuse in DT 0.8 0.8 0.8 0.8
+$\beta_{t}$ Hyper-parameter in DT 0.95 0.95 0.95 0.95
+$\epsilon$ Hyper-parameter in DT 1.3 1.3 1.3 1.3
+$\delta_{max}$ The upper-bound in DT 0.9 0.9 0.9 0.9
+$\delta_{mini}$ The lower-bound in DT 0.7 0.7 0.7 0.7
+$\beta_{s}$ Hyper-parameter in DS 0.7 0.75 0.6 0.7
+$\delta_{s}$ The threshold in DS 1.5 1.4 1.4 1.4
+$q$ Quantile value 0.01 0.001 0.001 0.01
+Table 7.
+Detailed hyper-parameters for each benchmark.
+“DT” and “DS” denotes dynamic threshold and dynamic skipping in adaptive monitoring.
+“City2City-C-Short” denotes the short-term Cityscapes-to-Cityscapes-C task, “City2City-C-Long” denotes long-term Cityscapes-to-Cityscapes-C task, and “ City2ACDC-Long” denotes long-term Cityscapes-to-ACDC task.
+Time $t\xrightarrow{\hskip 403.26341pt}$
+Round 1 2 3 4 5 /
+Condition F. M. S. B. D. F. M. S. B. D. F. M. S. B. D. F. M. S. B. D. F. M. S. B. D. / /
+Source 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 / /
+CoTTA 38.2 10.6 0.4 33.8 9.3 39.9 10.7 0.5 34.0 9.9 40.6 10.4 0.6 34.4 9.8 40.9 10.4 0.6 35.4 9.1 40.8 10.4 0.6 36.5 9.2 / /
+SVDP 36.8 8.8 0.6 43.5 11.2 43.9 11.8 1.7 47.9 14.9 46.2 13.9 2.5 50.0 17.8 46.2 14.4 3.2 49.0 19.6 45.0 15.4 3.8 47.8 19.9 / /
+IRG 37.9 9.0 0.7 44.3 12.2 46.1 12.9 2.0 49.2 16.6 47.6 15.2 3.0 49.5 19.5 45.2 17.3 4.6 48.6 21.4 45.6 17.7 6.0 46.7 21.9 / /
+MemCLR 37.7 8.9 0.8 45.5 13.5 45.8 13.6 2.3 49.8 17.7 47.3 16.4 4.1 50.2 20.3 46.4 17.4 4.8 47.8 21.9 45.0 18.1 5.1 46.8 22.4 / /
+Ours-unstop 39.0 10.4 0.8 48.0 13.8 48.2 13.7 3.1 52.9 18.1 49.3 15.9 5.3 53.3 21.3 48.5 17.1 7.1 52.2 23.5 49.4 18.6 7.7 51.7 25.0 / /
+Ours 39.2 9.9 0.8 46.7 12.7 46.8 13.5 2.7 52.3 17.9 49.1 15.8 4.5 53.9 21.2 48.8 17.7 6.9 53.8 23.1 49.4 18.5 8.0 52.9 24.7 / /
+Round 6 7 8 9 10 Mean Gain
+Source 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31 6.7 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 16.4 /
+CoTTA 39.6 10.3 0.6 34.9 9.4 40.2 10.8 0.5 35.3 10.0 40.0 10.8 0.5 35 9.7 40.1 10.5 0.5 34.1 9.5 40.8 10.4 0.5 36.3 9.6 19.1 +2.7
+SVDP 44.4 15.6 4.2 47.6 20.5 44.1 16.1 4.3 46.8 20.4 43.0 15.9 4.6 45.7 20.7 42.6 16.0 4.8 45.0 20.2 41.5 16.2 4.6 44.6 20.2 25.3 +8.9
+IRG 43.9 18 5.9 44.8 22.1 42.3 17.5 6.6 42.9 22.2 41.7 16.9 7.2 42.1 21.8 38.8 17 7.2 40.1 20.6 37.0 16.7 7.1 38.5 20.5 25.6 +9.2
+MemCLR 43.4 19.3 5.9 44.6 23.3 41.1 19.5 6.6 43.1 23.9 39.6 19.1 7.0 40.9 24.3 38.6 18.9 7.2 39.1 22.4 36.8 18.5 7.5 38.4 21.8 26.0 +9.6
+Ours-unstop 48.5 19.7 8.9 50.7 26.3 47.8 19.6 9.8 49.0 26.1 45.7 19.3 10.5 49.4 26.9 45.2 19.7 11.2 47.4 26.9 45.7 20.0 11.8 46.4 25.8 29.0 +12.6
+Ours 48.6 19.8 9.1 51.2 25.4 48.8 20.1 10.0 49.5 27.0 47.3 19.9 10.0 49.2 27.9 46.9 19.4 11.0 47.5 28.9 46.2 19.5 11.3 47.3 29.1 29.2 +12.8
+Table 8.
+Complete experimental results (mAP@0.5) of Cityscapes-to-Cityscapes-C long-term CTTA task.
+We evaluate the long-term adaptation performance by continually adapting the source model to the five corruption ten times with the largest corruption severity level 5.
+“F.”, “M.”, “S.”, “B.”, and “D.” represent Fog, Motion, Snow, Brightness, and Defocus, respectively.
+Appendix D Additional Related Work
+D.1. Continual Learning
 Continual learning, also known as incremental learning or life-long learning, typically involves enabling the model to retain previously acquired knowledge while learning from a sequential series of tasks, i.e.
@@
 Additionally, while the continual learning approaches aim to tackle catastrophic forgetting in sequences of new tasks, our work focuses on learning from different domains for a single task.
-3. METHOD
-3.1. Preliminary
-3.1.1. Problem Statement
-Table 1. Comparisons between different problem settings. ‘Online’ indicates whether the model has access to the target data or only predicts the incoming test samples immediately.
-Setting Source Data Target Training Data Target Distribution Train Loss Test Loss Online
-Continual Learning $\times$ $(x^{t},y^{t})$ Dynamic $\mathcal{L}(x^{t},y^{t})$$\times$ $\times$
-Unsupervised Domain Adaptation $(x^{s},y^{s})$ $(x^{t})$ Static $\mathcal{L}(x^{s},y^{s})+\mathcal{L}(x^{s},x^{t})$$\times$ $\times$
-Source-free Domain Adaptation $\times$ $(x^{t})$ Static $\mathcal{L}(x^{t})$$\times$ $\times$
-Test-time Adaptation $\times$ $\times$ Static $\times$ $\mathcal{L}(x^{t})$checkmark
-Continual Test-Time Adaptation $\times$ $\times$ Dynamic $\times$ $\mathcal{L}(x^{t})$checkmark
-Given a sequence of domain $D=\{D^{0},D^{1},...,D^{n}\}$, we define the first domain $D^{0}$ as the source domain and the subsequent domains as the target domain.
-The objective of CTTA is to enhance the performance of the model $f_{\theta_{0}}(x)$, where the parameters $\theta_{0}$ are pre-trained on source data $x^{D^{0}},y^{D^{0}}$ from $D^{0}$, in a continually changing target domain during inference time without using source data.
-For simplicity, we denote the data without the superscript about the specific domain in the following discussion.
-At time step $t$, unlabeled target data $x_{t}$ is provided sequentially, following the domain group order.
-The model is required to make a prediction $\hat{y}_{t}=f_{\theta_{t-1}}(x_{t})$ using the parameters $\theta_{t-1}$ which have been updated based on previous target data $x_{1},...,x_{t-1}$.
-Subsequently, $\hat{y}_{t}$ serves as the evaluation output at time step t, and the model will adapt itself toward $x_{t}$ as $\theta_{t-1}\rightarrow\theta_{t}$, which will only influence future inputs $x_{t+n}$.
-Moreover, We compare CTTA with other adaptation settings, as detailed in Table 1. These settings are developed to meet the diverse prerequisites and requirements for real-world applications.
-3.1.2. Mean-Teacher Framework
-In alignment with previous CTTA works (Wang et al., 2022; Gong et al., 2022; Döbler et al., 2023; Niloy et al., 2024), we build our method based on the mean-teacher framework (Tarvainen and Valpola, 2017).
-This framework is characterized by the interplay between a teacher model and a student model.
-Both networks are initialized with the source-trained model at the beginning of adaptation.
-The teacher model produces the pseudo-labels for the unlabeled target data, which serve as labels of the unlabeled data to supervise the student model.
-While the parameters of the student model are optimized via gradient descent, the parameters of the teacher model are updated following an Exponential Moving Average (EMA) strategy based on the student model.
-Formally, this process can be expressed as follows:
-(1) $\mathcal{L}_{pl}(x_{t})=\mathcal{L}_{rpn}(x_{t},\hat{y}_{t})+\mathcal{L}_{rcnn}(x_{t},\hat{y}_{t}),$
-(2) $\theta_{t}^{std}\leftarrow\theta_{t-1}^{std}+\gamma\frac{\partial(\mathcal{L}_{pl}(x_{t}))}{\partial\theta_{t-1}^{std}},$
-(3) $\theta_{t}^{tch}\leftarrow\alpha\theta_{t-1}^{tch}+(1-\alpha)\theta_{t}^{std},$
-where $x_{t}$ and $\hat{y}_{t}$ denote the unlabeled target data and corresponding pseudo-labels generated by the teacher network.
-$\theta^{std}_{t}$ and $\theta^{tch}_{t}$ symbolize the parameters of student and teacher networks at time step $t$, respectively.
-Moreover, the supervision loss $\mathcal{L}_{pl}$, which consists of the rpn loss $\mathcal{L}_{rpn}$ and rcnn loss $\mathcal{L}_{rcnn}$ in faster-rcnn (Ren et al., 2015), is utilized for pseudo-labeling.
-Additionally, $\gamma$ represents the student’s learning rate, and the EMA rate is denoted by $\alpha$.
-Hence, the teacher can be regarded as an ensemble of historical students, providing stable supervision.
-Despite the effectiveness of the mean-teacher framework under the static distribution shift, this framework encounters error accumulation and catastrophic forgetting in non-stationary environments.
-Therefore, we design CTAOD to improve the robustness of feature adaptation in the CTTA setting.
-Figure 2.
-The overview of the proposed CTAOD, which follows the mean-teacher framework.
-The teacher generates proposals and predictions for the weakly augmented image, whereas the student receives the strongly enhanced images.
-Our proposed CTAOD includes: (1) The Object-level Contrastive Learning (OCL) module compares region of interest features extracted from feature maps of both networks based on teacher proposals for contrastive learning; (2) The dynamic threshold (DT) strategy adjusts category-specific thresholds based on prediction scores; These pseudo-labels are utilized to supervise the student through the threshold.
-(3) The Data-driven stochastic restoration (DSR) reset the inactive parameter based on the gradient.
-3.2. Proposed Method
-3.2.1. Overview
-As indicated in Figure 2, this subsection presents a concise overview of our proposed CTAOD, consisting of Object-level Contrastive Learning (OCL), Dynamic Threshold (DT), and Data-driven Stochastic Restoration (DSR).
-Inspired by UDA methods for object detection (Li et al., 2022a; Liu et al., 2021), we adopt the Weak-Strong augmentation to enable the teacher model to generate reliable pseudo-labels without being affected by heavy augmentation.
-Specifically, the teacher model receives input with weak augmentations while we input student networks with strong enhanced images.
-The overall loss of the student model for gradient descent is defined as:
-(4) $\mathcal{L}_{all}=\mathcal{L}_{pl}(x_{t})+\lambda\mathcal{L}_{cl}(x_{t})+\mu\mathcal{L}_{kl}(x_{t}),$
-where $\mathcal{L}_{cl}(x_{t})$ denotes the contrastive learning (CL) loss, to be elaborated in section 3.2.2.
-The $\mathcal{L}_{kl}$ represents the Kullback-Leibler (KL) Divergence (Kullback and Leibler, 1951) loss, used to quantify the distinction between two probability distributions.
-The parameters $\lambda$ and $\mu$ are corresponding weights of the losses.
-The KL loss is defined as:
-(5) $\mathcal{L}_{kl}\left(P\parallel Q\right)=\sum_{x\in\mathcal{X}}P(x)\mathrm{~{}}\log\left(\frac{P(x)}{Q(x)}\right),$
-where $P(x)$ and $Q(x)$ represent different distribution. We employ the KL divergence loss to encourage the student model to approximate the teacher model closely.
-3.2.2. Object-Level Contrastive Learning
-In this subsection, we first briefly introduce SimCLR, a widely used CL approach for self-supervised visual representation learning (Chen et al., 2020).
-SimCLR aims to learn high-quality feature representation across differently augmented views of the same image.
-For the $i^{th}$ image $x_{i}$ in a batch, the CL loss in SimCLR is formulated as follows:
-(6) $\mathcal{L}_{cl}(x_{i})=-\log\left(\frac{\exp(sim(z_{i},z_{j})/\tau)}{\sum_{k=1,\ni_{k\neq i}}^{2N}\exp(sim(z_{i},z_{k})/\tau)}\right),$
-where $\tau>0$ is a temperature hyper-parameter.
-$z_{i}$ and $z_{j}$ denote the features of two different augmentations of the same images, serving as the positive pair.
-At the same time, $z_{k}$ represents one of the features of a different augmented sample in a batch size of $N$.
-Here, the $sim(\cdot)$ indicates the similarity function, i.e.
-cosine similarity.
-Notably, the SimCLR is initially designed for classification tasks, assuming each image pertains to a single category, and requires large batch sizes to ensure sufficient positive and negative pairs for feature representation learning.
-Consequently, the original SimCLR is not well-suited for object detection, where images typically contain multiple instances, thus requiring significant computational resources to accommodate large batch sizes.
-Motivated by SimCLR (Chen et al., 2020), we present an OCL module to extract teacher and student features for CL based on proposals generated from the region proposal network (RPN).
-Since proposals provide multiple cropped views around the object instance, this strategy does not require a large batch size, thus computationally efficient for online detection model updates.
-Specifically, given a weakly augmented image $f_{weak}(x_{t})$ at time step t, the teacher produces ROI proposals $p_{t}=\{p_{1},...,p_{l}\}$ via region proposal network.
-We then apply RoIAlign (He et al., 2017), a pooling operation for ROI, to extract corresponding teacher and student object-level features $F_{t}^{T}=\{f_{i}^{T}\in\mathbb{R}^{1\times C}\}_{i=1}^{l}$ and $F_{t}^{S}=\{f_{i}^{S}\in\mathbb{R}^{1\times C}\}_{i=1}^{l}$ based on the feature map from the backbone, respectively.
-Subsequently, the features are projected into the final object-level features with the same encoder using two fully connected layers.
-For simplicity, we still denote these two final features as $F_{t}^{T}$ and $F_{t}^{S}$.
-The features associated with the same proposal are considered positive pairs, otherwise negative pairs.
-we employ the CL loss to these features $F_{t}^{T}$ and $F_{t}^{S}$ derived from weak and strong enhanced images respectively by minimizing:
-(7) $\mathcal{L}_{cl}(x_{t})=\frac{1}{l}\sum_{i=1}^{l}-\log\frac{\exp(f_{i}^{T}\cdot f_{i}^{S}/\tau)}{\sum_{j=1}^{l}\exp(f_{i}^{T}\cdot f_{j}^{S}/\tau)},$
-where $l$ indicates the number of the features.
-This strategy encourages the model to learn fine-grained and localized representations on the target domain, without relying on accurate pseudo-labels.
-Moreover, the OCL is well integrated into the mean-teacher self-training paradigm as a drop-in enhancement for feature adaptation.
-3.2.3. Dynamic Threshold
-Due to the presence of domain shifts, pseudo-labels inevitably suffer from label noise.
-Additionally, the common practice is to use a fixed threshold to remove suspected noisy labels.
-However, model confidence could differ between different categories and domains, thereby hurting the performance using a fixed threshold.
-Taking inspiration from the threshold strategy (Wang et al., 2024; Zhao et al., 2023), we design the DT strategy to update the category-specific threshold based on the predicted scores dynamically to minimize noise and select a suitable threshold for each category across different domains in test-time.
-As shown in Figure 2, the teacher network are first to make the prediction $\hat{y}_{t}=f_{\theta_{t-1}}(f_{wek}(x_{t}))$ for the weakly enhanced images $f_{wek}(x_{t})$ The thresholds for each category are initialized with the same value $\delta_{0}$ at the start of the adaptation, which is updated every iteration by:
-(8) $\delta_{t}^{c}\leftarrow\beta\cdot\delta_{t-1}^{c}+(1-\beta)\cdot\epsilon\cdot(\overline{l_{t}^{c}})^{\frac{1}{2}},$
-where $\delta_{t}^{c}$ denotes the threshold of category $c$ at time step t, $\overline{l_{t}^{c}}$ is the mean predicted scores of the category $c$ at time step $t$, $\beta$ represents the update rate, and $\epsilon$ provides a linear projection.
-Furthermore, the threshold $\delta_{t}^{c}$ will not change if class $c$ does not exist in prediction $\hat{y}_{t}$, and we set a fixed upper and lower bound $\delta_{max}$ and $\delta_{mini}$.
-For unlabeled data from dynamic environments, this mechanism effectively prevents the threshold from being too high or too low while generating an appropriate threshold for each category.
-3.2.4. Data-Driven Stochastic Restoration
-Existing methods address catastrophic forgetting through stochastic reset mechanisms (Wang et al., 2022; Zhu et al., 2023) or pure data-driven restoration (Brahma and Rai, 2023).
-Nevertheless, the former may reset valuable parameters, potentially erasing essential knowledge relevant to the current domain.
-On the other hand, the latter restores the most unimportant parameters, which may retain noise parameters, thereby leading to error accumulation.
-Taking inspiration from (Wang et al., 2022; Brahma and Rai, 2023), we propose the DSR strategy to reset irrelevant parameters with higher possibilities, enabling robust adaptation.
-Following (Brahma and Rai, 2023), we adopt the diagonal approximation of Fisher Information Matrix (FIM) (Kirkpatrick et al., 2017) based on gradient, as a metric of parameter importance.
-The FIM approximates parameter importance by considering the sensitivity of the loss function to parameter changes (Kirkpatrick et al., 2017).
-Parameters with higher values in the FIM are those for which small changes result in significant increases in loss, indicating their importance for domain at hand.
-Specifically, consider a convolution layer of the student model with weight after gradient descent $W_{t-1}$ and its corresponding gradient matrix $G_{t}$ with the same shape as $W_{t-1}$ at time step $t$.
-First, we use a random matrix $R_{t}\sim\operatorname{Uniform}(0,1)$ with the same shape as the weights matrix, where each element follows a uniform distribution between 0 and 1.
-Moreover, we utilize the element-wise multiplication of the square of the gradient and the random matrix to approximate the FIM (Kirkpatrick et al., 2017):
-(9) $F_{t}=G_{t}\odot G_{t}\odot R_{t},$
-where $\odot$ represents the element-wise multiplication. Furthermore, the proposed DSR updates the parameters based on the FIM values, which is formulated as:
-(10) $M_{t}=F_{t}<\eta,$
-(11) $W_{t}=M_{t}\odot W_{0}+(1-M_{t})\odot W_{t-1},$
-where $M_{t}$ indicates the mask matrix of the same shape as $W_{t}$, $<$ represents the element-wise less than operation, and $\eta$ is the threshold value which is acquired by the $q$-quantile of $F_{t}$, i.e.
-$\eta=quantile(F_{t},q)$.
-Consequently, the elements in $M_{t}$ are set to 1, when the corresponding FIM value is less than $\eta$, indicating the corresponding parameters should be reset to the source weight $W_{0}$.
-This strategy enables the model to retain essential knowledge while introducing randomness to enhance robustness, thereby mitigating the forgetting issue in CTTA.
-4. EXPERIMENTS
-Table 2.
-Experimental results (mAP@0.5) of Cityscapes-to-Cityscapes-C short-term CTTA task.
-We evaluate the short-term adaptation performance by continually adapting the source-trained model to the twelve corruptions with the largest corruption severity level 5.
-Time $t\xrightarrow{\hskip 325.215pt}$ All
-Condition Defocus Glass Motion Zoom Snow Frost Fog Brightness Contrast Elastic Pixelate Jpeg Mean Gain
-Source (Ren et al., 2015) 6.8 8.1 8.0 1.5 0.2 6.8 34.6 30.7 3.0 50.2 17.6 13.5 15.1 /
-Tent (Wang et al., 2020) 6.8 7.8 7.7 1.3 0.2 6.1 33.1 28.0 2.2 51.1 14.8 11.0 14.2 -0.9
-CoTTA (Wang et al., 2022) 7.8 9.0 8.9 1.8 0.3 7.1 38.4 31.1 8.6 49.6 16.2 13.1 16.0 +0.9
-SVDP (Yang et al., 2024) 7.7 10.1 9.7 2.3 0.7 13.0 42.4 45.2 15.4 47.2 21.2 14.8 19.1 +4.0
-IRG (VS et al., 2023a) 8.0 11.0 9.3 3.4 1.2 13.0 37.9 41.3 15.9 38.9 16.9 13.4 17.5 +2.4
-MemCLR (VS et al., 2023b) 8.5 10.4 10.6 2.7 1.1 12.2 41.4 41.6 16.4 43.1 15.4 12.7 18.0 +2.9
-Ours 8.6 12.1 11.7 3.6 1.5 16.7 44.7 48.1 16.7 47.4 22.5 13.9 20.6 +5.5
-Table 3.
-Experimental results (mAP@0.5) of Cityscapes-to-Cityscapes-C long-term CTTA task.
-We evaluate the long-term adaptation performance by continually adapting the source model to the five corruption ten times with the largest corruption severity level 5.
-To save space, we display selected rounds of results, and full results are provided in the supplementary material.
-Time $t\xrightarrow{\hskip 390.25534pt}$
-Round 1 5 10 All
-Condition Fog Motion Snow Brightness Defocus Fog Motion Snow Brightness Defocus Fog Motion Snow Brightness Defocus Mean Gain
-Source (Ren et al., 2015) 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 16.4 /
-Tent (Wang et al., 2020) 35.8 8.1 0.2 29.2 6.2 30.7 10.2 0.8 27.9 12.4 20.4 5.0 0.1 11.5 2.4 12.0 -4.4
-CoTTA (Wang et al., 2022) 38.2 10.6 0.4 33.8 9.3 40.8 10.4 0.6 36.5 9.2 40.8 10.4 0.5 36.3 9.6 19.1 +2.7
-SVDP (Yang et al., 2024) 36.8 8.8 0.6 43.5 11.2 45.0 15.4 3.8 47.8 19.9 41.5 16.2 4.6 44.6 20.2 25.3 +8.9
-IRG (VS et al., 2023a) 37.9 9.0 0.7 44.3 12.2 45.6 17.7 6.0 46.7 21.9 37.0 16.7 7.1 38.5 20.5 25.6 +9.2
-MemCLR (VS et al., 2023b) 37.7 8.9 0.8 45.5 13.5 45.0 18.1 5.1 46.8 22.4 36.8 18.5 7.5 38.4 21.8 26.0 +9.6
-Ours 39.0 10.4 0.8 48.0 13.8 49.4 18.6 7.7 51.7 25.0 45.7 20.0 11.8 46.4 25.8 29.0 +12.6
-In this study, we rigorously evaluate our methodology across four benchmark tasks tailored for CTTA in object detection.
-Four tasks encompass continual adaptation to synthetic and real-world distribution shifts, evaluated over short and long-term periods.
-Inspired by the foundational work CoTTA (Wang et al., 2022), the short-term CTTA task entails the sequential adaptation to various target domains once.
-In contrast, the long-term CTTA task involves continually adapting the model toward groups of target domains cyclically.
-4.1. Datasets
-Cityscapes.
-The Cityscapes dataset (Cordts et al., 2016) is collected for urban scene understanding, encompassing 2,975 training images and 500 validation images with eight object types, i.e.
-person, rider, car, truck, bus, train, motorcycle, and bicycle.
-We utilize the model pre-trained on the training set of Cityscapes as the source model, and the source data is discarded during adaptation.
-Cityscapes-C.
-The Cityscapes-C (Hendrycks and Dietterich, 2019) is initially designed to assess robustness against various corruptions, introducing 15 types of corruption with 5 levels of severity.
-We create the dataset by applying these corruptions at the maximum severity level 5 to the validation set of clean cityscapes, treating each corruption type as an individual target domain comprising 500 images.
-For Cityscapes-C, both short-term and long-term CTTA tasks are explored.
-Our short-term CTTA task selectively focuses on the latter 12 corruptions, i.e.
-Defocus Blur, Frosted Glass Blur, Motion Blur, Zoom Blur, Snow, Frost, Fog, Brightness, Contrast, Elastic, Pixelate, and JPEG.
-For the long-term CTTA task, we prioritize the five corruptions with direct relevance to autonomous driving scenarios as the target domain groups following (Gan et al., 2023), namely Fog, Motion, Snow, Brightness, and Defocus.
-To mimic the scenario in real life where similar environments might be revisited, and to evaluate the forgetting effect of our methods in the long term, we repeat the adaptation to the target domain group 10 times.
-SHIFT.
-The SHIFT (Sun et al., 2022) is a synthetic dataset for autonomous driving, featuring real-world environmental changes.
-SHIFT can be categorized as clear, cloudy, overcast, rainy, and foggy, where each condition contains images taken at various times ranging from daytime to night.
-For SHIFT, short-term CTTA tasks are considered.
-We designate the clear condition as the source domain, with the remaining four conditions as the target domain groups including nearly 20k images in total.
-ACDC.
-The ACDC dataset (Sakaridis et al., 2021), namely the adverse conditions dataset, shares the same class types as Cityscapes and is collected in four different adverse visual conditions, including Fog, Night, Rain, and Snow.
-Following (Wang et al., 2022; Gan et al., 2023), we use these four conditions as the target domain group for the long-term CTTA task, with 400 unlabeled images from each condition.
-Similarly, we adapt the source-trained model to continually to the target domain groups 10 cycles for the long-term CTTA task.
-4.2. Implementation Details
-We adopt the Faster R-CNN (Ren et al., 2015) with ResNet50 (He et al., 2016) pre-trained on ImageNet (Krizhevsky et al., 2012) as the backbone.
-Aligning with (VS et al., 2023b; Wang et al., 2022), we maintain a batch size of 1, thereby emulating a real-world application scenario where the object detection model is required to adapt toward a continuous influx of images.
-The source and student models are trained using an SGD optimizer with a learning rate of 0.001 and a momentum of 0.9.
-Algorithms are implemented leveraging the Detectron2 platform (Wu et al., 2019).
-The metric of mAP at an IoU threshold of 0.5 (mAP@0.5) is employed for evaluation.
-Each experiment is conducted on 1 NVIDIA A800 GPU.
-More implementation details are provided in the supplementary material.
-4.3. Baselines and Compared Approaches
-To establish the efficacy of our proposed CTAOD, we compare it with five baselines across various setting types for a fair comparison, including Source (Ren et al., 2015), IRG (VS et al., 2023a), Memclr (VS et al., 2023b), Tent (Wang et al., 2020), and CoTTA (Wang et al., 2022).
-Specifically, ”Source” represents the source model, i.e., Faster R-CNN (Ren et al., 2015), solely pre-trained on the source domain.
-Moreover, Memclr (VS et al., 2023b) and IRG (Ren et al., 2015) are SFDA object detection methods.
-MemCLR (VS et al., 2023b) integrates cross-attention with contrastive learning (CL), while IRG (VS et al., 2023a) incorporates instance relation graph and CL.
-Furthermore, Tent (Wang et al., 2020) updates the affine parameters through entropy minimization in TTA.
-CoTTA (Wang et al., 2022) utilizes weight-averaged predictions and random recovery neurons to tackle CTTA, while SVDP (Yang et al., 2024) explores sparse visual prompts for CTTA dense prediction.
-Table 4. Experimental results (mAP@0.5) of SHIFT short-term CTTA task. We evaluate the adaptation performance by continually adapting the source model to the four conditions.
-Time $t\xrightarrow{\hskip 195.12767pt}$ All
-Condition Cloudy Overcast Rainy Foggy Mean Gain
-Source 51.8 41.5 43.8 33.9 42.7 /
-Tent 50.9 39.5 36.3 23.1 37.5 -5.2
-CoTTA 51.1 40.1 40.7 29.9 40.5 -2.2
-SVDP 52.0 41.0 43.8 35.9 43.2 +0.5
-IRG 51.9 40.6 42.7 34.3 42.4 -0.3
-MemCLR 51.8 40.4 42.6 35 42.4 -0.3
-Ours 52.3 41.3 44.1 36.8 43.6 +0.9
-4.4. Results and Analysis
-Table 5.
-Experimental results (mAP@0.5) of Cityscapes-to-ACDC long-term CTTA task.
-We evaluate the long-term adaptation performance by continually adapting the source model to the four conditions ten times.
-To save space, we display selected rounds of results, and full results are provided in the supplementary material.
-Time $t\xrightarrow{\hskip 368.57964pt}$
-Round 1 4 7 10 All
-Condition Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow Mean Gain
-Source (Ren et al., 2015) 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 36.0 /
-Tent (Wang et al., 2020) 52.4 18.6 33.4 38.9 51.7 17.4 31.4 36.0 45.8 14.6 28.1 28.5 35.0 9.5 21.5 18.3 30.5 -5.5
-CoTTA (Wang et al., 2022) 53.7 19.7 38.0 42.4 53.1 19.7 37.7 42.9 51.3 19.0 36.5 41.8 50.9 19.1 36.0 42.4 37.8 +1.8
-SVDP (Yang et al., 2024) 52.8 20.0 35.6 42.0 54.6 23.5 38.7 43.8 52.8 24.0 38.6 43.9 51.8 23.6 38.2 43.0 39.5 +3.5
-IRG (VS et al., 2023a) 52.7 20.6 36.0 42.9 53.6 23.2 38.1 45.6 51.0 23.1 37.1 42.8 49.7 22.5 36.5 40.7 38.9 +2.9
-MemCLR (VS et al., 2023b) 52.9 21.2 35.2 42.8 52.8 23.2 38.2 44.4 51.4 23.2 37.1 42.9 49.9 22.9 36.6 41.0 38.7 +2.7
-Ours 52.9 20.4 34.5 42.9 54.5 23.8 39.6 45.8 53.5 25.1 39.0 45.3 52.5 24.3 38.8 44.7 40.3 +4.3
-4.4.1. Synthetic Continual Distribution Shift
-We first evaluate the effectiveness of the proposed method on the synthetic continual shift dataset (i.e.
-Cityscapes-to-Cityscapes-C adaptation tasks) in the short term.
-As depicted in Table 2, Tent (Wang et al., 2020) undergoes a slight decline in performance, dropping from 15.1 to 14.2 mAP relative to the source model.
-This downturn may be attributed to its dependency on a large batch size to update the parameters of the batchnorm layer, making it suboptimal for online adaptation.
-In contrast, CoTTA (Wang et al., 2022), IRG (VS et al., 2023a), and Memclr (VS et al., 2023b) exhibit enhancements in performance, achieving 16.0, 17.5, and 18.0 mAP respectively.
-Although the SFDA methods like IRG and MemCLR assume a static target domain, employing a large momentum update rate $\alpha$ for the teacher enables a relatively stable adaptation in the short term.
-Consequently, CL leads to a better performance than CoTTA which employs weight-averaged predictions.
-However, SVDP that utilizes visual prompts achieves the sub-optimal performance of 19.1 mAP.
-Remarkably, our proposed method improves the performance to 20.6 mAP, consistently outperforming the above approaches.
-CTAOD ensures a more reliable adaptation to target domains characterized by intense and continual changes in the short term with high-quality pseudo-labels and feature representation.
-As presented in Table 9, the long-term task outcomes reveal the source model’s relatively poor performance, with an average mAP of 16.4.
-Tent’s performance is markedly deteriorative across the timeline.
-Despite the improvement of Memclr and IRG, their performance also begins to decline in the later rounds, failing to maintain stability over long-term adaptation.
-We believe this is due to the aforementioned methods not accounting for continual distribution shifts, resulting in error accumulation and catastrophic forgetting.
-Furthermore, CoTTA utilizes a stochastic restoration mechanism to mitigate forgetting, but its randomness may result in losing crucial information, thus limiting its performance at 19.1 mAP.
-In contrast, SVDP that employ sparse visual prompts raise the performance to 25.3 mAP.
-Particularly, CTAOD yields a remarkable 12.6 mAP enhancement over the Source and surpasses all comparative baselines, which employs the DSR mechanism to conserve valuable knowledge while eliminating noise from prior domains.
-These findings empirically validate the effectiveness of our proposed method in ensuring stable adaptation amidst synthetic continual distribution shifts, across both short-term and long-term adaptation scenarios.
-(a) Source
-(b) MemCLR
-(c) CTAOD (ours)
-Figure 3.
-Qualitative results.
-We compare the detection results of the Source model, MemCLR, and CTAOD in the $10^{th}$ round of adaption to brightness (row 1) and defocus (row 2) corruption on the long-term Cityscapes-to-Cityscapes-C task.
-4.4.2. Real-World Continual Distribution Shift
-We also evaluate our method on the real-world continual distribution shift dataset (i.e.
-SHIFT and Cityscapes-to-ACDC adaptation tasks).
-For the SHIFT short-term CTTA task, where each domain includes more complex images ranging from daytime to night, the results is shown in Table 4.
-While other baseline method suffer from performance decline, SVDP that employs fixed threshold for pseudo-labeling and prompts learning achieve a 0.5 mAP improvements.
-Moreover, CTAOD that introduces a DT strategy shows a superior performance of 43.6 mAP.
-The experimental results of ACDC long-term CTTA task are summarized in Table 5.
-Tent, CoTTA, IRG, and MemCLR potentially suffer from error accumulation and catastrophic forgetting, manifesting in a rapid decline in performance during the later stages.
-Similarly, SVDP achieves a sub-optimal performance of 39.5 mAP.
-Notably, CTAOD achieves a performance of 40.3 mAP, yielding an absolute improvement of 0.8 mAP over baselines.
-While CTAOD may not display a significant advantage over other methods in the initial round, its performance continually improves and remains stable throughout the long-term adaptation.
-This outcome underscores the capability of our method to foster robust adaptation toward the real-world continual distribution shift in the long term.
-4.4.3. Qualitative Results
-As shown in Figure 3, we provide some examples of detection results.
-We selectively compare the source model, MemCLR, and CTAOD in the final round adaptation to the brightness and defocus corruption on the long-term Cityscapes-to-Cityscapes-C tasks.
-Compared with Source and MemCLR, CTAOD guides the model to learn better feature representations while effectively mitigating forgetting in the long-term adaptation.
-Therefore, CTAOD assists the detector in distinguishing more foreground object categories and better locating them.
-More visualization analysis is presented in the supplementary material.
-Table 6.
-Alation experiment.
-“Mean-Teacher” represents the base mean teacher framework with weak-strong augmentation and KL divergence distillation.
-All experiments are done on long-term Cityscapes-to-Cityscapes-C tasks.
-Mean-Teacher OCL DT DSR Mean Gain
-1 $\checkmark$ 25.9 /
-2 $\checkmark$ $\checkmark$ 27.4 +1.5
-3 $\checkmark$ $\checkmark$ 27.1 +1.2
-4 $\checkmark$ $\checkmark$ 26.7 +0.8
-5 $\checkmark$ $\checkmark$ $\checkmark$ 28.5 +12.1
-6 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 29.0 +12.6
-4.4.4. Ablation Study
-We conduct an ablation study to empirically assess the impact of the proposed OCL, DT, and DSR.
-As shown in Table 6, the integration of the mean-teacher architecture, which incorporates weak-strong augmentation and KL divergence distillation, elevates the performance to 25.9 mAP.
-Building upon the mean-teacher foundation, the proposed OCL, DT, and DSR achieve an additional 1.5, 1.2, and 0.8 mAP improvement individually.
-Moreover, the combination of the OCL and DT modules raises the performance to 28.5 mAP.
-Furthermore, the inclusion of the DSR mechanism brings a 0.5 mAP boost in performance, culminating in a comprehensive model performance of 29.0 mAP
-5. CONCLUSION
-In this work, we proposed CTAOD, aimed at addressing the challenges of low-quality pseudo-labels and random errors resulting from stochastic recovery in continual test-tiem adaptation.
-Firstly, object-level contrastive learning leverages teacher ROI features for contrastive learning to refine the feature representation, tailored for object detection.
-Secondly, the dynamic threshold strategy selects higher-quality pseudo-labels by dynamically updating the category-specific threshold based on the predicted confidence scores.
-Lastly, the data-driven stochastic restoration mechanism selectively reset inactive parameters to mitigate forgetting while retaining valuable knowledge.
-The empirical results of four continual test-time adaptation tasks for object detection demonstrate the efficacy of CTAOD in both short-term and long-term adaption scenarios.
-Appendix A Appendix Overview
-In this supplementary material, we provide additional details and analysis.
-In Section B, we provide implementation details on CTAOD and other baseline methodologies.
-In Section C, we furnish complete results for the quantitative research and additional visualization examples for the qualitative research.
-In Section D, we discuss the limitations and prospective directions of our research.
-Appendix B Additional Implementation Details
-Following (VS et al., 2023b), we trained a Faster R-CNN (Ren et al., 2015) with ResNet50 (He et al., 2016) pre-trained on ImageNet (Krizhevsky et al., 2012) on the training set of Cityscapes (Cordts et al., 2016) as the source-trained model.
-We utilized an SGD optimizer with a learning rate of 0.001, a momentum of 0.9, and a batch size of 1 and 70k iterations.
-We will specify the implementation details of the CTAOD and other baseline methods in the following subsections.
-B.1. Implememtation of CTAOD
-Detailed hyper-parameters of CTAOD for each benchmark are listed in Table 8.
-Furthermore, we present a pseudo-code outline of our CTAOD framework in Algorithm 1.
-The three key components of CTAOD are highlighted, i.e., Object-level Contrastive Learning, Dynamic Threshold, and Data-driven Stochastic Restoration.
-B.2. Implememtation of Baseline Methods
-The Tent (Wang et al., 2020), CoTTA (Wang et al., 2022), and SVDP (Yang et al., 2024) are originally designed for the classification or segmentation tasks.
-We implement them based on their public code, with some modifications to fit into the objection detection.
-Specifically, the Tent approach is implemented by updating the parameters of the batch normalization in Faster R-CNN through entropy minimization.
-Meanwhile, we implement the Cotta method by constructing a mean-teacher framework based on Faster R-CNN, which incorporates a stochastic restoration mechanism.
-Similarly, we implement SVDP with visual prompts on the area of the bounding box in pseudo-labels based on mean-teacher framework.
-Further, the MemCLR (VS et al., 2023b) and IRG (VS et al., 2023a) are source-free domain adaptation methods for object detection.
-Therefore, we implement the MemCLR and IRG based on their public code, with little modifications to follow the CTTA setting (i.e., updating the model parameters during evaluation).
-For a fair comparison, we set the same momentum $\alpha$ of the Exponential Moving Average (EMA) for CoTTA, SVDP, MemCLR, and IRG as CTAOD, all of which apply the mean-teacher framework (Tarvainen and Valpola, 2017).
-Other hyper-parameters are retained as specified in their public code.
-Appendix C Additional Experiment Results
-C.1. Quantitative Results
-We present the complete experiment results on the two long-term CTTA tasks, i.e., Cityscapes-to-Cityscapes-C and Cityscapes-to-ACDC adaptation tasks, in Table 9 and 10.
-Experiments show that the proposed CTAOD can largely maintain strong performance in the long term.
-C.2. Qualitative Results
-As shown in figure 4 - 8, we present more visualization results to compare the CTAOD and other baseline methods based on the final round adaption on the long-term Cityscapes-to-Cityscapes-C task.
-The results further validate the effectiveness of CTAOD in distinguishing more foreground object categories and better locating them, compared with other baseline methods.
-C.3. Additional Ablation Study
-Table 7.
-Additional Ablation Study.
-The “FT” represents replacing dynamic threshold in CTAOD with a fixed threshold.
-The “SR” denotes the stochastic restoration (Wang et al., 2022), and the “DR” indicates the data-driven restoration (Brahma and Rai, 2023).
-All experiments are done on long-term Cityscapes-to-Cityscapes-C tasks.
-CTAOD FT of 0.9 FT of 0.8 FT of 0.7 SR DR Mean
-1 $\checkmark$ $\checkmark$ 27.7
-2 $\checkmark$ $\checkmark$ 28.2
-3 $\checkmark$ $\checkmark$ 28.5
-4 $\checkmark$ $\checkmark$ 27.1
-5 $\checkmark$ $\checkmark$ 28.3
-6 $\checkmark$ 29.0
-As presented in Table 7, we perform an additional ablation study to evaluate the effectiveness of the proposed Dynamic Threshold (DT) and Data-driven Stochastic Restoration (DSR).
-The findings indicate that the DT approach results in improvements of 1.3, 0.8, and 0.5 mAP over fixed thresholds set at 0.9, 0.8, and 0.7, respectively.
-While the stochastic restoration and data-driven restoration limit the performance to 27.1 and 28.3 mAP, the DSR mechanism raises the performance to 29.0 mAP.
-Appendix D Limitation
+Appendix E Limitation
 Firstly, our experiments are primarily based on a single detector backbone, i.e.
@@
 Consequently, a promising avenue for future research would be to apply our methodology in practical, real-world systems to further validate its effectiveness.
+Time $t\xrightarrow{\hskip 390.25534pt}$
+Round 1 2 3 4 5 /
+Condition Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow / /
+Source 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 / /
+CoTTA 53.7 19.7 38 42.4 53.6 20.3 37.7 42.2 53.7 20.1 38.1 43.3 53.1 19.7 37.7 42.9 52.6 19.5 37 42.2 / /
+SVDP 52.8 20.0 35.6 42.0 53.4 22.0 37.4 43.8 54.0 23.0 38.8 44.0 54.6 23.5 38.7 43.8 53.4 24.0 38.0 44.4 / /
+IRG 52.7 20.6 36 42.9 53.2 22.4 37.4 44.6 53.8 23.2 38.1 45.5 53.6 23.2 38.1 45.6 53.5 23.2 38.1 45.3 / /
+MemCLR 52.9 21.2 35.2 42.8 52.5 23.1 37.3 44.3 53.9 23.2 38.4 44.8 52.8 23.2 38.2 44.4 51.5 23.5 37.2 43.8 / /
+Ours-unstop 52.9 20.4 34.5 42.9 53 22.9 38.6 44.6 54.2 24 39.5 45.3 54.5 23.8 39.6 45.8 54.4 24.4 39.2 45.4 / /
+Ours 52.8 20.1 35.2 42.9 53.6 22.2 38.3 44.2 53.9 23.5 39.6 45.0 54.4 24.6 38.6 45.5 54.6 25.1 39.2 46.3 / /
+Round 6 7 8 9 10 Mean Gain
+Source 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 36 /
+CoTTA 52.3 19.1 36.3 42 51.3 19 36.5 41.8 51.2 18.9 36.1 41.4 51.5 19.2 36.4 41.3 50.9 19.1 36 42.4 37.8 +1.8
+SVDP 53.3 24.0 38.9 43.8 52.8 24.0 38.6 43.9 52.0 23.8 38.7 43.4 51.8 23.8 38.6 43.6 51.8 23.6 38.2 43.0 39.5 +3.6
+IRG 52.6 23.5 37.5 43.7 51 23.1 37.1 42.8 50.2 23.1 37 42.2 49.8 23.2 36.5 41.1 49.7 22.5 36.5 40.7 38.9 +2.9
+MemCLR 50.2 23.3 36.6 43 51.4 23.2 37.1 42.9 50.5 23.1 36.6 42.6 49.9 22.8 36.8 42 49.9 22.9 36.6 41 38.7 +2.7
+Ours-unstop 53.9 24.9 39.2 45.1 53.5 25.1 39 45.3 53.3 24.6 39.3 45.3 53.2 24.5 39.3 44.9 52.5 24.3 38.8 44.7 40.3 +4.3
+Ours 54.4 25.4 38.7 46.1 53.5 25.6 38.5 46.2 54.1 25.9 38.0 45.8 53.3 25.9 37.9 45.4 53.5 25.4 37.4 45.3 40.4 +4.4
 Table 9.
-Detailed hyper-parameters for each benchmark.
-“City2City-C-Short” denotes the short-term Cityscapes-to-Cityscapes-C task, “City2City-C-Long” denotes long-term Cityscapes-to-Cityscapes-C task, and “ City2ACDC-Long” denotes long-term Cityscapes-to-ACDC task.
-Hyper-parameter Description City2City-C-Short City2City-C-Long SHIFT-Short City2ACDC-Long
-$\gamma$ Learning rate 0.001 0.001 0.0001 0.001
-$\alpha$ Momentum in EMA 0.9996 0.9998 0.9999 0.9998
-$\tau$ Temperature in CL 0.07 0.07 0.07 0.07
-$\delta_{0}$ Initial valuse in DT 0.8 0.8 0.8 0.8
-$\beta$ Hyper-parameter in DT 0.95 0.95 0.95 0.95
-$\epsilon$ Hyper-parameter in DT 1.3 1.3 1.3 1.3
-$\delta_{max}$ The upper-bound in DT 0.9 0.9 0.9 0.9
-$\delta_{mini}$ The lower-bound in DT 0.7 0.7 0.7 0.7
-$q$ Quantile value 0.05 0.001 0.001 0.01
+Complete experimental results (mAP@0.5) of Cityscapes-to-ACDC long-term CTTA task.
+We evaluate the long-term adaptation performance by continually adapting the source model to the four conditions ten times with the largest corruption severity level 5.
 Algorithm 1 Pseudo-code for the CTAOD Method
 1:Initialization: Object detectors: The student and teacher model parameters $\theta_{0}^{std}$ and $\theta_{0}^{tch}$ initialized from source-trained model $\theta_{0}$.
-Hyper-parameters: learning rates $\gamma$, Momentum $\alpha$ in exponential moving average (EMA), temperature $\tau$, thresholds $\delta_{0}^{c}$, $\beta$ and $\epsilon$ in dynamic threshold, and quantile value $q$.
+Hyper-parameters: learning rates $\gamma$, Momentum $\alpha$ in exponential moving average (EMA), temperature $\tau$, thresholds $\delta_{0}^{c}$, $\beta_{t}$ and $\epsilon$ in dynamic threshold, $\beta_{s}$ and $\delta_{s}$ in dynamic skipping, and quantile value $q$.
 2:Input: Unlabeled test data $x_{t}$ for the target domain
@@
 5:       Generate predictions: $y_{t}=f_{\theta_{t-1}^{tch}}(f_{weak}(x_{t}))$ and proposal $p_{t}$ using the teacher model
-6:       // 1. Object-level Contrastive Learning
-7:       Extract teacher features $F_{t}^{T}$ and student features $F_{t}^{S}$ based on $p_{t}$ using RoIAlign.
-8:       Compute contrastive learning loss $\mathcal{L}_{cl}$:
-9:             $\mathcal{L}_{cl}(x_{t})=\frac{1}{l}\sum_{i=1}^{l}-\log\frac{\exp(f_{i}^{T}\cdot f_{i}^{S}/\tau)}{\sum_{j=1}^{l}\exp(f_{i}^{T}\cdot f_{j}^{S}/\tau)}$
-10:       // 2. Dynamic Threshold
-11:       for each category $c$ do
-12:             Compute the mean predicted scores $\overline{l_{t}^{c}}$ of category $c$
-13:             Update dynamic thresholds based on predictions:
-14:                   $\delta_{t}^{c}\leftarrow\beta\cdot\delta_{t-1}^{c}+(1-\beta)\cdot\epsilon\cdot(\overline{l_{t}^{c}})^{\frac{1}{2}}$
-15:       end for
-16:       Generate the pseudo-label $\hat{y}_{t}$ through the category-specific threshold $\delta_{0}^{c}$
-17:       Compute the supervised loss $\mathcal{L}_{pl}(x_{t})=\mathcal{L}_{rpn}(x_{t},\hat{y}_{t})+\mathcal{L}_{rcnn}(x_{t},\hat{y}_{t})$
-18:       Perform gradient descent to update the student model $\theta_{t}^{std}$ and get the gradient matrix $G_{t}$
-19:       Update teacher model $\theta_{t}^{tch}$ via EMA strategy
-20:       // 3. Data-driven Stochastic Restoration
-21:       Generate random matrix $R_{t}\sim\operatorname{Uniform}(0,1)$
-22:       Generate the fisher information matrix: $F_{t}=G_{t}\odot G_{t}\odot R_{t}$
-23:       Find the $q$-quantile of $F_{t}$: $\eta=quantile(F_{t},q)$
-24:       Generate the mask matrix: $M_{t}=F_{t}<\eta$
-25:       Reset the updated student model: $\theta_{t}^{std}=M_{t}\odot\theta_{0}^{std}+(1-M_{t})\odot\theta_{t}^{std}$
-26:end for
-27:Output: Teacher predictions $y_{t}$
-Table 9.
-Complete experimental results (mAP@0.5) of Cityscapes-to-Cityscapes-C long-term CTTA task.
-We evaluate the long-term adaptation performance by continually adapting the source model to the five corruption ten times with the largest corruption severity level 5.
-“F.”, “M.”, “S.”, “B.”, and “D.” represent Fog, Motion, Snow, Brightness, and Defocus, respectively.
-Time $t\xrightarrow{\hskip 403.26341pt}$
-Round 1 2 3 4 5 /
-Condition F. M. S. B. D. F. M. S. B. D. F. M. S. B. D. F. M. S. B. D. F. M. S. B. D. / /
-Source 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 / /
-Tent 35.8 8.1 0.2 29.2 6.2 34.6 7.7 0.1 27.5 5.5 33.5 7.8 0.1 25.5 5.1 32 7.3 0.1 23.2 4.6 30.5 6.8 0.1 21.0 3.9 / /
-CoTTA 38.2 10.6 0.4 33.8 9.3 39.9 10.7 0.5 34.0 9.9 40.6 10.4 0.6 34.4 9.8 40.9 10.4 0.6 35.4 9.1 40.8 10.4 0.6 36.5 9.2 / /
-SVDP 36.8 8.8 0.6 43.5 11.2 43.9 11.8 1.7 47.9 14.9 46.2 13.9 2.5 50.0 17.8 46.2 14.4 3.2 49.0 19.6 45.0 15.4 3.8 47.8 19.9 / /
-IRG 37.9 9.0 0.7 44.3 12.2 46.1 12.9 2.0 49.2 16.6 47.6 15.2 3.0 49.5 19.5 45.2 17.3 4.6 48.6 21.4 45.6 17.7 6.0 46.7 21.9 / /
-MemCLR 37.7 8.9 0.8 45.5 13.5 45.8 13.6 2.3 49.8 17.7 47.3 16.4 4.1 50.2 20.3 46.4 17.4 4.8 47.8 21.9 45.0 18.1 5.1 46.8 22.4 / /
-Ours 39.0 10.4 0.8 48.0 13.8 48.2 13.7 3.1 52.9 18.1 49.3 15.9 5.3 53.3 21.3 48.5 17.1 7.1 52.2 23.5 49.4 18.6 7.7 51.7 25.0 / /
-Round 6 7 8 9 10 Mean Gain
-Source 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31 6.7 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 16.4 /
-Tent 28.6 6.4 0.1 19.0 3.5 26.6 6.3 0.1 16.9 3.0 24.8 5.9 0.1 15.2 2.7 22.9 5.6 0.1 13.7 2.7 20.4 5.0 0.1 11.5 2.4 12.0 -4.4
-CoTTA 39.6 10.3 0.6 34.9 9.4 40.2 10.8 0.5 35.3 10.0 40.0 10.8 0.5 35 9.7 40.1 10.5 0.5 34.1 9.5 40.8 10.4 0.5 36.3 9.6 19.1 +2.7
-SVDP 44.4 15.6 4.2 47.6 20.5 44.1 16.1 4.3 46.8 20.4 43.0 15.9 4.6 45.7 20.7 42.6 16.0 4.8 45.0 20.2 41.5 16.2 4.6 44.6 20.2 25.3 +8.9
-IRG 43.9 18 5.9 44.8 22.1 42.3 17.5 6.6 42.9 22.2 41.7 16.9 7.2 42.1 21.8 38.8 17 7.2 40.1 20.6 37.0 16.7 7.1 38.5 20.5 25.6 +9.2
-MemCLR 43.4 19.3 5.9 44.6 23.3 41.1 19.5 6.6 43.1 23.9 39.6 19.1 7.0 40.9 24.3 38.6 18.9 7.2 39.1 22.4 36.8 18.5 7.5 38.4 21.8 26.0 +9.6
-Ours 48.5 19.7 8.9 50.7 26.3 47.8 19.6 9.8 49.0 26.1 45.7 19.3 10.5 49.4 26.9 45.2 19.7 11.2 47.4 26.9 45.7 20.0 11.8 46.4 25.8 29.0 +12.6
-Table 10.
-Complete experimental results (mAP@0.5) of Cityscapes-to-ACDC long-term CTTA task.
-We evaluate the long-term adaptation performance by continually adapting the source model to the four conditions ten times with the largest corruption severity level 5.
-Time $t\xrightarrow{\hskip 390.25534pt}$
-Round 1 2 3 4 5 /
-Condition Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow / /
-Source 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 / /
-Tent 52.4 18.6 33.4 38.9 52.4 18.1 32.5 37.7 52.2 17.8 32.7 36.7 51.7 17.4 31.4 36 50.4 16.5 31.7 32.9 / /
-CoTTA 53.7 19.7 38 42.4 53.6 20.3 37.7 42.2 53.7 20.1 38.1 43.3 53.1 19.7 37.7 42.9 52.6 19.5 37 42.2 / /
-SVDP 52.8 20.0 35.6 42.0 53.4 22.0 37.4 43.8 54.0 23.0 38.8 44.0 54.6 23.5 38.7 43.8 53.4 24.0 38.0 44.4 / /
-IRG 52.7 20.6 36 42.9 53.2 22.4 37.4 44.6 53.8 23.2 38.1 45.5 53.6 23.2 38.1 45.6 53.5 23.2 38.1 45.3 / /
-MemCLR 52.9 21.2 35.2 42.8 52.5 23.1 37.3 44.3 53.9 23.2 38.4 44.8 52.8 23.2 38.2 44.4 51.5 23.5 37.2 43.8 / /
-Ours 52.9 20.4 34.5 42.9 53 22.9 38.6 44.6 54.2 24 39.5 45.3 54.5 23.8 39.6 45.8 54.4 24.4 39.2 45.4 / /
-Round 6 7 8 9 10 Mean Gain
-Source 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 36 /
-Tent 49 15.8 29.5 30.6 45.8 14.6 28.1 28.5 42.4 12.6 25.8 24.9 38.9 11 23.6 21.6 35 9.5 21.5 18.3 30.5 -5.5
-CoTTA 52.3 19.1 36.3 42 51.3 19 36.5 41.8 51.2 18.9 36.1 41.4 51.5 19.2 36.4 41.3 50.9 19.1 36 42.4 37.8 +1.8
-SVDP 53.3 24.0 38.9 43.8 52.8 24.0 38.6 43.9 52.0 23.8 38.7 43.4 51.8 23.8 38.6 43.6 51.8 23.6 38.2 43.0 39.5 +3.6
-IRG 52.6 23.5 37.5 43.7 51 23.1 37.1 42.8 50.2 23.1 37 42.2 49.8 23.2 36.5 41.1 49.7 22.5 36.5 40.7 38.9 +2.9
-MemCLR 50.2 23.3 36.6 43 51.4 23.2 37.1 42.9 50.5 23.1 36.6 42.6 49.9 22.8 36.8 42 49.9 22.9 36.6 41 38.7 +2.7
-Ours 53.9 24.9 39.2 45.1 53.5 25.1 39 45.3 53.3 24.6 39.3 45.3 53.2 24.5 39.3 44.9 52.5 24.3 38.8 44.7 40.3 +4.3
+6:       // 1. Adaptive Monitoring
+7:       Compute the mean predicted scores of all instances $\overline{l}_{t}$ and the index $\frac{\overline{l}_{t}}{\overline{l}_{t}^{ema}}$
+8:       Udapate the exponentially moving average of mean scores $\overline{l}_{t}^{ema}$ through:
+9:             $\overline{l}_{t}^{ema}\leftarrow\beta_{s}\cdot\overline{l}_{t-1}^{ema}+(1-\beta_{s})\cdot\overline{l}_{t},$
+10:       if $\frac{\overline{l}_{t}}{\overline{l}_{t}^{ema}}$ not lies with $(\frac{1}{\delta_{s}},\delta_{s})$ then
+11:             Pause the current adaption
+12:       end if
+13:       for each category $c$ do
+14:             Compute the mean predicted scores $\overline{l_{t}^{c}}$ of category $c$
+15:             Update dynamic thresholds based on predictions:
+16:                   $\delta_{t}^{c}\leftarrow\beta_{t}\cdot\delta_{t-1}^{c}+(1-\beta_{t})\cdot\epsilon\cdot(\overline{l_{t}^{c}})^{\frac{1}{2}}$
+17:       end for
+18:       // 2. Object-level Contrastive Learning
+19:       Extract teacher features $F_{t}^{T}$ and student features $F_{t}^{S}$ based on $p_{t}$ using RoIAlign
+20:       Compute contrastive learning loss $\mathcal{L}_{cl}$:
+21:             $\mathcal{L}_{cl}(x_{t})=\frac{1}{l}\sum_{i=1}^{l}-\log\frac{\exp(f_{i}^{T}\cdot f_{i}^{S}/\tau)}{\sum_{j=1}^{l}\exp(f_{i}^{T}\cdot f_{j}^{S}/\tau)}$
+22:       Generate the pseudo-label $\hat{y}_{t}$ through the category-specific threshold $\delta_{0}^{c}$
+23:       Compute the supervised loss $\mathcal{L}_{pl}(x_{t})=\mathcal{L}_{rpn}(x_{t},\hat{y}_{t})+\mathcal{L}_{rcnn}(x_{t},\hat{y}_{t})$
+24:       Perform gradient descent to update the student model $\theta_{t}^{std}$ and get the gradient matrix $G_{t}$
+25:       Update teacher model $\theta_{t}^{tch}$ via EMA strategy
+26:       // 3. Data-driven Stochastic Restoration
+27:       Generate random matrix $R_{t}\sim\operatorname{Uniform}(0,1)$
+28:       Generate the fisher information matrix: $F_{t}=G_{t}\odot G_{t}\odot R_{t}$
+29:       Find the $q$-quantile of $F_{t}$: $\eta=quantile(F_{t},q)$
+30:       Generate the mask matrix: $M_{t}=F_{t}<\eta$
+31:       Reset the updated student model: $\theta_{t}^{std}=M_{t}\odot\theta_{0}^{std}+(1-M_{t})\odot\theta_{t}^{std}$
+32:end for
+33:Output: Teacher predictions $y_{t}$
 (a) Source
-(b) Tent
+(b) SVDP
 (c) CoTTA
@@
 (a) Source
-(b) Tent
+(b) SVDP
 (c) CoTTA
@@
 (a) Source
-(b) Tent
+(b) SVDP
 (c) CoTTA
@@
 (a) Source
-(b) Tent
+(b) SVDP
 (c) CoTTA
@@
 (a) Source
-(b) Tent
+(b) SVDP
 (c) CoTTA

### v4

@@
 Continual Test-Time Adaptation (CTTA) has recently emerged as a promising technique to gradually adapt a source-trained model to continually changing target domains.
-Despite recent advancements in addressing CTTA, two critical issues remain: 1) Fixed thresholds for pseudo-labeling in existing methodologies generate low-quality pseudo-labels, as model confidence varies across categories and domains; 2) Stochastic parameter restoration methods for mitigating catastrophic forgetting fail to effectively preserve critical information due to their intrinsic randomness.
-To tackle these challenges for detection models in CTTA scenarios, we present CTAOD, featuring three core components.
+Despite recent advancements in addressing CTTA, two critical issues remain: 1) Fixed thresholds for pseudo-labeling in existing methodologies lead to low-quality pseudo-labels, as model confidence varies across categories and domains; 2) Stochastic parameter restoration methods for mitigating catastrophic forgetting fail to preserve critical information effectively, due to their intrinsic randomness.
+To tackle these challenges for detection models in CTTA scenarios, we present AMROD, featuring three core components.
 Firstly, the object-level contrastive learning module extracts object-level features for contrastive learning to refine the feature representation in the target domain.
 Secondly, the adaptive monitoring module dynamically skips unnecessary adaptation and updates the category-specific threshold based on predicted confidence scores to enable efficiency and improve the quality of pseudo-labels.
-Lastly, the data-driven stochastic restoration mechanism selectively reset inactive parameters with higher possibilities, ensuring the retention of essential knowledge.
-We demonstrate the effectiveness of CTAOD on four CTTA object detection tasks, where CTAOD outperforms existing methods, especially achieving a 3.2 mAP improvement and a 20% increase in efficiency on the Cityscapes-to-Cityscapes-C CTTA task.
+Lastly, the adaptive randomized restoration mechanism selectively reset inactive parameters with higher possibilities, ensuring the retention of essential knowledge.
+We demonstrate the effectiveness of AMROD on four CTTA object detection tasks, where AMROD outperforms existing methods, especially achieving a 3.2 mAP improvement and a 20% increase in efficiency on the Cityscapes-to-Cityscapes-C CTTA task.
 The code will be released.
@@
 Figure 1.
-Motivation.
-The source-trained model adapts to dynamic environments without source data.
+The Motivation of our proposed method.
+The source-trained model adapts to continually changing target domains without source data.
 (1) AM: The bar chart shows mean model scores varying across categories and domains.
-Unlike existing methods that continually adapt to all test data using fixed thresholds, CTAOD dynamically skips unnecessary adaptations and updates the category-specific threshold.
-(2) DSR: The bottom illustrates the neuron reset process, where darker purple indicates higher activity, and blue represents source knowledge.
-While existing methods restore parameters randomly, CTAOD resets inactive parameters with higher possibilities.
+Unlike existing methods that continually adapt to all test data using fixed thresholds, AMROD dynamically skips unnecessary adaptations and updates the category-specific threshold.
+(2) ARR: The bottom illustrates the neuron reset process, where darker purple indicates higher activity, and blue represents source knowledge.
+While existing methods restore parameters randomly, AMROR resets inactive parameters with higher possibilities.
 Deep learning models have demonstrated immense potential across various data modalities such as image (He et al., 2016; Dosovitskiy et al., 2020), audio (Aytar et al., 2016), and language (Devlin et al., 2018).
@@
 Still, UDA falls short by necessitating access to source data, which is often inaccessible due to privacy constraints or data transmission barriers (Huang et al., 2021).
-This limitation catalyzes the exploration of Test-Time Adaptation (TTA) where the source-trained model directly adapts toward unlabeled test samples encountered during evaluation in an online manner, without the reliance on the source data (Wang et al., 2020).
+This limitation catalyzes the exploration of Test-Time Adaptation (TTA) where the source-trained model directly adapts toward unlabeled test samples encountered during evaluation in an online manner, without the reliance on the source data (Mummadi et al., 2021; Sinha et al., 2023; You et al., 2021; Wang et al., 2020).
 Nonetheless, the TTA methods, which assume a static target domain, face a more challenging and realistic problem, as real-world systems work in non-stationary environments.
 For example, a vehicle may encounter various continuous environmental changes such as fog, night, rain, and snow during its journey.
-Existing TTA methods are vulnerable to catastrophic forgetting of source knowledge and error accumulation when adaptation faces more than one distribution shift (Wang et al., 2022).
+Existing TTA methods are vulnerable to catastrophic forgetting of source knowledge and error accumulation when adaptation faces more than one distribution shift (Wang et al., 2022; Döbler et al., 2023).
 Recently, Wang et al.
@@
 Although randomly resetting parameters partially helps alleviate the forgetting of source knowledge (Wang et al., 2022; Zhu et al., 2023), its randomness may also contribute to losing crucial knowledge specific to the current domain (refer to Figure 1).
-Additionally, existing CTTA methods normally utilize pseudo-labeling through a fixed threshold for self-supervision (Wang et al., 2022; Döbler et al., 2023).
+Additionally, existing CTTA methods normally utilize pseudo-labeling through a fixed threshold for self-supervision (Wang et al., 2022; Döbler et al., 2023; Sójka et al., 2023).
 Given that model confidence may vary across categories and domains (refer to Figure 1), employing a uniformly fixed threshold could exclude high-quality pseudo-labels while incorporating incorrect ones.
@@
 Therefore, despite recent efforts to handle CTTA, two significant challenges persist: Current self-training-based methods suffer from low-quality pseudo-labels, leading to error accumulation (Challenge 1); Continual adaptation to dynamic environments struggles to effectively retain valuable knowledge about the current domain while mitigating the forgetting of source knowledge (Challenge 2).
-To tackle these challenges, we introduce CTAOD (Continual Test-time Adaption for Object Detection).
-Aligning with previous CTTA works for robust adaptation (Wang et al., 2022), CTAOD is constructed upon the mean-teacher framework (Tarvainen and Valpola, 2017) where the teacher model is an exponential moving average of the student model.
-Particularly, CTAOD comprises Object-level Contrastive Learning (OCL), Adaptive Monitoring (AM), and Data-driven Stochastic Restoration (DSR).
-More specifically, OCL extracts features based on proposals generated by the Region Proposal Network (RPN), which provides multiple cropped views around the object at different locations and scales.
+To tackle these challenges, we introduce AMROD (Adaptive Monitoring and Restoration for Object Detection).
+Aligning with previous CTTA works for robust adaptation (Wang et al., 2022; Gong et al., 2022; Döbler et al., 2023; Niloy et al., 2024), AMROR is constructed upon the mean-teacher framework (Tarvainen and Valpola, 2017) where the teacher model is an exponential moving average of the student model.
+Particularly, AMROR comprises Object-level Contrastive Learning (OCL), Adaptive Monitoring (AM), and Adaptive Randomized Restoration (ARR) modules.
+More specifically, OCL extracts object-level features based on proposals generated by the Region Proposal Network (RPN), which provides multiple cropped views around the object at different locations and scales.
 Subsequently, Contrastive Learning (CL) loss is applied on RPN cropped views to guide the model to encourage similar object instances to remain close while pushing dissimilar ones apart.
 The OCL is well integrated into the mean-teacher paradigm as a drop-in enhancement to acquire more fine-grained and localized feature representation.
-Furthermore, to make adaptation more efficient and stable and to improve the quality of pseudo-labels, i.e., addressing Challenge 1, we design an AM module to decide whether to pause or resume the adaptation and dynamically adjust the category-specific thresholds, based on the predicted confidence scores.
+Furthermore, to make adaptation more efficient and stable and to improve the quality of pseudo-labels, i.e., addressing Challenge 1, we design an AM module to decide whether to pause or resume the adaptation and dynamically adjust the category-specific thresholds, based on the mean predicted confidence scores.
 The dynamic nature of the AM method makes it better suited to address the effects of continuously changing distributions.
-Finally, the DSR mechanism is proposed to reset inactive parameters with a higher possibility than active ones by employing the gradients of parameters as weights for a random mask matrix.
-DSR not only helps prevent forgetting but preserves important information, i.e., addressing Challenge 2.
-On the other hand, its randomness also allows potentially falsely activated parameters to be reset, thereby making the model more stable.
+Finally, the ARR mechanism is proposed to reset inactive parameters with a higher possibility than active ones by employing the gradients of parameters as weights for a random mask matrix.
+ARR not only helps prevent forgetting but preserves important information, i.e., addressing Challenge 2.
+On the other hand, its randomness also allows certain falsely activated parameters to be reset as well, thereby resulting in greater stability for adaptation.
 We demonstrate the effectiveness of our proposed approach on four continual test-time adaptation benchmark tasks for object detection.
@@
 Our main contributions are summarized as follows:
-  • This study introduces CTAOD, which pioneers exploring CTTA for detection models.
+  • This study introduces AMROR, which pioneers exploring CTTA for detection models.
 Specifically, we propose to leverage object-level features for contrastive learning to refine feature representation in CTTA object detection.
-  • To address the two challenges in CTTA, our proposed CTAOD enables dynamic skipping and category-specific threshold updates based on predicted scores and reset the inactive parameter with higher possibilities.
-  • Empirical experiments demonstrate that CTAOD surpasses existing methods and facilitates efficient short-term and long-term adaptation in dynamic environments.
+  • To address the two challenges in CTTA, our proposed AMROR enables dynamic skipping and category-specific threshold updates based on the mean predicted scores and reset the inactive parameter with higher possibilities.
+  • Empirical experiments demonstrate that AMROR surpasses existing methods and facilitates efficient short-term and long-term adaptation in dynamic environments.
 2. RELATED WORK
 2.1. Source-Free Domain Adaptation
-UDA tackles the inter-domain divergence by aligning the distributions of source and target data (Ganin et al., 2016; Luo et al., 2020).
+UDA tackles the inter-domain divergence by aligning the distributions of source and target data (Ganin et al., 2016; Tzeng et al., 2017; Saito et al., 2018, 2019; Li et al., 2022a; Cao et al., 2023).
 Despite its effectiveness, the limitation of UDA lies in its requirement for access to the source domain data, which often raises concerns regarding data privacy and transmission efficiency.
-As a result, Source-Free Domain Adaptation (SFDA) received extensive research attention in object detection, where the source-trained detector is re-trained on the target training data (VS et al., 2023b, a; Li et al., 2021) before evaluation.
-For instance, MemCLR (VS et al., 2023a) employs a cross-attention-based memory bank for CL, while IRG (VS et al., 2023a) utilizes the object relations with instance relation graph network to explore the SFDA setting for object detection without the source data.
+As a result, Source-Free Domain Adaptation (SFDA) received extensive research attention in object detection, where the source-trained detector is adapted to the target training data without any source data (VS et al., 2023b, a; Li et al., 2022b, 2021; Chen et al., 2023) before evaluation.
+For instance, MemCLR (VS et al., 2023a) employs a cross-attention-based memory bank with CL for source-free detectors, while IRG (VS et al., 2023a) utilizes the object relations with instance relation graph network to explore the SFDA setting for object detection.
 However, the standard SFDA setting requires prior knowledge of the target domain, which is impractical in most real-world applications.
@@
 In this paper, we distinguish between TTA and SFDA based on evaluation protocol, although they can be transformed into each other in experiments.
-Furthermore, TTA methods improve the model performance under distribution shift commonly through pseudo-labeling (Sinha et al., 2023; Iwasawa and Matsuo, 2021), batchnorm statistics updating (Hu et al., 2021; You et al., 2021), or entropy regularization (Wang et al., 2020; Iwasawa and Matsuo, 2021) during testing.
+Furthermore, TTA methods improve the model performance under distribution shift commonly through pseudo-labeling (Mummadi et al., 2021; Sinha et al., 2023; Iwasawa and Matsuo, 2021; Sun et al., 2020), batchnorm statistics updating (Hu et al., 2021; You et al., 2021), or entropy regularization (Wang et al., 2020; Zhou and Levine, 2021; Iwasawa and Matsuo, 2021; Niu et al., 2022) during testing.
 For example, Tent (Wang et al., 2020) updates the batchnorm parameters with entropy minimization and demands a large batch size for optimization during test-time adaptation, which is unsuitable for real-time detection model deployment where images are processed sequentially.
-TTA methods assume a static target domain where the target data come from a single domain.
-However, the target distribution may exhibit a continual shift over time in practical scenarios.
+The above approaches assume a static target domain where the target data come from a single domain.
+However, in numerous practical scenarios, the distribution of target domain may exhibit a continual shift over time.
 2.3. Continual Test-time Adaptation
+Early works consider adaptation to continually changing domains by aligning the source and target data distribution (Hoffman et al., 2014; Wulfmeier et al., 2018) and the reliance on source data limits their applicability.
 Recently, Wang et al.
 (2022) introduce CoTTA, marking the first work tailored to the demands of CTTA by adapting a pre-trained model to sequences of domains without source data.
-Subsequently, research efforts have been dedicated to exploring CTTA, primarily focused on classification (Döbler et al., 2023; Brahma and Rai, 2023; gan2023decorate) and segmentation (Song et al., 2023; Ni et al., 2023; Liu et al., 2023) tasks.
+Subsequently, research efforts have been dedicated to exploring CTTA, primarily focused on classification (Döbler et al., 2023, 2023; Niloy et al., 2024; Brahma and Rai, 2023; Gan et al., 2023a) and segmentation (Song et al., 2023; Ni et al., 2023; Liu et al., 2023; Niloy et al., 2024; Zhu et al., 2023) tasks.
 For instance, similar to CoTTA, Zhu et al.
 (2023) apply a stochastic reset mechanism to prevent forgetting in the CTTA medical segmentation task.
-In contrast, PETAL (Brahma and Rai, 2023) utilizes the Fisher Information Matrix (FIM) as a metric of parameter importance to reset only the most irrelevant parameters.
+In contrast, PETAL (Brahma and Rai, 2023) utilizes the Fisher Information Matrix (FIM) as a metric of parameter importance to reset only the most irrelevant parameters across all layers.
 Nevertheless, these two methods are sub-optimal, as the randomness might lead to losing essential information, and pure data-driven restoration may retain false active parameters resulting from noise.
-Moreover, the mean-teacher framework (Tarvainen and Valpola, 2017) serves as a base architecture for most CTTA works (Wang et al., 2022; Gong et al., 2022), where the teacher generates pseudo-labels via a fixed threshold to supervise the student.
+Moreover, the mean-teacher framework (Tarvainen and Valpola, 2017) serves as a base architecture for most CTTA works (Wang et al., 2022; Gong et al., 2022; Döbler et al., 2023; Niloy et al., 2024), where the teacher model generates pseudo-labels via a fixed threshold to supervise the training of the student model.
 Nonetheless, these methods suffer from low-quality pseudo-labels with a uniform threshold since the model confidence varies across categories and domains.
@@
 However, this approach is unsuitable for object detection where smaller batch sizes are preferred for computational efficiency.
-To improve efficiency, yoo2024and presents a CTTA approach for object detection to update specific parts of the model in specific adaptation, which utilizes source statistics.
+To improve efficiency, Yoo et al.
+(2024) presents a CTTA approach for object detection to update specific parts of the model in specific adaptation, which utilizes source statistics.
 Although it effectively improved efficiency, its performance is limited.
 Additionally, Gan et al.
-(2023) presents a cloud-device collaborative continual adaptation paradigm in object detection by aligning source and target distribution.
+(2023b) presents a cloud-device collaborative continual adaptation paradigm in object detection by aligning source and target distribution.
 Nevertheless, the reliance on source data in both methods restricts their practical applicability.
-Therefore, a gap still exists in exploring the CTTA in object detection to improve the quality of pseudo-label (Challenge 1) and effectively reset noisy neurons (Challenge 2), without relying on any source data.
+Therefore, a gap still exists in exploring the CTTA in object detection to improve the quality of pseudo-label (Challenge 1) and effectively reset noisy neurons (Challenge 2), without relying on the source data.
 3. METHOD
@@
 Figure 2.
-Overview.
-CTAOD follows the mean-teacher framework, featuring: (1) The OCL module compares region of interest features extracted from feature maps of both networks based on teacher proposals for contrastive learning; (2) The AM module dynamically skips unnecessary adaptation and adjusts category-specific thresholds based on prediction scores; (3) The DSR reset the inactive parameter with higher possibilities based on the FIM.
+The overview of the proposed method.
+AMROD follows the mean-teacher framework, featuring: (1) The OCL module compares region of interest features extracted from feature maps of both networks based on teacher proposals for contrastive learning; (2) The AM module dynamically skips unnecessary adaptation and adjusts category-specific thresholds based on the mean prediction scores; (3) The ARR strategy reset the inactive parameter with higher possibilities based on the FIM.
 3.1.2. Mean-Teacher Framework
-In alignment with previous CTTA works (Wang et al., 2022; Gong et al., 2022), we build our method based on the mean-teacher framework (Tarvainen and Valpola, 2017).
+Following previous CTTA works (Wang et al., 2022; Gong et al., 2022), we build AMROD based on the mean-teacher framework (Tarvainen and Valpola, 2017).
 Specifically, both networks are first initialized with the source-trained model.
 The teacher model produces the pseudo-labels for the unlabeled target data to supervise the student model.
-While the parameters of the student model are optimized via gradient descent, the teacher model is updated following an Exponential Moving Average (EMA) strategy based on the student model.
+While the parameters of the student are optimized via gradient descent, the teacher is updated following an Exponential Moving Average (EMA) strategy based on the student.
 Formally, this process can be expressed as follows:
@@
 However, this framework encounters error accumulation and catastrophic forgetting in dynamic environments.
-Therefore, we design CTAOD to fit in the CTTA setting.
+Therefore, we design AMROD to fit in the CTTA setting.
 3.2. Proposed Method
 3.2.1. Overview
-Figure 2 presents a concise overview of CTAOD, consisting of Object-level Contrastive Learning (OCL), Adaptive Monitoring (AM), and Data-driven Stochastic Restoration (DSR).
-Inspired by UDA object detection methods (Li et al., 2022; Liu et al., 2021), we adopt the Weak-Strong augmentation to enable the teacher model to generate reliable pseudo-labels without being affected by heavy augmentation.
-Specifically, the teacher model receives input with weak augmentations while we input student networks with strong enhanced images.
-The overall loss of the student model for gradient descent is defined as:
+Figure 2 presents a concise overview of AMROD, consisting of Object-level Contrastive Learning (OCL), Adaptive Monitoring (AM), and Adaptive Randomized Restoration (ARR).
+Inspired by UDA object detection methods (Li et al., 2022a; Liu et al., 2021), we adopt the Weak-Strong augmentation to enable the teacher model to generate reliable pseudo-labels without being affected by heavy augmentation.
+Specifically, the teacher receives input with weak augmentations while we input student networks with strong enhanced images.
+The overall loss of the student for gradient descent is defined as:
 (4) $\mathcal{L}_{all}=\mathcal{L}_{pl}(x_{t})+\lambda\mathcal{L}_{cl}(x_{t})+\mu\mathcal{L}_{kl}(x_{t}),$
@@
 The $\mathcal{L}_{kl}$ represents the Kullback-Leibler (KL) Divergence (Kullback and Leibler, 1951) loss, used to quantify the distinction between two probability distributions.
-The parameters $\lambda$ and $\mu$ are corresponding weights of the losses.
-The KL loss is defined as:
+The parameters $\lambda$ and $\mu$ are corresponding weights.
+KL loss is defined as:
 (5) $\mathcal{L}_{kl}\left(P\parallel Q\right)=\sum_{x\in\mathcal{X}}P(x)\mathrm{~{}}\log\left(\frac{P(x)}{Q(x)}\right),$
@@
 SimCLR is a widely used CL approach for self-supervised visual representation learning (Chen et al., 2020).
-SimCLR learns high-quality feature representations from differently augmented views of the same image.
-Designed for classification tasks, it requires large batch sizes to ensure sufficient positive and negative pairs.
-Consequently, SimCLR is not well-suited for object detection, where images contain multiple instances and large batch sizes demand significant computational resources.
-Motivated by SimCLR (Chen et al., 2020), we present an OCL module to extract teacher and student features for CL based on proposals generated from the RPN.
-This strategy provides multiple cropped views around the object instance, eliminating the need for large batch sizes and ensuring computational efficiency for online detection model updates.
+It learns high-quality feature representations from differently augmented views of the same image, which requires large batch sizes to ensure sufficient positive and negative pairs for classification tasks.
+Consequently, SimCLR is less suitable for object detection, where images contain multiple instances and large batch sizes demand significant computational resources.
+Motivated by SimCLR (Chen et al., 2020), we present an OCL module to extract teacher and student features for CL based on proposals from the RPN.
+This strategy provides multiple cropped views around the object instance, eliminating the need for large batch sizes and ensuring computational efficiency for online detector updates.
 Specifically, given a weakly augmented image $f_{weak}(x_{t})$ at time step t, the teacher produces ROI proposals $p_{t}=\{p_{1},...,p_{l}\}$ via region proposal network.
-We then apply RoIAlign (He et al., 2017), a pooling operation for ROI, to extract corresponding teacher and student object-level features $F_{t}^{T}=\{f_{i}^{T}\in\mathbb{R}^{1\times C}\}_{i=1}^{l}$ and $F_{t}^{S}=\{f_{i}^{S}\in\mathbb{R}^{1\times C}\}_{i=1}^{l}$ based on the feature map from the backbone, respectively.
+OCL then apply RoIAlign (He et al., 2017) to extract corresponding teacher and student object-level features $F_{t}^{T}=\{f_{i}^{T}\in\mathbb{R}^{1\times C}\}_{i=1}^{l}$ and $F_{t}^{S}=\{f_{i}^{S}\in\mathbb{R}^{1\times C}\}_{i=1}^{l}$ based on the feature map from the backbone, respectively.
 The features associated with the same proposal are considered positive pairs, otherwise negative pairs.
-we employ the CL loss to these features $F_{t}^{T}$ and $F_{t}^{S}$ derived from weak and strong enhanced images respectively by minimizing:
+Then CL loss is applied to these features $F_{t}^{T}$ and $F_{t}^{S}$ derived from weak and strong enhanced images respectively by minimizing:
 (6) $\mathcal{L}_{cl}(x_{t})=\frac{1}{l}\sum_{i=1}^{l}-\log\frac{\exp(f_{i}^{T}\cdot f_{i}^{S}/\tau)}{\sum_{j=1}^{l}\exp(f_{i}^{T}\cdot f_{j}^{S}/\tau)},$
 where $\tau>0$ is the temperature, $l$ indicates the number of the features, and $f_{i}^{T}$ and $f_{i}^{S}$ denote the features of two different augmentations of the same object, serving as the positive pair.
-This strategy encourages the model to learn fine-grained and localized representations on the target domain, without relying on accurate pseudo-labels.
+This strategy encourages the model to learn fine-grained and localized feature representations on the target domain, without relying on accurate pseudo-labels.
 3.2.3. Adaptive Monitoring
 The AM module monitors the model’s status using the predicted scores, featuring dynamic skipping and dynamic threshold.
-The former halts unnecessary adaptations to save computational resources, while the latter updates the category-specific threshold.
+The former halts unnecessary adaptations to save computational resources, while the latter updates the category-specific threshold dynamically.
 As shown in Figure 2, the teacher network first make predictions $\hat{y}_{t}=f_{\theta_{t-1}}(f_{wek}(x_{t}))$ for the weakly enhanced images.
-Subsequently, We compute the mean scores of all category $\overline{l}_{t}$ in teacher’s predictions and maintain its exponentially moving average $\overline{l}_{t}^{ema}$ through:
+Subsequently, AM computes the overall mean scores across all predicted instances $\overline{l}_{t}$ in teacher’s predictions and maintains its exponentially moving average $\overline{l}_{t}^{ema}$ through:
 (7) $\overline{l}_{t}^{ema}\leftarrow\beta_{s}\cdot\overline{l}_{t-1}^{ema}+(1-\beta_{s})\cdot\overline{l}_{t},$
@@
 This dynamic skipping strategy effectively prevents unstable and noisy adaptation, thus conserving computational resources and mitigating error accumulation.
-For dynamic threshold, the thresholds for each category are initialized with the same value $\delta_{0}$ at first, which is updated every iteration by:
+For dynamic threshold, the thresholds for each category are initialized with the same value $\delta_{0}$ at first, which will be updated every iteration by:
 (8) $\delta_{t}^{c}\leftarrow\beta_{t}\cdot\delta_{t-1}^{c}+(1-\beta_{t})\cdot\epsilon\cdot(\overline{l}_{t}^{c})^{\frac{1}{2}},$
@@
 For unlabeled data from dynamic environments, this mechanism effectively generates an appropriate threshold for each category.
-3.2.4. Data-Driven Stochastic Restoration
-Existing methods address catastrophic forgetting through stochastic reset mechanisms (Wang et al., 2022) or pure data-driven restoration (Brahma and Rai, 2023).
+3.2.4. Adaptive Randomized Restoration
+Existing methods address catastrophic forgetting through stochastic reset (Wang et al., 2022) or pure data-driven restoration (Brahma and Rai, 2023).
 Nevertheless, the former may reset valuable parameters, potentially erasing essential knowledge relevant to the current domain.
 On the other hand, the latter restores the most unimportant parameters, which may retain noise parameters, thereby leading to error accumulation.
-Therefore, we propose the DSR strategy to reset irrelevant parameters with higher possibilities, enabling robust adaptation.
+Unlike PETAL (Brahma and Rai, 2023), which was designed for classification tasks and resets the least active parameters across all layers, our approach refines this for object detection by identifying inactive parameters within each layer individually.
+This layer-specific technique accounts for the varying parameter scales at different depths of the model, and by further incorporating randomness, we propose ARR to reset irrelevant parameters with higher possibilities to enhance stability and robustness in restoration.
 Following (Brahma and Rai, 2023), we adopt the diagonal approximation of Fisher Information Matrix (FIM) (Kirkpatrick et al., 2017) based on gradient, as a metric of parameter importance.
@@
 Parameters with higher values in the FIM are those for which small changes result in significant increases in loss, indicating their importance for the domain at hand.
-Specifically, consider a convolution layer with weight $W_{t-1}$ and its corresponding gradient matrix after gradient descent $G_{t}$ with the same shape as $W_{t-1}$ at time step $t$.
-First, we use a random matrix $R_{t}\sim\operatorname{Uniform}(0,1)$ with the same shape as the weights matrix, where each element follows a uniform distribution between 0 and 1.
-Moreover, we utilize the element-wise multiplication of the square of the gradient and the random matrix to approximate the FIM: $F_{t}=G_{t}\odot G_{t}\odot R_{t}$, where $\odot$ represents the element-wise multiplication.
-Furthermore, the proposed DSR updates the parameters based on the FIM values, which is formulated as:
+Specifically, given a layer with weight $W_{t}$ updated through gradient descent to $\hat{W}_{t}$ based on gradient matrix $G_{t}$ at time step $t$, ARR generates a random matrix $R_{t}\sim\operatorname{Uniform}(0,1)$ with the same shape as the weights matrix, where each element follows a uniform distribution between 0 and 1.
+Then the element-wise multiplication of the squared gradient matrix and the random matrix is applied to approximate the FIM.
+Furthermore, ARR updates the parameters based on the FIM values, formulated as:
+(9) $F_{t}=G_{t}\odot G_{t}\odot R_{t},$
 (10) $M_{t}=F_{t}<\eta,$
-(10) $W_{t}=M_{t}\odot W_{0}+(1-M_{t})\odot W_{t-1},$
-where $M_{t}$ indicates the mask matrix of the same shape as $W_{t}$, $<$ represents the element-wise less than operation, and $\eta$ is the threshold value which is acquired by the $q$-quantile of $F_{t}$, i.e.
-$\eta=quantile(F_{t},q)$.
+(11) $W_{t+1}=M_{t}\odot W_{0}+(1-M_{t})\odot\hat{W}_{t},$
+where $\odot$ is the element-wise multiplication, $M_{t}$ indicates the mask matrix, $<$ represents element-wise less than operation, and $\eta$ is the threshold value acquired by the $q$-quantile of $F_{t}$: $\eta=quantile(F_{t},q)$.
 Consequently, the elements in $M_{t}$ are set to 1, when the corresponding FIM value is less than $\eta$, indicating the corresponding parameters should be reset to the source weight $W_{0}$.
@@
 We evaluate the performance by continually adapting the source model to the twelve corruptions.
-“-unstop” denotes no skipping.
+“-unstop” denotes AMROD without skipping.
 Time $t\xrightarrow{\hskip 368.57964pt}$
@@
 We evaluate the performance by continually adapting the source model to the five corruption ten times.
-To save space, we display selected rounds of results, and full results are provided in the appendix.
-“-unstop” denotes no skipping.
+“-unstop” denotes AMROD without skipping.
+To save space, we display selected rounds of results, and full results are provided in the supplementary material.
 In this study, we rigorously evaluate our methodology across four benchmark tasks in CTTA object detection.
@@
 Inspired by the foundational work CoTTA (Wang et al., 2022), the short-term CTTA task entails the sequential adaptation to various target domains once.
-In contrast, the long-term CTTA task involves continually adapting the model toward groups of target domains cyclically.
+In contrast, the long-term CTTA task involves continually adapting the model toward a group of target domains cyclically.
 4.1. Datasets
@@
 Our short-term CTTA task selectively focuses on the latter 12 corruptions.
-For the long-term CTTA task, we prioritize the five corruptions related to autonomous driving scenarios as the target domain groups following (Gan et al., 2023), namely Fog, Motion, Snow, Brightness, and Defocus.
+For the long-term CTTA task, we prioritize the five corruptions related to autonomous driving scenarios as the target domain group following (Gan et al., 2023b), namely Fog, Motion, Snow, Brightness, and Defocus.
 We repeat the adaptation to the target domain group 10 times to evaluate long-term performance.
@@
 ACDC.
-The ACDC dataset (Sakaridis et al., 2021) shares the same class types as Cityscapes and is collected in four different adverse visual conditions, including Fog, Night, Rain, and Snow.
+The ACDC (Sakaridis et al., 2021) shares the same class types as Cityscapes and is collected in four different adverse visual conditions, including Fog, Night, Rain, and Snow.
 Following (Wang et al., 2022), we use these four conditions as the target domain group for the long-term CTTA task, with 400 unlabeled images per condition.
-Similarly, the source-trained model is continually adapted to the target domain group for 10 cycles.
+Similarly, the source model is continually adapted to the target domain group for 10 cycles.
 4.2. Implementation Details
 We adopt the Faster R-CNN with ResNet50 (He et al., 2016) pre-trained on ImageNet (Krizhevsky et al., 2012) as the backbone.
-Aligning with (VS et al., 2023b; Wang et al., 2022), we maintain a batch size of 1, thereby emulating a real-world application scenario where the detection model adapts toward a continuous influx of images.
+Aligning with (VS et al., 2023b; Wang et al., 2022), we maintain a batch size of 1 to emulate a real-world application scenario where the detection model adapts toward a continuous influx of images.
 The source models are trained using an SGD optimizer with a learning rate of 0.001 and a momentum of 0.9.
 Algorithms are implemented leveraging the Detectron2 platform (Wu et al., 2019).
-The metric of mAP at an IoU threshold of 0.5 is employed for evaluation.
+The metric of mAP at an IoU threshold of 0.5 (mAP@0.5) is employed for evaluation.
 Each experiment is conducted on 1 NVIDIA A800 GPU.
-More hyper-parameters and implementation details are provided in the appendix.
+More hyper-parameters and implementation details are provided in the supplementary material.
 4.3. Baselines and Compared Approaches
-We compare CTAOD with six baselines across various settings, including Source (Ren et al., 2015), IRG (VS et al., 2023a), Memclr (VS et al., 2023b), CoTTA (Wang et al., 2022), SVDP (Yang et al., 2024), and CTAOD-upstop.
+We compare AMROD with six baselines across various settings, including Source (Ren et al., 2015), IRG (VS et al., 2023a), Memclr (VS et al., 2023b), CoTTA (Wang et al., 2022), SVDP (Yang et al., 2024), and AMROD-upstop.
 Specifically, ”Source” represents the source model without adaptation.
@@
 MemCLR integrates cross-attention with CL, while IRG incorporates instance relation graph and CL.
-CoTTA utilizes weight-averaged predictions and random neurons recovery to tackle CTTA, while SVDP explores sparse visual prompts for CTTA dense prediction.
-“CTAOD-unstop” represents CTAOD without dynamic skipping.
+CoTTA utilizes a fixed threshold for pseudo-labeling and random neuron recovery to tackle CTTA, while SVDP explores sparse visual prompts for CTTA dense prediction.
+“AMROD-unstop” represents AMROD without dynamic skipping.
 4.4. Results and Analysis
+4.4.1. Synthetic Continual Distribution Shift
+We first evaluate the effectiveness of AMROD on the short-term Cityscapes-to-Cityscapes-C adaptation tasks.
+As depicted in Table 1, CoTTA, IRG, and Memclr exhibit enhancements in performance, achieving 16.0, 17.5, and 18.0 mAP respectively.
+Although the SFDA methods like IRG and MemCLR assume a static target domain, employing a large momentum update rate $\alpha$ for the teacher enables a relatively stable adaptation in the short term.
+Consequently, CL leads to a better performance than CoTTA which solely employs weight-averaged predictions.
+Moreover, SVDP, which utilizes visual prompts, achieves a sub-optimal performance of 19.1 mAP.
+Remarkably, our proposed method improves the performance to 20.8 mAP, consistently outperforming the above approaches.
+Additionally, the dynamic skipping strategy in AM reduces adaptation iterations by 20% while achieving a 0.2 mAP improvement compared to AMROD-unstop.
+AMROD ensures a more reliable and efficient adaptation to target domains characterized by intense and continual changes in the short term.
 Time $t\xrightarrow{\hskip 368.57964pt}$
@@
 We evaluate the performance by continually adapting the source model to the four conditions ten times.
-To save space, we display selected rounds of results, and full results are provided in the appendix.
-“-unstop” denotes no skipping.
+“AMROD-unstop” represents AMROD without dynamic skipping.
+To save space, we display selected rounds of results, and full results are provided in the supplementary material.
 (a) Source
 (b) MemCLR
-(c) CTAOD (ours)
+(c) AMROD (ours)
 Figure 3.
 Qualitative results.
-We compare the detection results of the Source model, MemCLR, and CTAOD in the $10^{th}$ round of adaption to defocus corruption on the long-term Cityscapes-to-Cityscapes-C task.
-4.4.1. Synthetic Continual Distribution Shift
-We first evaluate the effectiveness of CTAOD on the short-term Cityscapes-to-Cityscapes-C adaptation tasks.
-As depicted in Table 1, CoTTA, IRG, and Memclr exhibit enhancements in performance, achieving 16.0, 17.5, and 18.0 mAP respectively.
-Although the SFDA methods like IRG and MemCLR assume a static target domain, employing a large momentum update rate $\alpha$ for the teacher enables a relatively stable adaptation in the short term.
-Consequently, CL leads to a better performance than CoTTA which solely employs weight-averaged predictions.
-Moreover, SVDP, which utilizes visual prompts, achieves a sub-optimal performance of 19.1 mAP.
-Remarkably, our proposed method improves the performance to 20.8 mAP, consistently outperforming the above approaches.
-Additionally, the dynamic skipping strategy in AM reduces adaptation iterations by 20% while achieving a 0.2 mAP improvement compared to CTAOD-unstop.
-CTAOD ensures a more reliable and efficient adaptation to target domains characterized by intense and continual changes in the short term.
+We compare the detection results of the Source model, MemCLR, and AMROD in the $10^{th}$ round of adaption to defocus corruption on the long-term Cityscapes-to-Cityscapes-C task.
 As presented in Table 8, the long-term task outcomes reveal the source model’s poor performance, with an average mAP of 16.4.
@@
 In contrast, SVDP which employs sparse visual prompts raises the performance to 25.3 mAP.
-Particularly, CTAOD yields a remarkable 12.8 mAP enhancement over the Source and surpasses all comparative baselines using 80% adaptation iterations, which employs the DSR mechanism to conserve valuable knowledge while eliminating noise from prior domains.
+Particularly, AMROD yields a remarkable 12.8 mAP enhancement over the Source and surpasses all comparative baselines using 80% adaptation iterations, which employs the ARR mechanism to conserve valuable knowledge while eliminating noise from prior domains.
 These findings empirically validate the effectiveness of our proposed method in ensuring stable and efficient adaptation amidst synthetic continual distribution shifts, across both short-term and long-term scenarios.
@@
 The results of the SHIFT short-term CTTA task are shown in Table 4.
-While other baseline methods suffer from performance decline, SVDP which employs a fixed threshold for pseudo-labeling and prompts learning achieves a 0.5 mAP improvement.
-Moreover, CTAOD which introduces a dynamic threshold strategy in AM shows a superior performance of 43.9 mAP with 16% improvement in efficiency.
-The experimental results of ACDC long-term CTTA task are summarized in Table 3.
+While other baseline methods suffer from performance decline, SVDP which employs a fixed threshold for pseudo-labeling, stochastic restoration, and prompts learning achieves a 0.5 mAP improvement.
+Moreover, AMROD which introduces AM and ARR modules shows a superior performance of 43.9 mAP with 16% improvement in efficiency.
+The experimental results of the ACDC long-term CTTA task are summarized in Table 3.
 CoTTA, IRG, and MemCLR potentially suffer from error accumulation and catastrophic forgetting, manifesting in a rapid decline in performance during the later stages.
 Similarly, SVDP achieves a sub-optimal performance of 39.5 mAP.
-Notably, CTAOD achieves a performance of 40.4 mAP, yielding an absolute improvement of 0.9 mAP over baselines.
-While CTAOD may not display a significant advantage over other methods in the initial round, its performance continually improves and remains stable throughout the long-term adaptation.
+Notably, AMROD achieves a performance of 40.4 mAP, yielding an absolute improvement of 0.9 mAP over baselines.
+While AMROD may not display a significant advantage over other methods in the initial round, its performance continually improves and remains stable throughout the long-term adaptation.
 These results underscore the capability of our method to foster robust adaptation toward the real-world continual distribution shift in both the short and long term.
-4.4.3. Qualitative Results
-As shown in Figure 3, we selectively compare the source model, MemCLR, and CTAOD in the final round adaptation to the brightness on the long-term Cityscapes-to-Cityscapes-C tasks.
-CTAOD guides the model in learning better feature representations while effectively mitigating forgetting in the long-term adaptation.
-Therefore, CTAOD assists the detector in distinguishing more foreground object categories and better locating them.
-More visualization analysis is presented in the appendix.
-Mean-Teacher OCL AM DSR Mean Gain
+Mean-Teacher OCL AM ARR Mean Gain
 1 $\checkmark$ 25.9 /
@@
 Table 5.
-Alation experiment.
+Main ablation experiment.
 “Mean-Teacher” represents the base mean teacher framework with weak-strong augmentation and KL divergence distillation.
 All experiments are done on long-term Cityscapes-to-Cityscapes-C tasks.
+4.4.3. Qualitative Results
+As shown in Figure 3, we selectively compare the source model, MemCLR, and AMROD in the final round adaptation to the brightness on the long-term Cityscapes-to-Cityscapes-C tasks.
+AMROD guides the model in learning better feature representations while effectively mitigating forgetting in the long-term adaptation.
+Therefore, AMROD assists the detector in distinguishing more foreground object categories and better locating them.
+More visualization analysis is presented in the supplementary material.
+AMROD FT of 0.9 FT of 0.8 FT of 0.7 SR DR Mean
+1 $\checkmark$ $\checkmark$ 27.7
+2 $\checkmark$ $\checkmark$ 28.2
+3 $\checkmark$ $\checkmark$ 28.5
+4 $\checkmark$ $\checkmark$ 28.7
+5 $\checkmark$ $\checkmark$ 28.3
+6 $\checkmark$ 29.2
+Table 6.
+Fine-grained ablation study.
+The “FT” represents replacing AM module in AMROD with a fixed threshold.
+The “SR” and the “DR” indicate replacing ARR with the stochastic restoration (Wang et al., 2022) and the data-driven restoration (Brahma and Rai, 2023), respectively.
+All experiments are done on long-term Cityscapes-to-Cityscapes-C tasks.
 4.4.4. Ablation Study
-We conduct an ablation study to empirically assess the impact of the OCL, AM, and DSR.
-As shown in Table 5, the integration of the mean-teacher architecture elevates the performance to 25.9 mAP.
-Building upon the mean-teacher foundation, the proposed OCL, AM, and DSR achieve an additional 1.5, 1.4, and 0.8 mAP improvement individually.
+We conduct an ablation study to evaluate the impact of the OCL, AM, and ARR.
+As shown in Table 5, integrating the mean-teacher architecture achieves the performance of 25.9 mAP.
+Building upon this foundation, the proposed OCL, AM, and ARR achieve an additional 1.5, 1.4, and 0.8 mAP improvement individually.
 Moreover, the combination of the OCL and AM modules raises the performance to 28.6 mAP.
-Furthermore, the inclusion of the DSR mechanism brings a 0.6 mAP boost in performance, culminating in a comprehensive model performance of 29.2 mAP.
-More fine-grained comparisons with AM and DSR are provided in the appendix.
+Furthermore, the inclusion of the ARR mechanism brings a 0.6 mAP boost in performance, culminating in a comprehensive model performance of 29.2 mAP.
+As presented in Table 6, we further perform a fine-grained comparison study to assess the effectiveness of the AM and ARR modules.
+This is done by replacing AM and ARR with a Fixed threshold (FT), stochastic restoration (SR) (Wang et al., 2022), or data-driven restoration (DR) (Brahma and Rai, 2023).
+The findings indicate that the dynamic nature of the AM module improves performance by 1.5, 1.0, and 0.7 mAP over FT set at 0.9, 0.8, and 0.7, respectively.
+While the randomness of SR and the gradient-only approach in DR limit the performance to 28.7 and 28.3 mAP, the ARR mechanism achieves the performance of 29.2 mAP.
+These results highlight the significance of the adaptability of AM and ARR over existing methods in dynamic environments.
+Additional ablation studies on hyper-parameters are provided in the appendix.
 5. CONCLUSION
-In this work, we proposed CTAOD, aimed at addressing the challenges of low-quality pseudo-labels and random errors resulting from stochastic recovery in continual test-tiem adaptation.
-Firstly, object-level contrastive learning leverages teacher ROI features for contrastive learning to refine the feature representation, tailored for object detection.
-Secondly, the dynamic threshold strategy selects higher-quality pseudo-labels by dynamically updating the category-specific threshold based on the predicted confidence scores.
-Lastly, the data-driven stochastic restoration mechanism selectively reset inactive parameters to mitigate forgetting while retaining valuable knowledge.
-The empirical results of four continual test-time adaptation tasks for object detection demonstrate the efficacy of CTAOD in both short-term and long-term adaption scenarios.
+In this work, we proposed AMROD to address the two challenges in Continual Test-Time Adaptation (CTTA).
+Firstly, object-level contrastive learning leverages ROI features for contrastive learning to refine the feature representation, tailored for object detection.
+Secondly, the adaptive monitoring module enables efficient adaptation and high-quality pseudo-labels by dynamically skipping unstable adaptation and updating the category-specific threshold, based on the predicted confidence scores.
+Lastly, the adaptive randomized restoration mechanism selectively reset inactive parameters to mitigate forgetting while retaining valuable knowledge.
+The empirical results of four CTTA tasks for object detection demonstrate the efficiency and efficacy of AMROD in both short-term and long-term adaptation.
 Appendix A Appendix Overview
-In this appendix, we provide additional details and analysis.
-In the first section, we provide implementation details on CTAOD and other baseline methodologies.
-In the second section C, we furnish complete results for the quantitative research and additional visualization examples for the qualitative research.
-In the third section, we further discuss the continual learning.
-In the last section E, we discuss the limitations and prospective directions of our research.
+We provide additional details and analysis in this appendix.
+In section B, we provide implementation details on AMROD and other baseline methodologies.
+In section C, we furnish complete results for the quantitative research, additional visualization examples, and additional ablation studies on hyper-parameters.
+In section D, we further discuss the continual learning.
+In section E, we discuss the limitations and prospective directions of our research.
 Appendix B Additional Implementation Details
@@
 Similarly, we trained another source model on the clear condition of the SHIFT (Sun et al., 2022) with 30 epochs.
-We will specify the implementation details of the CTAOD and other baseline methods in the following subsections.
-B.1. Implememtation of CTAOD
-Detailed hyper-parameters of CTAOD for each benchmark are listed in Table 7.
-Furthermore, we present a pseudo-code outline of our CTAOD framework in Algorithm 1.
-The three key components of CTAOD are highlighted, i.e., Object-level Contrastive Learning, Adaptive Monitoring, and Data-driven Stochastic Restoration.
+We will specify the implementation details of the AMROD and other baseline methods in the following subsections.
+B.1. Implememtation of AMROD
+Detailed hyper-parameters of AMROD for each benchmark are listed in Table 7.
+Furthermore, we present a pseudo-code outline of our AMROD framework in Algorithm 1.
+The three key components of AMROD are highlighted, i.e., Object-level Contrastive Learning, Adaptive Monitoring, and Adaptive Randomized Restoration.
 B.2. Implememtation of Baseline Methods
@@
 We implement them based on their public code, with some modifications to fit into the objection detection.
-Specifically, we implement the Cotta method by constructing a mean-teacher framework based on Faster R-CNN, which incorporates a stochastic restoration mechanism.
-Similarly, we implement SVDP with visual prompts on the area of the bounding box in pseudo-labels based on mean-teacher framework.
+Specifically, we implement the Cotta method by constructing a mean-teacher framework based on Faster R-CNN, which incorporates a fixed threshold for pseudo-labeling and a stochastic restoration mechanism.
+Similarly, we implement SVDP with visual prompts on the area of the bounding box in pseudo-labels based on the mean-teacher framework.
 Further, the MemCLR (VS et al., 2023b) and IRG (VS et al., 2023a) are source-free domain adaptation methods for object detection.
 Therefore, we implement the MemCLR and IRG based on their public code, with little modifications to follow the CTTA setting (i.e., updating the model parameters during evaluation).
-For a fair comparison, we set the same momentum $\alpha$ of the Exponential Moving Average (EMA) for CoTTA, SVDP, MemCLR, and IRG as CTAOD, all of which apply the mean-teacher framework (Tarvainen and Valpola, 2017).
+For a fair comparison, we set the same momentum $\alpha$ of the Exponential Moving Average (EMA) for CoTTA, SVDP, MemCLR, and IRG as AMROD, all of which apply the mean-teacher framework (Tarvainen and Valpola, 2017).
 Other hyper-parameters are retained as specified in their public code.
@@
 We present the complete experiment results on the two long-term CTTA tasks, i.e., Cityscapes-to-Cityscapes-C and Cityscapes-to-ACDC adaptation tasks, in Table 8 and 9.
-Experiments show that the proposed CTAOD can largely maintain strong performance in the long term.
+Experiments show that the proposed AMROD can largely maintain strong performance in the long term.
 C.2. Qualitative Results
-As shown in figure 4 - 8, we present more visualization results to compare the CTAOD and other baseline methods based on the final round adaption on the long-term Cityscapes-to-Cityscapes-C task.
-The results further validate the effectiveness of CTAOD in distinguishing more foreground object categories and better locating them, compared with other baseline methods.
+As shown in figure 4 - 8, we present more visualization results to compare the AMROD and other baseline methods based on the final round adaption on the long-term Cityscapes-to-Cityscapes-C task.
+The results further validate the effectiveness of AMROD in distinguishing more foreground object categories and better locating them, compared with other baseline methods.
 C.3. Additional Ablation Study
-CTAOD-unstop FT of 0.9 FT of 0.8 FT of 0.7 SR DR Mean
-1 $\checkmark$ $\checkmark$ 27.7
-2 $\checkmark$ $\checkmark$ 28.2
-3 $\checkmark$ $\checkmark$ 28.5
-4 $\checkmark$ $\checkmark$ 27.1
-5 $\checkmark$ $\checkmark$ 28.3
-6 $\checkmark$ 29.0
-Table 6.
-Additional Ablation Study.
-The “CTAOD-unstop” denotes CTAOD without dynamic skipping.
-The “FT” represents replacing dynamic threshold in CTAOD with a fixed threshold.
-The “SR” denotes the stochastic restoration (Wang et al., 2022), and the “DR” indicates the data-driven restoration (Brahma and Rai, 2023).
-All experiments are done on long-term Cityscapes-to-Cityscapes-C tasks.
-As presented in Table 6, we perform an additional ablation study to evaluate the effectiveness of the proposed Dynamic Threshold (DT) strategy in Adaptive Monitoring and Data-driven Stochastic Restoration (DSR).
-The findings indicate that the DT approach results in improvements of 1.3, 0.8, and 0.5 mAP over fixed thresholds set at 0.9, 0.8, and 0.7, respectively.
-While the stochastic restoration and data-driven restoration limit the performance to 27.1 and 28.3 mAP, the DSR mechanism raises the performance to 29.0 mAP.
+To investigate the impact of hyperparameters, we conducted an ablation study on four key values on the long-term Cityscapes-to-Cityscapes-C task, including $\beta_{s}=0.65,0.85$, $\delta_{s}=1.3,1.5$, $\beta_{t}=0.85,0.75$, and $\eta=0.002,0.0005$.
+The chosen hyper-parameter configuration in this paper of $\beta_{s}=0.75$, $\delta_{s}=1.4$, $\beta_{t}=0.85$, and $\eta=0.001$ yielded a performance of 29.2 mAP.
+The ablation results, corresponding to different configurations, produced the following mAP scores: 28.8, 29.2, 29.0, 28.9, 29.1, 29.0, and 28.9, respectively.
+These findings highlight that the selected hyperparameters offer a balanced performance and confirm the stability of CTAOD, as it demonstrates minimal sensitivity to hyperparameter variations.
 Hyper-parameter Description City2City-C-Short City2City-C-Long SHIFT-Short City2ACDC-Long
@@
 Faster R-CNN (Ren et al., 2015).
-Future work will extend to other advanced detector backbone networks to assess the universality and generalization capabilities of CTAOD.
-Moreover, the proposed DSR mechanism to reset the model every iteration increases the computational cost.
+Future work will extend to other advanced detector backbone networks to assess the universality and generalization capabilities of AMROD.
+Moreover, the proposed ARR mechanism to reset the model every iteration increases the computational cost.
 Future work could explore more efficient mechanisms for restoration, such as restoring the model only when significant changes in the distribution of the target domain are detected.
@@
 25:        Update teacher model $\theta_{t}^{tch}$ via EMA strategy
-26:       // 3. Data-driven Stochastic Restoration
+26:        // 3. Adaptive Randomized Restoration
 27:        Generate random matrix $R_{t}\sim\operatorname{Uniform}(0,1)$
@@
 (e) MemCLR
-(f) CTAOD (ours)
+(f) AMROD (ours)
 Figure 4.
 Complete qualitative results.
-We compare the detection results of the CTAOD and other baseline methods in the $10^{th}$ round of adaption to Brightness corruption on the long-term Cityscapes-to-Cityscapes-C task.
+We compare the detection results of the AMROD and other baseline methods in the $10^{th}$ round of adaption to Brightness corruption on the long-term Cityscapes-to-Cityscapes-C task.
 (a) Source
@@
 (e) MemCLR
-(f) CTAOD (ours)
+(f) AMROD (ours)
 Figure 5.
 Complete qualitative results.
-We compare the detection results of the CTAOD and other baseline methods in the $10^{th}$ round of adaption to Defocus corruption on the long-term Cityscapes-to-Cityscapes-C task.
+We compare the detection results of the AMROD and other baseline methods in the $10^{th}$ round of adaption to Defocus corruption on the long-term Cityscapes-to-Cityscapes-C task.
 (a) Source
@@
 (e) MemCLR
-(f) CTAOD (ours)
+(f) AMROD (ours)
 Figure 6.
 Complete qualitative results.
-We compare the detection results of the CTAOD and other baseline methods in the $10^{th}$ round of adaption to Fog corruption on the long-term Cityscapes-to-Cityscapes-C task.
+We compare the detection results of the AMROD and other baseline methods in the $10^{th}$ round of adaption to Fog corruption on the long-term Cityscapes-to-Cityscapes-C task.
 (a) Source
@@
 (e) MemCLR
-(f) CTAOD (ours)
+(f) AMROD (ours)
 Figure 7.
 Complete qualitative results.
-We compare the detection results of the CTAOD and other baseline methods in the $10^{th}$ round of adaption to Motion corruption on the long-term Cityscapes-to-Cityscapes-C task.
+We compare the detection results of the AMROD and other baseline methods in the $10^{th}$ round of adaption to Motion corruption on the long-term Cityscapes-to-Cityscapes-C task.
 (a) Source
@@
 (e) MemCLR
-(f) CTAOD (ours)
+(f) AMROD (ours)
 Figure 8.
 Complete qualitative results.
-We compare the detection results of the CTAOD and other baseline methods in the $10^{th}$ round of adaption to Snow corruption on the long-term Cityscapes-to-Cityscapes-C task.
+We compare the detection results of the AMROD and other baseline methods in the $10^{th}$ round of adaption to Snow corruption on the long-term Cityscapes-to-Cityscapes-C task.